### v1

# EVINCE: Optimizing Adversarial LLM Dialogues via Conditional Statistics and Information Theory

Abstract
This paper introduces EVINCE (Entropy and Variation IN Conditional Exchanges), a dialogue framework advancing Artificial General Intelligence (AGI) by enhancing versatility, adaptivity, and reasoning in large language models (LLMs). Leveraging adversarial debate and a novel dual entropy theory, EVINCE improves prediction accuracy, robustness, and stability in LLMs by integrating statistical modeling, information theory, and machine learning to balance diverse perspective exploration with strong prior exploitation. The framework’s effectiveness is demonstrated through consistent convergence of information-theoretic metrics, particularly improved mutual information, fostering productive LLM collaboration. We apply EVINCE to healthcare, showing improved disease diagnosis, and discuss its broader implications for decision-making across domains. This work provides theoretical foundations and empirical validation for EVINCE, paving the way for advancements in LLM collaboration and AGI development.
EVINCE: Optimizing Adversarial LLM Dialogues
via Conditional Statistics and Information Theory
Edward Y. Chang Computer Science Clinical Mind AI Stanford University echang@cs.stanford.edu Draft: 2/15/2024; Revised: 8/15/2024
1 Introduction
The pursuit of Artificial General Intelligence (AGI) remains a central goal of AI research. We propose a paradigm shift in this quest: utilizing multiple Large Language Models (LLMs) engaged in synergistic dialogues as a crucial step towards AGI. This approach, we contend, addresses key limitations of current AI systems and provides a novel pathway to more robust, versatile, and capable artificial intelligence. Specifically, our work targets three core AGI characteristics: versatility, adaptivity, and reasoning capability.
Current LLMs, despite their remarkable capabilities, face significant challenges, including hallucination (generating false or nonsensical information), bias (reflecting and potentially amplifying societal prejudices), and limited reasoning (difficulties in complex problem-solving and logical inference). We posit that multi-agent dialogue systems offer a promising avenue to address these challenges. By fostering diversity and debate among LLMs, these systems can mitigate biases and promote enhanced reasoning capabilities. Furthermore, the iterative nature of multi-round dialogues allows for continuous context enrichment, enabling LLMs to access more precise information and formulate more accurate responses, thus reducing the occurrence of hallucinations.
Our previous work, SocraSynth Chang (2023a), addresses LLM limitations through structured multi-agent dialogues. By leveraging both adversarial and collaborative interactions between LLMs, SocraSynth demonstrates quantifiable improvements across various domains. In healthcare, it achieves a 5% improvement in diagnostic accuracy Chang and et al. (2023), while in news analysis, it identifies and explains biases in articles Chang (2024c). These advancements showcase enhanced versatility, adaptivity, and reasoning capability—key AGI characteristics.
SocraSynth’s adversarial component promotes the exploration of diverse perspectives, while its collaborative component fosters rigorous reasoning to reach well-reasoned conclusions. This synergy has yielded measurable gains beyond healthcare and bias mitigation, extending to geopolitical analysis Chang (2023b), corporate planning Tsao (2023), investment banking Chang (2024b), and emotional behavior modeling Chang (2024a). These results demonstrate SocraSynth’s effectiveness in mitigating LLM limitations and achieving substantial performance improvements across various applications, highlighting its potential for advancing towards AGI’s generalized problem-solving capabilities.
Although SocraSynth has demonstrated effectiveness, it relies on “contentiousness” as a qualitative measure to moderate the linguistic behaviors of participating LLMs. For instance, a high contentiousness value (0.9 out of 1.0) might lead LLMs to challenge each other’s assumptions and propose alternative solutions, while a low value (below 0.3) could encourage them to synthesize their viewpoints and find common ground. While the concept of contentiousness has proven useful in guiding SocraSynth dialogues, its qualitative nature limits its precision and explainability.
In this paper, we extend SocraSynth by introducing three theoretical pillars to quantify “contentiousness” and moderate dialogues based on statistical and information theories. These pillars, collectively referred to as EVINCE (Entropy and Variation IN Conditional Exchanges), provide quantitative measures for justifiable and explainable multi-agent dialogue moderation and evaluation.
These pillars are:
  1. 1.
    Inclusiveness Exploration: We develop methods to ensure dialogues explore all potential perspectives. We use conditional statistics via in-context learning Xie et al. (2021) to “free” an LLM agent from its default “maximal likelihood” next-token prediction behavior, allowing it to adopt specific stances. We introduce a dual entropy optimality theory to balance the exploration of new ideas with adherence to priors, thus optimizing information exchange between agents for comprehensive and stable discourse.
  2. 2.
    Information Flow Dynamics: We introduce information theory-based metrics to quantify and optimize dialogue dynamics. These measure information diversity (entropy), novelty (statistical divergence scores), and inter-agent persuasion (mutual information). These metrics enable us to assess and enhance the quality and efficiency of information flow within the multi-agent system, fostering rich and productive exchanges.
  3. 3.
    Reasoning Quality and Coherence: We establish frameworks to assess the logical structure and coherence of multi-agent reasoning. This pillar evaluates argument validity, analytical depth, and dialogue coherence. We synergistically integrate our previously developed CRIT system Chang (2023c) into SocraSynth. CRIT, which combines Socratic methods with formal reasoning techniques, enhances our ability to conduct critical thinking through evaluating argument quality, information-source credibility, and overall “reasonableness” within the dialogue. This integration ensures that the collective reasoning of LLM agents is not only diverse but also logically sound and aligned with the dialogue’s objectives.
The core strength of SocraSynth and EVINCE in advancing towards AGI lies in their ability to enhance key AGI characteristics through multi-agent dialogues. By employing conditional statistics and information theory, they boost versatility and adaptivity, allowing LLMs to transcend their typical “maximal likelihood” behaviors and mimic how humans adapt their linguistic behaviors to complete tasks. The framework’s debate structure fosters advanced reasoning and critical thinking capabilities. Together, these advancements represent significant steps towards achieving the complex, intricate capabilities required for AGI.
The main contributions of this paper are:
  1. 1.
    EVINCE Framework Design: Unlike approaches that use debate merely to improve accuracy via redundancy, EVINCE facilitates information discovery, bias mitigation, and decision-making that requires both breadth and depth of information.
  2. 2.
    Theoretical Foundations: We establish a theoretical basis for EVINCE, rooted in conditional Bayesian statistics, mutual information, and dual entropy. These principles are applied to measure, monitor, and modulate collaborative LLM interactions, contributing to a deeper understanding of how LLMs can effectively cooperate for improved decision-making. The dual entropy theory is novel and ground-breaking, illustrating how a productive decision-making process should start with room for diverse input and stable objectives, and then, through information exchange, converge to optimal decision/prediction.
  3. 3.
    Empirical Validation: We provide empirical validation of EVINCE’s underlying theories, highlighting the framework’s effectiveness in balancing exploration and exploitation to enhance prediction accuracy. We also introduce a set of maxims derived from our empirical findings, offering practical guidance for optimizing mutual information and minimizing various divergence measures.
The remainder of this paper is structured as follows: Section 2 discusses related work and addresses common misconceptions about our approach. Section 3 presents the three pillars of EVINCE, including its maxims, theories, and algorithms. Section 4 provides empirical evidence of EVINCE’s benefits by revisiting and expanding upon previous SocraSynth experiments in healthcare and bias detection. Section 5 offers concluding remarks and discusses future directions.
2 Related Work
The core objective of adversarial debate, as embodied in SocraSynth with EVINCE, is to foster diverse opinions and challenge assumptions, ultimately leading to more comprehensive and informed decision-making. This contrasts with traditional ensemble learning methods, which prioritize error diversity for improved accuracy. This section surveys ensemble learning and adversarial learning, highlighting the unique aspects of EVINCE and its theoretical underpinnings.
2.1 Ensemble Learning
Ensemble methods, such as Bagging Breiman (1996), Boosting Freund and Schapire (1997), and Mixtures of Experts Jacobs et al. (1991), traditionally aim to enhance machine learning performance by combining predictions from multiple models to achieve error diversity. Similarly, early LLM debate frameworks, e.g., Michael et al. (2023); Chan et al. (2023); Liang et al. (2023); Du et al. (2023) have focused primarily on improving answer accuracy rather than fostering a comprehensive exploration of information for critical decision-making.
SocraSynth distinguishes itself from traditional ensemble methods by prioritizing the generation of diverse predictions over the mere avoidance of errors. This is achieved through a dynamic protocol that adaptively adjusts the “contentiousness” level of the debate, encouraging models to initially explore a wide range of perspectives and rigorously assess the quality of arguments. Through constructive exchange of ideas in subsequent iterations, SocraSynth converges on a more informed and robust decision. Its approach, grounded in EVINCE’s dual entropy theory, allows for the exploration of various ideas by one LLM while maintaining stability through the other, effectively preventing premature convergence on a narrow set of perspectives.
2.2 Addressing Common Misconceptions about Adversarial LLM Dialogue
The rapid growth in academic conference submissions has led to an expanded reviewer pool, potentially impacting review quality. In this context, some misconceptions about our work have arisen, which we address here:
  1. 1.
    Misconception: Multiple LLMs cannot mitigate hallucination.
    Some reviewers have expressed skepticism about the ability of multiple LLMs to reduce hallucination. However, our approach does not simply aggregate potentially flawed outputs. Instead, SocraSynth leverages the strengths of multiple models to cross-check, challenge, and refine information. This process is analogous to peer review in scientific discourse, where collective scrutiny enhances the reliability of conclusions.
  2. 2.
    Misconception: Multiple LLM dialogues cannot lead to knowledge discovery.
    Despite evidence provided in Chang (2023a) and Chang and Chang (2023), some reviewers remain skeptical about the potential for knowledge discovery through LLM dialogues. Our work demonstrates that carefully structured interactions between LLMs can indeed lead to novel insights and knowledge synthesis by leveraging the polydisciplinary knowledge representation inherent in LLMs.
  3. 3.
    Misconception: The approach lacks theoretical foundation.
    Our approach is grounded in conditional statistics and information theory, which we use to optimize LLM interactions. This theoretical foundation ensures that the dialogue converges towards more accurate and less biased outcomes, addressing concerns about the rigor of our methodology.
By addressing these misconceptions, we aim to clarify the scientific basis and potential of our approach. The synergistic interaction between LLMs in SocraSynth, coupled with EVINCE’s theoretical pillars, represents a novel method for enhancing the capabilities of language models. This approach has applications in reducing bias, mitigating hallucination, and potentially advancing the field towards more robust AI systems.
Metric Pros Cons Remedies
Cross Entropy (CE) Shore and Johnson (1980) Measures how well the predictions of one model fit the actual distribution of another model’s outputs (asymmetric). Computationally intensive especially with large models and data sets; sensitive to the exact nature of probability distributions. Optimize computation strategies; use approximations or sampling methods to manage large data sets or complex models.
Entropy Shannon
Shannon (1948)
Indicates level of diversity; high suggests exploration of possibilities, and low for confidence on few choices High entropy might indicate noise rather than useful diversity; low entropy might mask important variability. Use critical reading methods (Appendix A) to assess argument quality; implement noise detection to differentiate between useful diversity and noise.
Jensen-Shannon
Divergence (JS)
Lin (1991)
Symmetric and bounded (0 to 1), providing an interpretable measure of distributional differences. May be less sensitive to small differences between distributions. Increase sensitivity settings or resolution of the metric; combine with other metrics to capture finer distinctions between distributions.
KL Divergence
Kullback (1951)
Measures difference between two probabilistic distributions. Asymmetric; not well-defined if a distribution has zero probabilities Use smoothing techniques to avoid zero probabilities; consider symmetric alternatives like JS divergence
Mutual Info (MI) Shore and Johnson (1980) Measures reduction of uncertainty; symmetric. Does not indicate the directionality of information flow. Supplement with directional information metrics; normalized with max entropy of A and B.
Wasserstein
Distance (WD)
Kantorovich (1942)
Direct measure of how similar or different the model outputs are; it depicts symmetric relationship. Not bounded but can be normalized or bounded for consistent interpretation. Define context-specific bounds for low, medium, and high divergence; consider normalizing it for non-directional comparisons.
Table 1: Summary of metrics for assessing LLM debates (equations are presented in Appendix D)
3 Algorithm, Maxims, and Theories
Problem Statement:
Organize a structured debate between two equally competent large language models (LLMs), LLMA and LLMB, to conduct $t$ rounds. At each round $t$, each model produces a probability distribution, denoted as $P_{A}^{(t)}$ and $P_{B}^{(t)}$, over $C$ possible outcomes, accompanied by supporting arguments $R_{A}^{(t)}$ and $R_{B}^{(t)}$. The goal is to design an iterative debate process that leverages the structured exchange of arguments to enable the models to converge on an optimal prediction distribution $P^{*}$ across the $C$ classes. Optimality is defined as achieving the highest possible accuracy on the prediction task.
To address this problem, we first introduce key metrics from information theory that form the basis of our approach.
3.1 Preliminaries in Information Theory
This section presents the key metrics used to measure information diversity, similarity, divergence, and other relevant factors within EVINCE. These metrics serve three primary objectives:
3.1.1 Fostering diversity of perspectives while ensuring reasoning quality
  • Wasserstein Distance (WD): Measures distribution difference between predictions to identify exploration opportunities Kantorovich (1942); Rubner et al. (2000); Villani (2008).
  • Shannon Entropy or Relative Entropy: Measures diversity of perspectives Cover and Thomas (2006); Shannon (1948).
  • Reasoning Quality: The CRIT algorithm evaluates the logical soundness and persuasiveness of supporting arguments Chang (2023c), helping to identify and mitigate hallucinations and poorly-reasoned arguments.
3.1.2 Exploring new possibilities while adhering to the dialogue subject
  • Correlation Coefficients: Tracks the evolution of opinions and assesses debate stability Brown et al. (2005) toward the goal of the dialogue.
  • Mutual Information (MI): Quantifies information overlap to ensure focused and productive debates Cover and Thomas (2006) and measures the degree of agreement/disagreement.
3.1.3 Examining information convergence and establishing termination criteria
  • Jensen-Shannon (JS) Divergence: Assesses similarity between probability distributions Lin (1991) (symmetric).
  • Cross Entropy (CE): Measures the asymmetric difference between prediction distributions Shore and Johnson (1980).
  • Kullback-Leibler (KL) Divergence: Reveals asymmetric differences between probability distributions Kullback (1951).
Table 1 summarizes these metrics and their pros and cons. Next, we present the EVINCE algorithm and explain how these metrics are used to moderate an LLM dialogue to achieve a balance between exploration and exploitation, leading to optimal prediction outcomes. In Section 3.3, we will introduce several maxims and prove the dual entropy theorem, which provides a theoretical foundation for the EVINCE algorithm.
[…]
Figure 1: Specifications of Algorithm EVINCE. Key points: 1) Asymmetric Start: In Step #1, LLMA initiates with opening arguments based solely on the given information, while LLMB starts with access to LLMA’s prediction and arguments for refutation. 2) Termination Criteria: The while loop in Step #2 considers three factors: Wasserstein distance, mutual information, and argument quality. EVINCE terminates if the dialogue ceases to make significant progress. 3) Further Details: Maxims #1 to #4 provide additional explanations of the algorithm’s principles. 4) Argument Evaluation: For details on the CRIT algorithm used to assess argument quality, refer to Chang (2023c).
3.2 EVINCE Algorithm Specifications
Figure 1 formally specifies the detailed operations of EVINCE.
Setup and Initialization:
EVINCE employs two equally competent LLM instances, $LLM_{A}$ and $LLM_{B}$, which can be different models (e.g., GPT and Claude) or independent instances of the same model. Given an information set $S$ and a class-label set $C$, EVINCE outputs a probability distribution over $C$ with justifications. For example, $S$ could represent a patient’s symptoms and $C$ a set of diseases, with EVINCE moderating a dialogue to predict the patient’s disease(s).
The debate “contentiousness” is initially set to 90% to encourage disagreement and explore diverse perspectives. This high level is maintained for the first two iterations, then gradually reduced by $x\%$ in each subsequent iteration to shift from exploration to exploitation [^1] The setting of $x$ at the beginning of a dialogue iteration can depend on the values of all moderation subroutines. Our empirical experience suggests $x=25\%$ because a typical dialogue concludes in five iterations, and $x=20\%$ would set contentiousness to be 90%, 90%, 68%, 50%, and 38%.. Table 2 demonstrates how GPT adapts its linguistic behavior based on the “contentiousness” parameter, showcasing the model’s ability to adjust to different contextual settings. Importantly, this table is generated by GPT itself, not by human moderators, as detailed in Chang (2023a) and further explored in Section 4.
C.L. Tone Emphasis Language
$0.9$ Highly confrontational; focused on raising strong ethical, scientific, and social objections. Flagging risks and downsides; ethical quandaries, unintended consequences, and exacerbation of inequalities. Definitive and polarizing, e.g., “should NOT be allowed,” “unacceptable risks,” “inevitable disparities.”
$0.7$ Still confrontational but more open to potential benefits, albeit overshadowed by negatives. Acknowledging that some frameworks could make it safer or more equitable, while cautioning against its use. Less polarizing; “serious concerns remain,” “needs more scrutiny.”
0.5 Balanced; neither advocating strongly for nor against gene editing. Equal weight on pros and cons; looking for a middle ground. Neutral; “should be carefully considered,” “both benefits and risks.”
0.3 More agreeable than confrontational, but maintaining reservations. Supportive but cautious; focus on ensuring ethical and equitable use. Positive but careful; “transformative potential,” “impetus to ensure.”
0.0 Completely agreeable and supportive. Fully focused on immense potential benefits; advocating for proactive adoption. Very positive; “groundbreaking advance,” “new era of possibilities.”
Table 2: Changes in linguistic behaviors of LLMs at different contentiousness levels.
Moderation Subroutines:
EVINCE employs four key subroutines to manage the exploration-exploitation tradeoff, ensuring information diversity, quality, and stability:
  • CRIT: Evaluates argument quality and source credibility. Low scores can trigger dialogue termination.
  • WD (Wasserstein distance): Assesses prediction diversity, expected to decrease over time.
  • MI (Mutual Information): Evaluates dialogue convergence. Stagnation can trigger termination.
  • Additional metrics: KL divergence, Jensen-Shannon divergence, and cross entropy ensure evaluation consistency (see Table 1).
When all metrics reach a plateau, indicating no further improvement, EVINCE terminates the dialogue. At this point, both LLM instances are prompted to jointly deliver a conciliatory conclusion with comprehensive arguments and counterarguments. Additionally, they provide a list of missing information that, if obtained, could potentially improve prediction accuracy and reliability.
Dialogue Iterations (Steps 1 and 2):
Given the contentiousness level set for a dialogue iteration, each LLM generates a new prediction distribution on $C$ with arguments, based on the other’s last prediction and accumulated reasons as the context.
Final Output (Step 3):
  • Combine final predictions from both LLMs.
  • Weight predictions based on supporting argument quality (evaluated by CRIT).
  • Handle uncertainty by soliciting missing information when final prediction entropy is high (see Chang (2024b) for RAG application).
EVINCE creates a structured debate between AI models, balancing diverse viewpoints with consensus-building to produce accurate, well-reasoned predictions.
3.3 Maxims with Theoretical Foundations
Progress towards the optimality goal is guided and measured by metrics introduced in Section 3.1. This section explains how these metrics can be used in complementary ways to facilitate proper trade-offs between diversity and convergence, exploration and exploitation, and several other factors. In the EVINCE algorithm presented in Figure 1, we have annotated the steps to which these four maxims are applied.
Maxim #1: Orchestrate Two Equally Competent LLMs in Structured Debate:
Integrating two equally competent LLMs ensures a balanced exchange of insights and avoids bias. This adversarial setup fosters diversity in predictions, each supported by justifications, promoting critical evaluation and uncovering potential blind spots.
How? Choosing LLMs with comparable performance on a shared validation set, a balanced debate can be ensured. Suitable models include GPT-4, Claude, and Gemini. Conditioning different instances of the same LLM to support opposing stances on a subject matter can also be effective due to the theoretical justification of in-context learning with conditional Bayesian statistics Xie et al. (2021).
Maxim #2: Encourage the Accurate Rather Than the “Popular” Prediction:
Typically, LLMs, with their maximum likelihood next-token prediction objective, tend to favor the most popular predictions. By conditioning LLMs within specific contexts, we can prioritize specific stance over popularity, mitigating confirmation biases.
How? Using the proxy metrics in Table 1, EVINCE dynamically adjusts the “contentiousness” level in debates. These metrics quantify agreement, diversity, and mutual information, promoting productive information exchange and enhancing prediction quality.
Maxim #3. Evaluating the Convergence Rate of the Predictions Across the Iterations:
This aspect focuses on measuring how quickly and effectively the predictions from the LLMs converge over successive rounds, assessing the efficiency of the debate and aggregation mechanisms.
How? Convergence is assessed by measuring mutual information and using proxy metrics such as Wasserstein distance and cross entropy. When the mutual information is high or the similarity between predictions is high, the debate is considered to be converging.
Maxim #4. Combine Predictions Weighted by Diversity and Quality:
Weighting the probability distributions from two LLMs based on diverse probabilistic insights and argument quality.
How? Following these three maxims:
  • Maxim #4.1 Prediction Reliability: Estimate the reliability of predictions using entropy-based measures to quantify uncertainty and information content. Typically, lower entropy indicates higher confidence in a prediction, suggesting higher reliability.
  • Maxim #4.2 Argument Quality: Evaluate the quality of supporting arguments using techniques inspired by the Socratic method. This includes identifying logical fallacies, assessing the relevance and credibility of evidence.
  • Maxim #4.3 Aggregation: Employ a weighted aggregation method, such as a Bayesian model to combine weighted predictions accounting for both probabilistic insights and the quality of supporting arguments.
3.4 Entropy Duality Theorem (EDT)
Theorem EDT: Optimal Pairing of LLMs for Probabilistic Prediction Accuracy.
The optimal pairing of LLMs for diagnosis accuracy, in terms of stability, accuracy, and robustness, occurs when the LLMs are 1) equivalent in the quality of the information they process, and 2) exhibit contrasting entropy values in their prediction distributions—one high and one low.
Proof.
Please see Table 3. ∎
Theorem EDT: Optimal Pairing of LLMs for Probabilistic Prediction Accuracy.
The optimal pairing of LLMs for diagnosis accuracy, in terms of stability, accuracy, and robustness, occurs when the LLMs are 1) equivalent in the quality of the information they process, and 2) exhibit contrasting entropy values in their prediction distributions—one high and one low.
[Proof]:
[…]
Step 6: Conclude the proof. By combining the prediction distributions of LLMA and LLMB, with one having high information entropy and the other low, we achieve an optimal balance that maximizes the lower bound of the combined information entropy. This balance between exploration (high entropy) and exploitation (low entropy) optimizes overall prediction accuracy and robustness, as proved through information theory and the properties of entropy. Thus, the theorem is established. $\square$
Table 3: Proof of Entropy Dual Theorem (EDT)
4 Empirical Study
This empirical study investigates the application of EVINCE to disease diagnosis, leveraging large language models (LLMs) as diagnostic tools. We aim to validate the following three hypotheses:
  1. 1.
    Contentiousness & Prediction Quality: Initial LLM disagreement (measured by Wasserstein distance) increases with higher initial contentiousness but decreases as debate progresses. Individual LLM prediction uncertainty (Shannon entropy) will follow a similar pattern.
  2. 2.
    EDT Effectiveness & Confusion Matrices: LLM pairs following the Entropy Duality Theorem (EDT) will have complementary error patterns, leading to higher combined prediction accuracy than non-EDT pairs.
  3. 3.
    EVINCE & Historical Misdiagnoses: EVINCE, applied to real-world data, will improve diagnostic accuracy and identify potential misdiagnoses or ambiguities within the ground truth.
(a) GPT4 pairs Claude
(b) GPT4 pairs Gemini
Figure 2: Pre-/post-debate accuracy on all patients on all diseases shows EVINCE helps
Problem Statement:
Given a set of symptoms $S$ and a context $\kappa$, the goal is to predict a probability distribution of top-$k$ diseases over a set of $C$ possible diseases. This is represented as $P=\text{LLM}(S,\kappa)$, where each LLM generates top-$k$ predictions on $C$ ($k\leq|C|$) based on the input symptoms $S$ and context $\kappa$.
$P=\left(p(\text{top}\,1\,\text{to}\,k\in C\mid S,\kappa)\right).$
Context $\kappa$ is where dual entropy is adjusted through three parameters: temperature, the value of $k$ in top-$k$, and the contentious level $\Delta$. A distribution tends to have high entropy when all three parameters are set high, and vice versa.
Resources, Dataset & Data Preparation:
Our study utilizes a dataset from Kaggle Patil (2020), comprising 4,921 patient records. Each record includes the diagnosed disease and up to 17 symptoms (e.g., fever, cough, fatigue, itchiness, difficulty breathing). After removing duplicates, we obtain 304 unique diagnostic instances spanning 40 diseases. (The refined dataset is provided as supplementary data.) Each instance serves as a test case where EVINCE leverages the inherent knowledge of LLMs (GPT-4, Gemini, and Claude3) without training them through few-shot techniques on this specific dataset. Our computing resources are sponsored by Azure through a Stanford grant.
Evaluation:
We assess prediction quality using the top-k Mean Reciprocal Rank (MRR). If one of the top-k predicted diseases matches the ground truth diagnosis, the score is the reciprocal of its rank (1 for the top prediction, 1/2 for the second, 1/3 for the third, etc.). If none of the top-k predictions are correct, the score is 0.
4.1 Study #1: Post vs. Pre-Debate Accuracy
We employed GPT-4, Gemini, and Claude3 to perform independent disease predictions on 304 patient instances, then used EVINCE to pair them and evaluate performance gain.
Experimental Setup:
We set $k=5$ for both LLM agents, with one agent at high temperature and the other at low temperature. The contentiousness level was set very high ($\Delta=0.9$ out of 1) to encourage significant cross entropy. Setting $k=5$ ensures some minimal common ground, fostering meaningful interaction. High contentiousness promotes counterarguments and information exchange.
Pre- and Post-Debate Evaluation:
We conducted two sets of experiments:
1. Constrained Prediction (Baseline): We limited disease predictions to the 40 labels in the dataset, mimicking common supervised learning assumptions. This yielded high accuracy (95-97%) but is unrealistic for real-world diagnosis where general practitioners consider all possibilities. This constraint highlights LLMs’ flexibility, as they’re less prone to overfitting erroneous labels (further discussed in subsequent studies).
2. Unconstrained Prediction: We removed the label constraint to better simulate real-world conditions. All 304 patient cases yielded stable results across GPT-4, Gemini-3, and Claude-3, with a standard deviation of just 1.5%. Prior to debate (light blue bars in Figure 2), GPT-4 led in accuracy (82.8%), followed by Gemini (80.3%) and Claude (79.5%).
EVINCE Performance:
Implementing EVINCE with GPT-4 and Claude-3 pairing and GPT-4 and Gemini-3 pairing consistently improved accuracy by 4-5 percentage points (green bars in Figure 2). The GPT-4 and Claude-3 pairing achieved 87.5% accuracy (Figure 2(a)), rivaling state-of-the-art clinical performance like the REFUEL algorithm Peng et al. (2018).
Discussion:
The remaining 12.5% inaccuracy for the GPT-Claude pairing might not be solely attributable to EVINCE. Considering the potential 11% US misdiagnosis rate reported by Johns Hopkins Newman-Toker et al. (2023b), this discrepancy could indicate mislabeled data in the original dataset. This presents a groundbreaking opportunity: EVINCE could potentially identify and correct errors in existing datasets, a concept we explore further in Section 4.3.
4.2 Study #2: Confusion vs. Opportunities
(a) GPT liver c-matrix
(b) Claude liver c-matrix
Figure 3: Confusion matrices
(a) Entropy
(b) WD %
(c) Norm. MI
Figure 4: Entropy, WD, and normalized MI
Two primary factors contribute to EVINCE’s improved diagnostic accuracy: (1) structured debates with reasoning encourage LLMs to explore alternative diagnoses in both breadth and depth, leading to more comprehensive analysis and decision-making (see Appendices B and C); and (2) pairing high- and low-entropy LLMs balances exploratory diversity with exploitative stability, resulting in more robust and high-quality decisions, as demonstrated in this second study.
Analysis of Confusion Matrices:
We use confusion matrices to analyze two LLMs’ performance in diagnosing Hepatitis types A to E (Fig. 3):
  • GPT-4 shows limited accuracy, particularly for types C and D, achieving only 50% accuracy for types A and B.
  • Claude exhibits a wider spread of predictions across all Hepatitis types.
These matrices highlight how Claude’s flexibility in exploring diverse diagnostic hypotheses aids the debate process. Claude’s initial uncertainty (high entropy) brings new information, potentially challenging and correcting GPT-4’s more confident (low entropy) predictions. This dynamic interplay exemplifies EVINCE’s balance between exploration and exploitation, leading to more accurate and comprehensive diagnoses.
Observations from Information Metrics:
  • Entropy Stabilization: Figure 4(a) shows entropy levels of both LLMs stabilizing after three debate rounds, indicating convergence towards a similar, stable entropy state.
  • Wasserstein Distance Improvement: Figure 4(b) demonstrates consistent improvement in Wasserstein distance (WD) between the two models’ predictions over successive rounds.
  • Mutual Information Increase: Figure 4(c) reveals a 14% improvement in normalized mutual information (MI) between GPT-4 and Claude’s prediction distributions, suggesting increased shared information throughout the debate.
  • Convergence of Divergence Metrics: Figure 5 shows consistent convergence of all divergence metrics.
Comparative Performance:
EVINCE demonstrates a 5% higher accuracy rate in diagnosing specific types of liver diseases compared to a baseline approach (Figure 2(a)), underscoring its capability to handle complex diagnostic scenarios effectively.
Figure 5: Convergence of all metrics
4.3 Study #3: Ground-Truth Remediation
This study illustrates how EVINCE can identify potential misdiagnoses, explain the reasoning behind them, and recommend corrective actions. Traditionally, machine learning scientists rely on labeled data as “ground truth.” However, as evidenced by research like that of Newman-Toker et al. (2021) Newman-Toker et al. (2023a) from Johns Hopkins, misdiagnosis is a widespread issue in healthcare systems globally. These erroneous diagnoses, often treated as ground truth, can be perpetuated by supervised learning algorithms, exacerbating the problem within the healthcare system.
(a) GPT & Claude top-5
(b) WD %
(c) Mutual Info.
Figure 6: Remediation: Jaundice to Hepatitis
In the debate scenario detailed in Appendix C, where Jaundice is the ground truth diagnosis, Figure 6(a) illustrates initial differences between GPT-4 and Claude’s predictions. Jaundice is absent in GPT-4’s top-5 (with 0% in red), while ranked third by Claude. Although Claude influences GPT-4 to include Jaundice in its third prediction in the second round, subsequent rounds see both LLMs drop Jaundice to the fourth position of 10%.
Meanwhile, Hepatitis A, initially GPT-4’s top prediction (30% in dark blue), is quickly demoted to fifth and eventually drops out of the top-5 entirely due to Claude’s influence. Hepatitis B, initially ranked second by GPT-4 and top by Claude, stabilizes in the second position in rounds 3 and 4 (in light blue). Notably, Hepatitis C rises from second place on both lists to the top position and remains there (in black).
Figure 7: Convergence of all metrics
As demonstrated in the previous case study, Wasserstein distance (WD) effectively measures the divergence between LLM predictions and assesses debate convergence. Figures 6(b) and 6(c) show that WD stabilizes after three debate rounds, coinciding with a plateau in normalized mutual information (MI) between GPT-4 and Claude. This stabilization suggests their predictions converge.
Figure 7 illustrates the convergence of all divergence metrics—including Jensen-Shannon divergence, cross-entropy, and Kullback-Leibler divergence—particularly between the second and third rounds. Although the final joint prediction for Hepatitis C reached a high consensus of 37.5%, it deviates from the actual condition of Jaundice, which the Kaggle dataset reports with 10% confidence. EVINCE provides general practitioners with alerts and suggests remedial actions (see Appendices D.9 and C.8) to address this discrepancy. Recommended actions include querying additional symptoms from the patient and conducting specific laboratory tests.
4.4 Experiment Remarks
EVINCE initiates debates with high contentiousness, encouraging dual prediction entropy between LLMs, as supported by the EDT theorem. It utilizes normalized mutual information (MI) to track shared knowledge accumulation throughout the debate, while Wasserstein distance (WD) and Jensen-Shannon divergence (JSD) quantify dissimilarity between LLM predictions.
These metrics (EDT, WD, JSD, MI) provide a comprehensive view of debate progress. WD and JSD assess the potential for further communication and refinement, while MI monitors shared understanding, aiding in determining the optimal stopping point.
The asymmetric nature of KL divergence and cross entropy warrants further investigation. Despite eventual convergence in our case studies, discrepancies observed in the second round (one direction increasing while the other decreases) suggest potential value in exploring asymmetric information. Future work will re-evaluate the use of these metrics if asymmetry proves beneficial.
Besides generating final joint disease predictions, EVINCE provides:
  • Recommendations for additional symptom inquiries and lab tests to improve accuracy.
  • Suggestions to query symptom onset, duration, severity, trends, and associated symptoms (documented in Appendices B.8 and C.9).
These recommendations have been verified by general practitioners to be valuable.
5 Concluding Remarks
This paper introduces EVINCE, a framework designed to facilitate collaborative dynamics among Large Language Models (LLMs) through structured, adversarial debates. Our research demonstrates that EVINCE significantly advances the pursuit of Artificial General Intelligence (AGI) by enhancing three core characteristics: versatility, adaptivity, and reasoning capability. By addressing key limitations of current AI systems, including hallucination, bias, and limited reasoning, EVINCE provides a novel pathway towards more robust and capable artificial intelligence.
The core strength of EVINCE lies in its three theoretical pillars: Inclusiveness Exploration, Information Flow Dynamics, and Reasoning Quality and Coherence. These pillars, grounded in conditional statistics and information theory, enable LLMs to transcend their typical “maximal likelihood” behaviors, mirroring human adaptability in linguistic tasks. The integration of the CRIT system, which combines Socratic methods with formal reasoning techniques, further enhances critical thinking and ensures logically sound and objective-aligned collective reasoning.
Our empirical validation demonstrates EVINCE’s effectiveness in improving prediction accuracy across various domains, notably achieving a 5% improvement in medical diagnosis tasks. The framework has also shown promise in identifying biases in news articles Chang (2024c), showcasing its potential for broader applications in fields such as geopolitical analysis Chang (2023b) corporate planning Tsao (2023), and emotional behavior modeling Chang (2024a).
While EVINCE has demonstrated significant potential, our future work will focus on refining key aspects of the system. We aim to enhance the contentiousness parameter by tying it more rigorously to metrics such as Wasserstein distance, mutual information, and CRIT scores. Additionally, we plan to explore alternative methods for inducing dual-entropy conditions, moving beyond current parameters like temperature and top-k selection.
Comprehensive ablation studies will be crucial in determining optimal configurations for various applications, ensuring EVINCE’s adaptability to diverse contexts and challenges. Through these continued refinements and explorations, we aim to push the boundaries of collaborative AI frameworks, bringing us closer to realizing the full potential of artificial general intelligence.
In conclusion, EVINCE represents a significant step forward in AI research, offering a structured approach to leveraging the collective intelligence of multiple LLMs. By addressing current limitations and enhancing key AGI characteristics, EVINCE paves the way for more versatile, adaptive, and capable AI systems, bringing us closer to the goal of artificial general intelligence.
Acknowledgements
This work has been significantly shaped by insights from several key figures in the field. Microsoft CSO Eric Horvitz’s concept of “polydisciplinary” representation in LLMs was a major influence. Demis Hassabis of DeepMind provided valuable input on end-to-end architecture without human interference. Prof. Stuart Russell’s assertion that LLMs mimic human linguistic behavior to achieve goals was pivotal in our approach to modeling linguistic behaviors and utilizing Bayesian frameworks.
My colleague Prof. Vaughan Pratt’s feedback on the initial paper inspired the use of “contentiousness” to foster adversarial debate, a key component of our framework. I am grateful to Prof. Monica Lam for her collaborative support throughout this research. Additionally, I thank my colleagues at Stanford Clinical Mind AI for their invaluable domain knowledge in disease diagnosis.
Appendix A: Evaluative Phase of EVINCE
EVINCE uses the Socratic method to evaluate the “reasonableness” of a set of arguments that support a subject matter. The Socratic method is a questioning technique used in teaching and philosophy to encourage critical thinking and self-discovery Wikipedia (2023). The method involves asking a series of questions to explore complex ideas and help individuals arrive at their own understanding of a concept. It is based on the belief that knowledge cannot be simply imparted, but must be discovered through a process of questioning and dialogue.
To illustrate how these methods can practically be applied, let’s use the example of critical reading. Critical reading is a crucial component of critical thinking, which involves evaluating the quality and credibility of written materials, from research papers to blog posts Lai et al. (2017); Paul and Binker (1990). It requires a systematic and analytical approach, asking relevant questions, and using effective prompts to gain deeper understanding of the text Elder and Paul (2010).
Function $\Gamma$ = CRIT($d$)
Input. $d$: document; Output. $\Gamma$: validation score;
Vars. $\Omega$: claim; $R$ $R^{\prime}$: reason & counter reason set;
Subroutines. $Claim$(), $FindDoc$(), $Validate$();
Begin
#1   Identify in $d$ the claim statement $\Omega$;
#2   Find a set of supporting reasons $R$ to $\Omega$;
#3   For $r\in R$ eval $r\Rightarrow\Omega$
     If $Claim$($r$), ($\gamma_{r}$, $\theta_{r}$) = CRIT($FindDoc$($r$));
     else, ($\gamma_{r}$, $\theta_{r}$) = $V$($r\Rightarrow\Omega$);
#4   Find a set of rival reasons $R^{\prime}$ to $\Omega$;
#5   For $r^{\prime}\in R^{\prime}$, ($\gamma_{r^{\prime}}$, $\theta_{r^{\prime}}$) = V($r^{\prime}\Rightarrow\Omega$) eval rival arguments;
#6   Compute weighted sum $\Gamma$, with $\gamma_{r}$, $\theta_{r}$, $\gamma_{r^{\prime}}$, $\theta_{r^{\prime}}$.
#7   Analyze the arguments to arrive at the $\Gamma$ score.
#8   Reflect on and synthesize CRIT in other contexts.
End
Table 4: CRIT Pseudo-code. (The symbol $\Rightarrow$ denotes both inductive and deductive reasoning.)
To aid in critical reading, we introduce a prompt template called CRIT Chang (2023c), which stands for Critical Reading Inquisitive Template. Given a document $d$, CRIT evaluates it and produces a validation score $\Gamma$. Let $\Omega$ denote the conclusion or claim of $d$, and let $R$ be the set of reasons supporting the claim. We define ($\gamma_{r},\theta_{r}$) = V($r\Rightarrow\Omega$) as the causal validation function, where $\gamma_{r}$ denotes the validation score, $\theta_{r}$ the source credibility score, for each reason-to-conclusion argument $r\Rightarrow\Omega$. Table 4 presents the pseudo-code of $\Gamma$ = CRIT($d$), which generates the final validation score $\Gamma$ for document $d$ with justifications.
EVINCE uses CRIT to evaluate argument quality of the participating LLMs involved in the debate. The input to CRIT from each LLM is first its stance on the debate subject, e.g., a set of predicted diseases, and the arguments are its reasons to arrive at the prediction. Each document in the case of EVINCE is the prediction set as the conclusion $\Omega$, the arguments as set $R$, and the opposing LLM’s counterarguments as $R^{\prime}$. With this document, CRIT is able to produce validity and credibility scores in $\Gamma$ for the LLM.
For detailed prompts, examples, and an empirical study verifying the effectiveness of CRIT, please consult Chang (2023c).
Appendix B: EVINCE Debate - Dengue Fever vs. Chikungunya
This experiment addresses a diagnostic challenge involving multiple potential diseases. A patient presents with the following 12 symptoms: [skin rash, joint pain, vomiting, fatigue, high fever, headache, nausea, loss of appetite, pain behind the eyes, back pain, malaise, muscle pain, red spots over the body]. In this case, GPT-4 provides the initial diagnosis, with Gemini following suit. The confirmed diagnosis for this experiment is Dengue Fever. Given the satisfactory diversity in predictions from the two LLMs, the debate progresses and converges smoothly.
B.1 Moderator’s Prompt
“Identify potential diseases linked to the symptoms (skin rash, joint pain, vomiting, etc.). Provide top-3 predictions with justifications, supplementary symptom inquiries, and relevant lab tests.”
B.2 GPT-4’s Opening Round
Top-3 Predictions: Dengue Fever (60%), Chikungunya (25%), and Zika Virus (15%).
Justifications: Most symptoms are classic signs of Dengue Fever. Red spots over the body could correspond to petechiae, often seen in Dengue due to thrombocytopenia (a decrease in platelet count).
Chikungunya can cause similar symptoms to Dengue, such as fever, severe joint pain, muscle pain, headache, nausea, and rash. Joint pain from Chikungunya is often very debilitating and can be more prolonged than in Dengue.
Zika Virus symptoms can include fever, rash, joint pain, malaise, and conjunctivitis, with a milder presentation compared to Dengue or Chikungunya. The presence of a rash and conjunctivitis is notable, although Zika is less likely to cause severe muscle pain or high fever.
B.3 Gemini’s Opening Round
Top-3 Predictions: Viral Infection (60%), Autoimmune Disease (20%), Bacterial Infection (15%).
Justifications: Symptoms common in viral infections. Autoimmune diseases is possible, but less likely due to vomiting and nausea. Bacterial infection is less probable given rash and muscle pain.
B.4 GPT-4’s Second Round
Top-3 Predictions: Dengue Fever (60%), Chikungunya (25%), and Zika Virus (15%).
Justifications: GPT-4 maintains its predictions, countering Gemini’s assessment with three points. First, while viral infections can cause these symptoms, the specific combination suggests a more targeted diagnosis. Second, the acute onset of high fever, joint pain, and rash is less typical for autoimmune diseases. Finally, the rash and red spots are not characteristic of bacterial meningitis.
B.5 Gemini’s Second Round
Top-3 Predictions: Dengue Fever (50%), Chikungunya (30%), and Viral Infection (20%).
Justifications: Gemini’s initial prediction considers a Viral Infection as the most likely cause due to its broad range of potential symptoms. However, GPT-4’s focus on Dengue Fever and Chikungunya, with their specific symptom profiles, provides a more precise explanation for the patient’s condition. Here’s what is lacking in the first set of predictions:
  • Specificity: While a Viral Infection could explain most symptoms, it does not account for joint pain and the presence of a rash, which are more characteristic of DF or Chikungunya.
  • Detailed justifications: The first analysis does not delve as deeply into the reasons behind each disease possibility. GPT-4’s explanation for Dengue Fever and Chikungunya is more comprehensive, strengthening the connection between symptoms and diagnosis.
  • Importance of additional information: Emphasize how geographic location and travel history can significantly impact the likelihood of these specific diseases.
B.6 GPT-4’s Third Round
Top-3 Predictions: Dengue Fever (60%), Chikungunya (35%), and Zika Virus (5%).
Justifications: Though GPT-4 slightly adjusts probabilities, it considers the adjustment arbitrary without new information that would specifically shift the likelihood between Dengue and Chikungunya. GPT-4 refutes Gemini’s Viral Infection prediction as lacking specificity. The symptoms described are quite distinctive and align closely with specific viral infections transmitted by mosquitoes.
B.7 Gemini’s Third Round
Top-3 Predictions: Dengue Fever (60%), Chikungunya (35%), and Zika Virus (5%).
Justifications: Agreeing that GPT’s critique is spot-on, and reiterating the need for more info.
B.8 Final Joint Recommendations to GP
After seeking further input from GPT-4, Gemini fully concurs with GPT-4’s final suggestions, even though the moderator instructs Gemini to maintain its original stance. Both LLMs jointly recommend additional symptom examinations and lab tests:
A comprehensive diagnostic approach for Dengue Fever requires a combination of clinical assessment, laboratory testing, and consideration of the patient’s travel and vaccination history, as well as the epidemiological context. Clinical symptoms such as skin rash, joint pain, vomiting, and red spots over the body provide valuable clues to the diagnosis, but additional symptoms like nausea, vomiting, diarrhea, bleeding, bruising, and drowsiness can indicate the severity and progression of the disease.
Recommended Laboratory Tests:
  • Complete Blood Count (CBC): Identify thrombocytopenia or leukopenia.
  • Serology Tests: Detect specific IgM and IgG antibodies for Dengue, Chikungunya, and Zika.
  • Polymerase Chain Reaction (PCR): Direct detection of viral RNA.
  • NS1 Antigen Test for Dengue: Early and specific detection of Dengue virus.
  • Urine Test for Zika: Viral presence in urine.
Appendix C: EVINCE Debate - Jaundice vs. Hepatitis
In this experiment, GPT-4 and Claude receive an identical list of symptoms from the moderator, aimed at guiding disease diagnosis, facilitating debate, and shaping subsequent recommendations. The “ground-truth” disease is Jaundice. (We do not shorten the debate content for review.)
C.1 Moderator’s Opening Prompt
“What potential disease could be associated with the following symptoms: itching, vomiting, fatigue, weight loss, high fever, yellowish skin, dark urine, and abdominal pain? Please offer top-five predictions with probabilities normalized to one, supported by justifications. Additionally, in the end of debate, include a list of supplementary symptom inquiries and recommend relevant lab tests to strengthen the confidence in your prediction.”
C.2 GPT-4 Opening Round
  • Hepatitis C (HCV): 40% Justification: Hepatitis C is a viral infection that can cause both acute and chronic hepatitis. The combination of symptoms like fatigue, abdominal pain, nausea, vomiting, jaundice, dark urine, and weight loss strongly suggests HCV. Itching may also occur due to liver dysfunction.
  • Hepatitis B (HBV): 30% Justification: Hepatitis B is a viral infection that can cause acute or chronic liver inflammation. Symptoms are similar to HCV, including fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, and itching. The presence of these symptoms warrants testing for both HBV and HCV.
  • Cirrhosis: 15% Justification: Cirrhosis is a late stage of liver scarring that can result from various liver diseases. The combination of symptoms like fatigue, weight loss, abdominal pain, jaundice, dark urine, and itching suggests advanced liver disease, possibly cirrhosis.
  • Obstructive Jaundice: 10% Justification: Obstructive jaundice occurs when there is a blockage in the bile ducts. The presence of jaundice, dark urine, itching, abdominal pain, nausea, vomiting, and weight loss suggests a possible obstruction. However, the absence of fever makes an infectious cause like cholangitis less likely.
  • Acute Liver Failure: 5% Justification: Acute liver failure is a rare but life-threatening condition characterized by rapid deterioration of liver function. The presence of jaundice, abdominal pain, nausea, vomiting, fatigue, and itching, along with the acute onset of symptoms, may indicate acute liver failure. However, the absence of altered mental status (hepatic encephalopathy) makes this less likely.
C.3 Claude Opening Round
  • Hepatitis B (HBV): 35% Reasoning: Hepatitis B is a viral infection that can cause acute or chronic liver inflammation. Symptoms include fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, joint pain, and itching.
  • Hepatitis C (HCV): 25% Reasoning: Hepatitis C is a viral infection that can cause both acute and chronic hepatitis. Symptoms are similar to HBV but may also include weight loss and fever. Jaundice is a common symptom.
  • Obstructive Jaundice (due to gallstones, pancreatic cancer, or bile duct obstruction): 20% Reasoning: Obstructive jaundice occurs when there is a blockage in the bile ducts, preventing bile from reaching the intestines. This can lead to yellowing of the skin and eyes, dark urine, itching, abdominal pain, nausea, and vomiting.
  • Alcoholic Hepatitis: 15% Reasoning: Alcoholic hepatitis is liver inflammation caused by excessive alcohol consumption. Symptoms include abdominal pain, nausea, vomiting, fatigue, dark urine, and possibly fever and weight loss.
  • Hepatitis A (HAV): 5% Reasoning: Hepatitis A is an acute, self-limiting liver infection transmitted through contaminated food or water. Symptoms include fatigue, abdominal pain, nausea, vomiting, dark urine, and jaundice.
C.4 GPT-4 Second Round
  • Hepatitis B (HBV): 35% Symptoms: fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice (yellowish skin), joint pain, itching
  • Hepatitis C (HCV): 25% Symptoms: fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice (yellowish skin), weight loss, fever, itching
  • Obstructive Jaundice (due to gallstones, pancreatic cancer, or bile duct obstruction): 20% Symptoms: yellowish skin (jaundice), dark urine, itching, abdominal pain, nausea, et al.
  • Alcoholic Hepatitis: 15% Symptoms: abdominal pain, nausea, vomiting, fatigue, jaundice (yellowish skin), dark urine, fever, weight loss
  • Hepatitis A (HAV): 5% Symptoms: fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice (yellowish skin)
C.5 Claude Second Round
  • Hepatitis C (HCV): 40% Reasoning: Hepatitis C is a viral infection that can cause both acute and chronic hepatitis. The combination of symptoms like fatigue, abdominal pain, nausea, vomiting, jaundice, dark urine, and weight loss strongly suggests HCV. Itching may also occur due to liver dysfunction.
  • Hepatitis B (HBV): 30% Reasoning: Hepatitis B is a viral infection that can cause acute or chronic liver inflammation. Symptoms are similar to HCV, including fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, and itching. The presence of these symptoms warrants testing for both HBV and HCV.
  • Cirrhosis (due to various causes such as chronic hepatitis, alcohol abuse, or non-alcoholic fatty liver disease): 15% Reasoning: Cirrhosis is a late stage of liver scarring that can result from various liver diseases. The combination of symptoms like fatigue, weight loss, abdominal pain, jaundice, dark urine, and itching suggests advanced liver disease, possibly cirrhosis.
  • Obstructive Jaundice (due to gallstones, pancreatic cancer, or bile duct obstruction): 10% Reasoning: Obstructive jaundice occurs when there is a blockage in the bile ducts. The presence of jaundice, dark urine, itching, abdominal pain, nausea, vomiting, and weight loss suggests a possible obstruction. However, the absence of fever makes an infectious cause like cholangitis less likely.
  • Acute Liver Failure (due to various causes such as viral hepatitis, drug-induced liver injury, or autoimmune hepatitis): 5% Reasoning: Acute liver failure is a rare but life-threatening condition characterized by rapid deterioration of liver function. The presence of jaundice, abdominal pain, nausea, vomiting, fatigue, and itching, along with the acute onset of symptoms, may indicate acute liver failure. However, the absence of altered mental status (hepatic encephalopathy) makes this less likely.
C.6 GPT-4 Third Round
  • Hepatitis C (HCV): 35% Reasoning: Hepatitis C is indeed a major cause of the symptoms listed, including fatigue, abdominal pain, nausea, vomiting, jaundice, dark urine, and weight loss. Itching can also occur due to liver dysfunction.
  • Hepatitis B (HBV): 25% Reasoning: Hepatitis B shares many symptoms with Hepatitis C, such as fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, and itching. Testing for both HBV and HCV is warranted given the overlap in symptoms.
  • Cirrhosis: 20% Reasoning: Cirrhosis can result from chronic hepatitis, alcohol abuse, or non-alcoholic fatty liver disease. The symptoms of fatigue, weight loss, abdominal pain, jaundice, dark urine, and itching strongly suggest advanced liver disease, which could be cirrhosis.
  • Obstructive Jaundice: 15% Reasoning: Obstructive jaundice, caused by gallstones, pancreatic cancer, or bile duct obstruction, is consistent with jaundice, dark urine, itching, abdominal pain, nausea, vomiting, and weight loss. The absence of fever reduces the likelihood of an infectious cause like cholangitis.
  • Acute Liver Failure: 5% Reasoning: Acute liver failure is less common but serious, characterized by rapid deterioration of liver function. The symptoms of jaundice, abdominal pain, nausea, vomiting, fatigue, and itching, along with the acute onset of symptoms, suggest this possibility. However, the absence of altered mental status (hepatic encephalopathy) makes it less likely.
C.7 Claude Third Round
  • Hepatitis C (HCV): 35% Reasoning: I slightly lowered the probability of HCV from 40% to 35%. While the combination of symptoms strongly suggests HCV, the absence of a few key risk factors (e.g., history of intravenous drug use or blood transfusions) in the given information makes me less confident in assigning a higher probability.
  • Hepatitis B (HBV): 30% Reasoning: I agree with the 30% probability assigned to HBV. The symptoms are indeed similar to HCV, and testing for both viruses is warranted.
  • Cirrhosis: 20% Reasoning: I increased the probability of cirrhosis from 15% to 20%. The presence of jaundice, dark urine, itching, and weight loss, along with the other symptoms, strongly suggests advanced liver disease. Cirrhosis can be a consequence of chronic liver diseases like hepatitis B or C, which have been assigned higher probabilities.
  • Obstructive Jaundice: 10% Reasoning: I agree with the 10% probability for obstructive jaundice. The absence of fever makes an infectious cause like cholangitis less likely, but the other symptoms still suggest a possible obstruction.
  • Acute Liver Failure: 5% Reasoning: I agree with the 5% probability for acute liver failure. While the combination of symptoms could indicate acute liver failure, the absence of altered mental status (hepatic encephalopathy) makes this diagnosis less likely.
C.8 Joint Final Round
  1. 1.
    Hepatitis C (HCV) - 35% Justification: The combination of symptoms strongly suggests HCV, but without additional risk factors, a slightly lower probability is appropriate.
  2. 2.
    Hepatitis B (HBV) - 30% Justification: The symptoms are similar to HCV, and testing for both viruses is necessary.
  3. 3.
    Cirrhosis - 20% Justification: The presence of jaundice, dark urine, itching, weight loss, and other symptoms indicates advanced liver disease, making cirrhosis a more likely possibility.
  4. 4.
    Obstructive Jaundice - 10% Justification: The absence of fever reduces the likelihood of an infectious cause like cholangitis, but the other symptoms still suggest a potential biliary obstruction.
  5. 5.
    Acute Liver Failure - 5% Justification: While the combination of symptoms could indicate acute liver failure, the absence of altered mental status makes this diagnosis less probable.
C.9 Final Joint Recommendations to GP
Gemini and GPT-4 jointly recommend the following additional symptom inquiries and lab tests to confirm a hepatitis diagnosis:
Additional Symptom Inquiries:
  1. 1.
    Onset and Duration: Determining the onset and duration of symptoms can clarify the illness timeline and its progression.
  2. 2.
    Severity and Trend: Evaluating symptom severity and whether they are worsening or improving aids in understanding the disease’s trajectory and treatment responses.
  3. 3.
    Associated Symptoms: Checking for additional symptoms like nausea, vomiting, fever, joint pain, or urine color changes can pinpoint other hepatitis indicators and exclude other conditions.
Recommended Lab Tests:
  1. 1.
    Liver Function Tests (LFTs): Critical for assessing liver health, LFTs evaluate enzyme levels such as alanine aminotransferase (ALT) and aspartate aminotransferase (AST), where abnormalities can signify liver inflammation.
  2. 2.
    Hepatitis Panel: This test checks for hepatitis A, B, and C viruses, vital for determining the specific type and guiding treatment strategies.
  3. 3.
    Serology Testing: Useful for distinguishing between acute and chronic hepatitis by identifying specific antibodies or antigens.
  4. 4.
    Imaging Studies: Ultrasound or MRI can provide visual insights into the liver’s state, detecting inflammation, fibrosis, or cirrhosis, thereby complementing blood-based diagnostics.
Appendix D: Formulas of Metrics
This appendix outlines the mathematical formulas for various data analysis metrics used in probabilistic and statistical modeling.
Kullback-Leibler Divergence
The Kullback-Leibler Divergence measures the difference between two probability distributions:
$D_{KL}(P\|Q)=\sum_{x\in\mathcal{X}}P(x)\log\left(\frac{P(x)}{Q(x)}\right).$
Jensen-Shannon Divergence
The Jensen-Shannon Divergence is a symmetrized and smoothed version of the KL Divergence:
$JSD(P\|Q)=\frac{1}{2}D_{KL}(P\|M)+\frac{1}{2}D_{KL}(Q\|M)$
where $M=\frac{1}{2}(P+Q)$.
Wasserstein Distance
The Wasserstein Distance, also known as the Earth Mover’s Distance (EMD), measures the distance between two probability distributions:
[…]
Cross Entropy
Cross Entropy measures the average number of bits required to identify an event from a set of possibilities, under a specific model:
$H(P,Q)=-\sum_{x\in\mathcal{X}}P(x)\log(Q(x)).$
Mutual Information
Mutual Information measures the amount of information that one random variable contains about another random variable:
$I(X;Y)=\sum_{y\in\mathcal{Y}}\sum_{x\in\mathcal{X}}p(x,y)\log\left(\frac{p(x,y)}{p(x)p(y)}\right).$
Normalized Mutual Information
Normalized Mutual Information is calculated as the mutual information divided by the maximum of the entropies of the variables:
$NMI(X;Y)=\frac{I(X;Y)}{\max(H(X),H(Y))}.$
Appendix E: Symptom Checking
This is the typical procedure of a GP to perform patient symptom checking.
  1. 1.
    Patient History: The GP begins by reviewing the patient’s medical history, including previous illnesses, chronic conditions, medications, allergies, and family medical history.
  2. 2.
    Symptom Assessment: The patient describes their current symptoms, including starting time and severity. This is an interactive process as the GP queries the patient for additional symptoms to their reported ones to disambiguate several possibilities.
  3. 3.
    Physical Examination: The GP performs simple physical exams, which may include checking vital signs (e.g., blood pressure, heart rate, temperature), examining specific body parts or systems, and assessing overall physical health.
  4. 4.
    Suggest Lab Tests: Depending on the symptoms and physical examination findings, the doctor may order diagnostic tests such as blood tests, X-rays, ultrasound, or other studies. These tests can help confirm or rule out certain conditions.
  5. 5.
    Diagnosis: Based on the information gathered, the doctor formulates a preliminary diagnosis.
  6. 6.
    Treatment or Management Plan: Once a diagnosis is established, the doctor develops a treatment or management plan. This may include prescribing medications, recommending lifestyle changes, suggesting physical therapy, or providing guidance on managing chronic conditions.
  7. 7.
    Referrals: In some cases, the GP may refer the patient to specialists for further evaluation and treatment. Specialists have expertise in specific areas of medicine, such as cardiology, orthopedics, or dermatology.
Appendix F: The EnToPPS Framework
EnToPPS integrates predictions from two LLMs, denoted as A and B, each providing probability distributions over C classes. The following steps outline the EnToPPS process:
  1. 1.
    Obtain Top-C Predictions: For each LLM (A and B), obtain the predicted probabilities for all C classes, denoted as $P_{A}$ and $P_{B}$:
    $P_{A}=[p_{A1},p_{A2},\ldots,p_{AC}],\quad P_{B}=[p_{B1},p_{B2},\ldots,p_{BC}],$
    where $p_{Ai}$ and $p_{Bi}$ represent the predicted probability of class $i$ by LLM A and B, respectively.
  2. 2.
    Select Top-k Predictions: For each LLM (A and B), select the top-k predicted classes based on their probabilities:
    $T_{A}=[t_{A1},t_{A2},\ldots,t_{Ak}],\quad T_{B}=[t_{B1},t_{B2},\ldots,t_{Bk}],$
    where $t_{Ai}$ and $t_{Bi}$ represent the class index of the $i^{th}$ top prediction by A and B, respectively.
  3. 3.
    Combine Top-k Predictions: Combine the top-k predictions from both LLMs to create a set of unique predicted classes:
    $T_{C}=T_{A}\cup T_{B}=[t_{C1},t_{C2},\ldots,t_{Cm}],k\leq m\leq 2k.$
  4. 4.
    Backfill Missing Probabilities: For each class in the combined set $T_{C}$, backfill its probability from the original probability distributions $P_{A}$ and $P_{B}$:
    • If a class $t_{Ci}$ is present in $T_{A}$, assign its probability from $P_{A}$: $p_{Ci}=p_{Ai}$.
    • If a class $t_{Ci}$ is present in $T_{B}$, assign its probability from $P_{B}$: $p_{Ci}=p_{Bi}$.
    • If a class $t_{Ci}$ is present in both $T_{A}$ and $T_{B}$, assign the average probability: $p_{Ci}=\frac{p_{Ai}+p_{Bi}}{2}$.
  5. 5.
    Normalize Probabilities: Normalize the probabilities of the classes in the combined set $T_{C}$ to ensure they sum up to 1:
    $P_{C}=[p_{C1},p_{C2},\ldots,p_{Cm}],\quad\text{where }p_{Ci}=\frac{p_{Ci}}{\sum_{j=1}^{m}p_{Cj}}.$

### v4

@@
-EVINCE: Optimizing Adversarial LLM Dialogues via Conditional Statistics and Information Theory
+EVINCE: Optimizing Multi-LLM Dialogues Using Conditional Statistics and Information Theory
 Abstract
-This paper introduces EVINCE (Entropy and Variation IN Conditional Exchanges), a dialogue framework advancing Artificial General Intelligence (AGI) by enhancing versatility, adaptivity, and reasoning in large language models (LLMs).
-Leveraging adversarial debate and a novel dual entropy theory, EVINCE improves prediction accuracy, robustness, and stability in LLMs by integrating statistical modeling, information theory, and machine learning to balance diverse perspective exploration with strong prior exploitation.
-The framework’s effectiveness is demonstrated through consistent convergence of information-theoretic metrics, particularly improved mutual information, fostering productive LLM collaboration.
-We apply EVINCE to healthcare, showing improved disease diagnosis, and discuss its broader implications for decision-making across domains.
-This work provides theoretical foundations and empirical validation for EVINCE, paving the way for advancements in LLM collaboration and AGI development.
-EVINCE: Optimizing Adversarial LLM Dialogues
-via Conditional Statistics and Information Theory
-Edward Y. Chang Computer Science Clinical Mind AI Stanford University echang@cs.stanford.edu Draft: 2/15/2024; Revised: 8/15/2024
+$\mathsf{EVINCE}$ (Entropy and Variation IN Conditional Exchanges) is a novel framework for optimizing multi-LLM dialogues using conditional statistics and information theory.
+It addresses limitations in multi-agent debate (MAS) frameworks, where multiple LLMs “chat” without behavior modulation or mutual information quality assessment.
+Using dual entropy optimization to balance perspective diversity and prior knowledge, $\mathsf{EVINCE}$ provides quantitative tools to dynamically regulate LLM linguistic behaviors.
+When mutual information is low and both cross-entropy and Wasserstein distance are high, $\mathsf{EVINCE}$ promotes contentious dialogues to expose diverse perspectives and uncover inconsistencies.
+Conversely, as cross-entropy decreases and mutual information stabilizes, it transitions discussions into a conciliatory phase, encouraging compromise and acknowledgment of valid points.
+Using information-theoretic metrics and optimizing mutual information, $\mathsf{EVINCE}$ emerges as a structured and highly effective framework for multi-LLM collaboration.
+Machine Learning, Multiple Agents
 1 Introduction
-The pursuit of Artificial General Intelligence (AGI) remains a central goal of AI research.
-We propose a paradigm shift in this quest: utilizing multiple Large Language Models (LLMs) engaged in synergistic dialogues as a crucial step towards AGI.
-This approach, we contend, addresses key limitations of current AI systems and provides a novel pathway to more robust, versatile, and capable artificial intelligence.
-Specifically, our work targets three core AGI characteristics: versatility, adaptivity, and reasoning capability.
-Current LLMs, despite their remarkable capabilities, face significant challenges, including hallucination (generating false or nonsensical information), bias (reflecting and potentially amplifying societal prejudices), and limited reasoning (difficulties in complex problem-solving and logical inference).
-We posit that multi-agent dialogue systems offer a promising avenue to address these challenges.
-By fostering diversity and debate among LLMs, these systems can mitigate biases and promote enhanced reasoning capabilities.
-Furthermore, the iterative nature of multi-round dialogues allows for continuous context enrichment, enabling LLMs to access more precise information and formulate more accurate responses, thus reducing the occurrence of hallucinations.
-Our previous work, SocraSynth Chang (2023a), addresses LLM limitations through structured multi-agent dialogues.
-By leveraging both adversarial and collaborative interactions between LLMs, SocraSynth demonstrates quantifiable improvements across various domains.
-In healthcare, it achieves a 5% improvement in diagnostic accuracy Chang and et al.
-(2023), while in news analysis, it identifies and explains biases in articles Chang (2024c).
-These advancements showcase enhanced versatility, adaptivity, and reasoning capability—key AGI characteristics.
-SocraSynth’s adversarial component promotes the exploration of diverse perspectives, while its collaborative component fosters rigorous reasoning to reach well-reasoned conclusions.
-This synergy has yielded measurable gains beyond healthcare and bias mitigation, extending to geopolitical analysis Chang (2023b), corporate planning Tsao (2023), investment banking Chang (2024b), and emotional behavior modeling Chang (2024a).
-These results demonstrate SocraSynth’s effectiveness in mitigating LLM limitations and achieving substantial performance improvements across various applications, highlighting its potential for advancing towards AGI’s generalized problem-solving capabilities.
-Although SocraSynth has demonstrated effectiveness, it relies on “contentiousness” as a qualitative measure to moderate the linguistic behaviors of participating LLMs.
-For instance, a high contentiousness value (0.9 out of 1.0) might lead LLMs to challenge each other’s assumptions and propose alternative solutions, while a low value (below 0.3) could encourage them to synthesize their viewpoints and find common ground.
-While the concept of contentiousness has proven useful in guiding SocraSynth dialogues, its qualitative nature limits its precision and explainability.
-In this paper, we extend SocraSynth by introducing three theoretical pillars to quantify “contentiousness” and moderate dialogues based on statistical and information theories.
-These pillars, collectively referred to as EVINCE (Entropy and Variation IN Conditional Exchanges), provide quantitative measures for justifiable and explainable multi-agent dialogue moderation and evaluation.
-These pillars are:
+Current large language models (LLMs) (e.g.
+(Anthropic, 2024; OpenAI, 2023; Gemini, 2023)) demonstrate remarkable capabilities but face significant challenges, including hallucination (generating false information), bias (reflecting societal prejudices), and limited reasoning (difficulties in complex problem solving and logical inference).
+These challenges critically limit LLMs’ reliability and applicability in real-world scenarios.
+Multi-agent dialogue systems present a promising approach to addressing these limitations.
+By fostering diversity and structured debate among LLMs, these systems can mitigate biases and enhance reasoning through collaborative intelligence.
+However, most prior works in Multi-Agent Debate (MAD) (e.g., (Abdelnabi et al., 2024; Chan et al., 2023; Fu et al., 2023; Li et al., 2023; Liang et al., 2023; Michael et al., 2023; Smit et al., 2024)) primarily follow the traditional Mixture of Experts (MoE) paradigm (Jacobs et al., 1991).
+These methods rely on redundancy to mask errors (Freund & Schapire, 1997) and leverage error diversity to improve response quality.
+As a result, their debates often suffer from repetitive arguments, reinforcement of flawed reasoning, and inconsistent conclusions (Wang et al., 2024; Zhang et al., 2024).
+Moreover, these systems lack dynamic modulation of contentious intensity during debates, leading to discussions that are either uniformly mild or overly contentious, ultimately producing suboptimal results.
+Recent work by SocraSynth (Chang, 2023a) explored the dynamic modulation of debate “contentiousness” to generate diverse perspectives and then refine with reasoning to improve results.
+However, its qualitative approach to modulating contentiousness in linguistic behavior limits precise control and optimization of these interactions.
+In this paper, we advance MAD by introducing $\mathsf{EVINCE}$ (Entropy and Variation IN Conditional Exchanges), which provides quantitative foundations for multi-agent dialogue moderation.
+$\mathsf{EVINCE}$ builds on three theoretical pillars:
   1. 1.
-    Inclusiveness Exploration: We develop methods to ensure dialogues explore all potential perspectives.
-We use conditional statistics via in-context learning Xie et al.
-(2021) to “free” an LLM agent from its default “maximal likelihood” next-token prediction behavior, allowing it to adopt specific stances.
-We introduce a dual entropy optimality theory to balance the exploration of new ideas with adherence to priors, thus optimizing information exchange between agents for comprehensive and stable discourse.
+    Inclusive Exploration: We develop methods to ensure dialogues comprehensively explore diverse perspectives.
+Using conditional statistics, we adjust an LLM agent’s behavior beyond its default “maximum likelihood” next-token prediction, enabling it to adopt specific stances via in-context learning.
+To balance idea exploration with adherence to prior knowledge, we introduce a dual-entropy optimization framework, improving information exchange for richer discourse.
   2. 2.
-    Information Flow Dynamics: We introduce information theory-based metrics to quantify and optimize dialogue dynamics.
-These measure information diversity (entropy), novelty (statistical divergence scores), and inter-agent persuasion (mutual information).
-These metrics enable us to assess and enhance the quality and efficiency of information flow within the multi-agent system, fostering rich and productive exchanges.
+    Information Flow Dynamics: We quantify and optimize dialogue interactions using information-theoretic metrics, measuring information diversity (entropy), novelty (divergence scores), and inter-agent persuasion (mutual information).
+These metrics enhance the quality and efficiency of information flow in multi-agent settings, fostering richer, more productive exchanges.
   3. 3.
-    Reasoning Quality and Coherence: We establish frameworks to assess the logical structure and coherence of multi-agent reasoning.
-This pillar evaluates argument validity, analytical depth, and dialogue coherence.
-We synergistically integrate our previously developed CRIT system Chang (2023c) into SocraSynth.
-CRIT, which combines Socratic methods with formal reasoning techniques, enhances our ability to conduct critical thinking through evaluating argument quality, information-source credibility, and overall “reasonableness” within the dialogue.
-This integration ensures that the collective reasoning of LLM agents is not only diverse but also logically sound and aligned with the dialogue’s objectives.
-The core strength of SocraSynth and EVINCE in advancing towards AGI lies in their ability to enhance key AGI characteristics through multi-agent dialogues.
-By employing conditional statistics and information theory, they boost versatility and adaptivity, allowing LLMs to transcend their typical “maximal likelihood” behaviors and mimic how humans adapt their linguistic behaviors to complete tasks.
-The framework’s debate structure fosters advanced reasoning and critical thinking capabilities.
-Together, these advancements represent significant steps towards achieving the complex, intricate capabilities required for AGI.
-The main contributions of this paper are:
+    Reasoning Quality and Coherence: We establish frameworks to evaluate the logical structure and coherence of multi-agent reasoning.
+This includes assessing argument validity, analytical depth, and dialogue consistency.
+We integrate the CRIT algorithm (Chang, 2023b), which combines Socratic methods with formal reasoning, to enhance argument evaluation and ensure logically sound, goal-oriented discourse.
+The contributions of this paper are as follows:
   1. 1.
-    EVINCE Framework Design: Unlike approaches that use debate merely to improve accuracy via redundancy, EVINCE facilitates information discovery, bias mitigation, and decision-making that requires both breadth and depth of information.
+    Framework Design: Unlike prior debate-based methods that enhance accuracy primarily through redundancy, $\mathsf{EVINCE}$ enables structured information discovery, bias mitigation, and decision-making by integrating both broad exploratory analysis and deep evaluative reasoning.
   2. 2.
-    Theoretical Foundations: We establish a theoretical basis for EVINCE, rooted in conditional Bayesian statistics, mutual information, and dual entropy.
-These principles are applied to measure, monitor, and modulate collaborative LLM interactions, contributing to a deeper understanding of how LLMs can effectively cooperate for improved decision-making.
-The dual entropy theory is novel and ground-breaking, illustrating how a productive decision-making process should start with room for diverse input and stable objectives, and then, through information exchange, converge to optimal decision/prediction.
+    Theoretical Foundations: $\mathsf{EVINCE}$ establishes a theoretical foundation rooted in Bayesian statistics, mutual information, and dual entropy.
+The dual entropy framework formalizes how decision-making should begin with diverse, high-entropy inputs and stable objectives, then iteratively refine through information exchange to achieve optimal predictions.
   3. 3.
-    Empirical Validation: We provide empirical validation of EVINCE’s underlying theories, highlighting the framework’s effectiveness in balancing exploration and exploitation to enhance prediction accuracy.
-We also introduce a set of maxims derived from our empirical findings, offering practical guidance for optimizing mutual information and minimizing various divergence measures.
-The remainder of this paper is structured as follows: Section 2 discusses related work and addresses common misconceptions about our approach.
-Section 3 presents the three pillars of EVINCE, including its maxims, theories, and algorithms.
-Section 4 provides empirical evidence of EVINCE’s benefits by revisiting and expanding upon previous SocraSynth experiments in healthcare and bias detection.
-Section 5 offers concluding remarks and discusses future directions.
-2 Related Work
-The core objective of adversarial debate, as embodied in SocraSynth with EVINCE, is to foster diverse opinions and challenge assumptions, ultimately leading to more comprehensive and informed decision-making.
-This contrasts with traditional ensemble learning methods, which prioritize error diversity for improved accuracy.
-This section surveys ensemble learning and adversarial learning, highlighting the unique aspects of EVINCE and its theoretical underpinnings.
-2.1 Ensemble Learning
-Ensemble methods, such as Bagging Breiman (1996), Boosting Freund and Schapire (1997), and Mixtures of Experts Jacobs et al.
-(1991), traditionally aim to enhance machine learning performance by combining predictions from multiple models to achieve error diversity.
-Similarly, early LLM debate frameworks, e.g., Michael et al.
-(2023); Chan et al.
-(2023); Liang et al.
-(2023); Du et al.
-(2023) have focused primarily on improving answer accuracy rather than fostering a comprehensive exploration of information for critical decision-making.
-SocraSynth distinguishes itself from traditional ensemble methods by prioritizing the generation of diverse predictions over the mere avoidance of errors.
-This is achieved through a dynamic protocol that adaptively adjusts the “contentiousness” level of the debate, encouraging models to initially explore a wide range of perspectives and rigorously assess the quality of arguments.
-Through constructive exchange of ideas in subsequent iterations, SocraSynth converges on a more informed and robust decision.
-Its approach, grounded in EVINCE’s dual entropy theory, allows for the exploration of various ideas by one LLM while maintaining stability through the other, effectively preventing premature convergence on a narrow set of perspectives.
-2.2 Addressing Common Misconceptions about Adversarial LLM Dialogue
-The rapid growth in academic conference submissions has led to an expanded reviewer pool, potentially impacting review quality.
-In this context, some misconceptions about our work have arisen, which we address here:
+    Empirical Validation: Through experiments in disease diagnosis and news debiasing (Section 3), we validate $\mathsf{EVINCE}$’s principles and demonstrate how optimizing information flow improves predictive accuracy and reasoning robustness.
+2 EVINCE Algorithm, Maxims, and Theories
+Problem Statement:
+Organize a structured debate between two equally competent large language models (LLMs), LLMA and LLMB, to conduct $t$ rounds.
+At each round $t$, each model generates confidence scores for its top-k predictions, denoted as $P_{A}^{(t)}$ and $P_{B}^{(t)}$, over $C$ possible outcomes, accompanied by supporting arguments $R_{A}^{(t)}$ and $R_{B}^{(t)}$.
+The goal is to design an iterative debate process that leverages the structured exchange of arguments to enable the models to converge on an optimal prediction ranking $P^{*}$ across the $C$ classes.
+Optimality is defined as achieving the highest possible accuracy on the prediction task while maintaining strong reasoning support.
+2.1 Preliminaries in Information Theory
+This section summarizes the key metrics used in $\mathsf{EVINCE}$ to evaluate information diversity, similarity, divergence, and other relevant factors.
+A detailed comparison, including advantages and limitations, is provided in Appendix A.
+Notably, computational complexity is not a concern in our context, as we work with probability distributions of up to ten categories.
+These metrics serve three primary objectives:
+A. Fostering diversity of perspectives while ensuring reasoning quality:.
+    Wasserstein Distance (WD): Measures the difference between prediction distributions to identify exploration opportunities (Kantorovich, 1942; Rubner et al., 2000; Villani, 2008).
+    Shannon Entropy or Relative Entropy: Measures diversity of perspectives (Shannon, 1948).
+    Reasoning Quality: The $\mathsf{CRIT}$ (Critical Thinking) algorithm evaluates the logical soundness and persuasiveness of supporting arguments (depicted in Appendix B), helping to identify and mitigate hallucinations and poorly-reasoned arguments (Chang, 2023b).
+B. Exploring new possibilities while maintaining stability toward the goal of the dialogue:.
+    Correlation Coefficients: Tracks the evolution of opinions and assesses debate stability (Brown et al., 2005) toward the goal of the dialogue.
+    Mutual Information (MI): Quantifies information overlap to ensure focused and productive debates (Cover & Thomas, 2006b) and measures the degree of agreement/disagreement.
+C. Examining information convergence and establishing dialogue termination criteria:.
+    Jensen-Shannon (JS) Divergence: Assesses similarity between probability distributions (symmetric).
+    Cross Entropy (CE): Measures the asymmetric difference between prediction distributions.
+    Kullback-Leibler (KL) Divergence: Reveals asymmetric differences between probability distributions.
+Next, we present the $\mathsf{EVINCE}$ algorithm and explain how these metrics are used to moderate an LLM dialogue to achieve a balance between exploration and exploitation, leading to optimal prediction outcomes.
+2.2 Algorithm Specifications
+Figure 1 specifies the detailed algorithm of $\mathsf{EVINCE}$.
+It employs two equally competent LLM instances, $LLM_{A}$ and $LLM_{B}$, which can be different models (e.g., GPT and Claude) or independent instances of the same model.
+Given an information set $S$ and a class-label set $C$, $\mathsf{EVINCE}$ outputs a confidence distribution over $C$ with justifications.
+For example, $S$ could represent a patient’s symptoms and $C$ a set of diseases, with $\mathsf{EVINCE}$ moderating a dialogue to predict the patient’s disease(s).
+[…]
+Figure 1: Specifications of Algorithm $\mathsf{EVINCE}$.
+Key points: 1) Asymmetric Start: In Step #1, LLMA initiates with opening arguments based solely on the given information, while LLMB starts with access to LLMA’s prediction and arguments for refutation.
+2) Termination Criteria: The while loop in Step #2 considers three factors: Wasserstein distance, mutual information, and argument quality.
+$\mathsf{EVINCE}$ terminates if the dialogue ceases to make significant progress.
+3) Further Details: Maxims #1 to #4 provide additional explanations of the algorithm’s principles.
+4) Argument Evaluation: Step #2.2 evaluates of argument quality, and the while loop examines if argument quality continues to improve.
+5) Update($\Delta$) modulates contentiousness, and see Maxim #3 for its specifications.
+Moderation Subroutines:
+$\mathsf{EVINCE}$ employs four key subroutines from the previous subsection to manage the exploration-exploitation tradeoff, ensuring information diversity, quality, and stability:
+  • $\mathsf{CRIT}$: Evaluates argument quality and source credibility. Low scores can trigger dialogue termination.
+  • WD (Wasserstein distance): Assesses prediction diversity, expected to decrease over time.
+  • MI (Mutual Information): Evaluates dialogue convergence. Stagnation can trigger termination.
+  • Additional metrics: KL Jensen-Shannon divergence, and cross entropy ensure evaluation consistency.
+When all metrics plateau, indicating no further improvement, $\mathsf{EVINCE}$ terminates the dialogue.
+(See the while loop in the second step of the algorithm).) At this stage, both LLM instances are prompted to collaboratively deliver a conciliatory conclusion that includes comprehensive arguments and counterarguments.
+They also provide a list of missing information that, if obtained, could enhance prediction accuracy and reliability, potentially using Retrieval-Augmented Generation (RAG) techniques.
+Dialogue Iterations (Steps 1 and 2):
+Given the contentiousness level set for a dialogue iteration, each LLM generates a new prediction distribution on $C$ with supporting arguments, considering the other LLM’s previous prediction and accumulated reasoning as context.
+The contentiousness level updates follow Maxims #2.2 and #3.
+The top-k predictions from LLMs are then calibrated using reasoning quality assessment following Maxim #4.
+Final Output (Step 3):
+  • Combine final predictions from both LLMs.
+  • Weight predictions based on supporting argument quality (evaluated by $\mathsf{CRIT}$).
+  • Handle uncertainty by soliciting missing information when final prediction entropy is high (Maxim #4.4).
+$\mathsf{EVINCE}$ facilitates a structured debate between AI models, unearthing all perspectives related to the prediction tasks at hand while balancing diverse viewpoints and fostering consensus to produce accurate, well-reasoned predictions.
+2.3 Maxims with Theoretical Foundations
+Progress towards the optimality goal is guided and measured by metrics introduced in Section 2.1.
+This section explains how these metrics can be used in complementary ways to facilitate proper trade-offs between diversity and convergence, exploration and exploitation, and several other factors.
+In the $\mathsf{EVINCE}$ algorithm presented in Figure 1, we have annotated the steps to which these maxims are applied.
+Maxim #1. Orchestrate Two Equally Competent LLMs in Structured Debate:
+Integrating two equally competent LLMs ensures a balanced exchange of insights and avoids bias.
+This adversarial setup fosters diversity in predictions, each supported by justifications, promoting critical evaluation and uncovering potential blind spots.
+Methods: Choosing LLMs with comparable performance on a shared validation set, a balanced debate can be ensured.
+Suitable models (in 2024) include GPT-4, Claude, and Gemini.
+Conditioning different instances of the same LLM to support opposing stances on a subject matter can also be effective due to the theoretical justification of in-context learning with conditional Bayesian statistics (Xie et al., 2021).
+Maxim #2. Fostering Exploration through Diverse Perspectives:
+While LLMs default to maximum likelihood predictions favoring popular outcomes, context conditioning can shift them toward exploration over exploitation.
+High initial contentiousness encourages dynamic debate and challenges to prevailing views, mitigating confirmation bias through contrary queries and diverse top-k predictions.
+Methods: Set high initial contentiousness to stimulate exploration, enabling LLMs to challenge each other’s predictions. Monitor mutual information and agreement diversity to ensure productive debate.
+Maxim #3. Refining High-Quality Perspectives:
+Once new insights plateau (measured by mutual information, Wasserstein distance, and KL divergence), shift from exploration to exploitation by reducing contentiousness.
+Methods: Algorithm Update($\Delta$) guides this transition to a conciliatory stage, focusing on strengthening well-supported arguments for the final output.
+Maxim #4. Combine Predictions Weighted by Diversity and Quality:
+Combine the probability distributions from two LLMs by weighting them based on distributional diversity and argument quality.
+While LLMs face fundamental challenges in generating accurate probabilities ((Peng & Zhang, 2024; Scholten et al., 2024)), we can use reasoning-calibrated confidence scores as an effective proxy for weighting predictions.
+Methods: Following these four maxims:
+  • Maxim #4.1 Prediction Reliability: Use entropy-based measures to estimate reliability. Lower entropy suggests higher confidence and greater reliability in predictions.
+  • Maxim #4.2 Argument Quality: Evaluate argument quality using $\mathsf{CRIT}$, identifying logical fallacies and assessing the relevance and credibility of evidence.
+  • Maxim #4.3 Aggregation: Apply a weighted aggregation method, such as a Bayesian model, to combine predictions, accounting for both probabilistic insights and argument quality.
+  • Maxim #4.4 Diagnosis and RAG: If the final prediction has high entropy, indicating uncertainty, diagnose the issue and perform Retrieval-Augmented Generation (RAG) to obtain missing information.
+2.4 Entropy Duality Theorem (EDT)
+Theorem EDT: Optimal Pairing of LLMs for Probabilistic Prediction Accuracy.
+The optimal pairing of LLMs for diagnosis accuracy, in terms of stability, accuracy, and robustness, occurs when the LLMs are 1) equivalent in the quality of the information they process, and 2) exhibit contrasting entropy values in their prediction distributions—one high and one low.
+Proof.
+Please see Appendix C. ∎
+3 Experiments on Healthcare & News
+This study evaluates $\mathsf{EVINCE}$’ versatility by examining two distinct applications: 1.
+Disease Diagnosis, using LLMs as diagnostic tools, and 2.
+News Debiasing, which addresses bias in language generation.
+Together, these experiments demonstrate the framework’s capability to handle diverse and complex tasks.
   1. 1.
-    Misconception: Multiple LLMs cannot mitigate hallucination.
-    Some reviewers have expressed skepticism about the ability of multiple LLMs to reduce hallucination.
-However, our approach does not simply aggregate potentially flawed outputs.
-Instead, SocraSynth leverages the strengths of multiple models to cross-check, challenge, and refine information.
-This process is analogous to peer review in scientific discourse, where collective scrutiny enhances the reliability of conclusions.
+    Contentiousness & Convergence: The initial disagreement between the LLM (measured by the Wasserstein distance) reflects the diversity of perspectives, with disagreements decreasing as the debates progress and the LLMs reach consensus.
+Individual uncertainty (measured via Shannon entropy) follows a similar trend, starting high and reducing as LLMs refine their predictions.
   2. 2.
-    Misconception: Multiple LLM dialogues cannot lead to knowledge discovery.
-    Despite evidence provided in Chang (2023a) and Chang and Chang (2023), some reviewers remain skeptical about the potential for knowledge discovery through LLM dialogues.
-Our work demonstrates that carefully structured interactions between LLMs can indeed lead to novel insights and knowledge synthesis by leveraging the polydisciplinary knowledge representation inherent in LLMs.
+    EDT Effectiveness: LLM pairs optimized under the Entropy Duality Theorem (EDT) exhibit complementary error patterns.
+This synergy leads to improved overall prediction accuracy compared to non-EDT configurations, demonstrating $\mathsf{EVINCE}$’s ability to balance exploration and consensus.
   3. 3.
-    Misconception: The approach lacks theoretical foundation.
-    Our approach is grounded in conditional statistics and information theory, which we use to optimize LLM interactions.
-This theoretical foundation ensures that the dialogue converges towards more accurate and less biased outcomes, addressing concerns about the rigor of our methodology.
-By addressing these misconceptions, we aim to clarify the scientific basis and potential of our approach.
-The synergistic interaction between LLMs in SocraSynth, coupled with EVINCE’s theoretical pillars, represents a novel method for enhancing the capabilities of language models.
-This approach has applications in reducing bias, mitigating hallucination, and potentially advancing the field towards more robust AI systems.
+    Anomaly Identification: $\mathsf{EVINCE}$ enhances the detection of anomalies or inconsistencies in the generated predictions, uncovering ambiguities in input data or errors in existing labels in various applications.
+General Problem Statement:
+Given an input context $\kappa$ and a set of characteristics $F$, the objective is to generate and rank the top-$k$ predictions (e.g., diseases for the first problem and biased news for the second) from a set of possible outcomes $C$.
+This can be expressed as $P=\text{LLM}(F,\kappa)$, where each LLM outputs confidence scores for its top-$k$ predictions ($k\leq|C|$), leveraging the input features and context.
+$P=\left(p(\text{top},1,\text{to},k\in C\mid F,\kappa)\right).\vspace{-0.1in}$ (1)
+(a) GPT4 pairs Claude
+(b) GPT4 pairs Gemini
+Figure 2: EVINCE improves diagnosis accuracy markedly.
+Context $\kappa$ enables dual entropy adjustment via temperature, top-$k$, and contentiousness $\Delta$, with higher settings increasing entropy and lower settings reducing it.
+Note: Maxim #4 states that the initial confidence distribution $P$ from LLMs (Eq.1) represents confidence scores rather than absolute probabilities.
+In the $\mathsf{EVINCE}$ framework, these scores are calibrated through multi-iteration debates where predictions must be supported by arguments and examined by counterarguments.
+The $\mathsf{CRIT}$ algorithm (detailed in Appendix B) evaluates reasoning quality to adjust these confidence scores—reducing credibility when high-confidence predictions lack strong supporting evidence.
+Since $\mathsf{EVINCE}$ and $\mathsf{CRIT}$ can employ different LLMs, and $\mathsf{CRIT}$ focuses solely on evaluating reasoning quality via Socratic methods, this collaborative approach uses reasoning quality as a proxy for confidence calibration.
+This sidesteps the fundamental challenge of generating reliable absolute probabilities from LLMs ((Peng & Zhang, 2024; Scholten et al., 2024)).
+For example, if an LLM outputs confidence scores $P$ = (0.5, 0.3, 0.2) for its top-3 predicted diseases, and $\mathsf{CRIT}$ evaluates their supporting arguments as having strength (50%, 100%, 50%), the calibrated scores become (0.25, 0.3, 0.1).
+After normalization, the final distribution is $P$ = (0.38, 0.46, 0.16).
+(a) GPT liver c-matrix
+(b) Claude liver c-matrix
+Figure 3: Confusion matrices
+3.1 Experiment #1 Disease Diagnosis
+Resources & Dataset:
+We use a Kaggle dataset (Patil, 2020) of 4,921 patient records, each containing a diagnosis and up to 17 symptoms.
+After deduplication, 304 unique cases across 40 diseases remain (dataset provided as supplement).
+We evaluate GPT-4, Gemini, and Claude-3 using their inherent knowledge, without few-shot training.
+Computing resources provided by Azure research grant.
+Metric:
+We use top-k Mean Reciprocal Rank (MRR): if a top-k prediction matches ground truth, the score is the reciprocal of its rank (1 for first, 1/2 for second, etc.); 0 if no match.
+3.1.1 EVINCE vs. Baseline Accuracy
+We evaluated $\mathsf{EVINCE}$ through pairwise collaborations of GPT-4, Gemini, and Claude-3 on 304 patient cases.
+Parameter Settings:
+Each LLM pair uses $k=5$ predictions with contrasting temperatures (high/low) and high contentiousness ($\Delta=0.9$).
+This ensures meaningful interaction while maximizing cross entropy for information exchange.
+Evaluations:
+We conducted two types of experiments:
+1.
+Constrained Prediction (Baseline): Limiting predictions to the dataset’s 40 disease labels yielded high accuracy (95-97%).
+While unrealistic for general practice where physicians consider all possibilities, this constraint demonstrates LLMs’ flexibility in avoiding overfitting to potentially erroneous labels.
+2.
+Unconstrained Prediction: Removing label constraints to simulate real-world conditions, all 304 cases showed consistent results across LLMs (standard deviation = 1.5%).
+Individual pre-debate accuracies (Figure 2, light blue bars) were: GPT-4 (82.8%), Gemini (80.3%), and Claude (79.5%).
+(a) Entropy
+(b) WD %
+(c) Norm. MI
+Figure 4: Entropy, WD, and normalized MI
+Performance:
+$\mathsf{EVINCE}$ pairings (GPT-4/Claude-3 and GPT-4/Gemini-3) improved accuracy by 4-5 percentage points (Figure 2, green bars).
+The GPT-4/Claude-3 pairing achieved 87.5% accuracy (Figure 2(a)), comparable to state-of-the-art clinical algorithms (Peng et al., 2018).
+Discussion:
+While $\mathsf{EVINCE}$ significantly outperforms single LLM predictions, the remaining 12.5% inaccuracy merits deeper analysis.
+Given the reported 11% US misdiagnosis rate (Newman-Toker et al., 2023b), some discrepancies might reflect errors in the original dataset rather than $\mathsf{EVINCE}$ limitations.
+This observation suggests $\mathsf{EVINCE}$’s potential for identifying and correcting dataset errors, which we explore in Section 3.1.3.
+3.1.2 Analysis: Explorative vs. Exploitative
+(a) Study #2 convergence
+(b) Study #3 convergence
+Figure 5: Convergence of all information metrics.
+Analysis of Confusion Matrices:
+The confusion matrices for Hepatitis types A-E diagnosis (Fig. 3) reveal:
+  • GPT-4: Limited accuracy (50%) for types A/B, poor performance on C/D.
+  • Claude: Broader distribution of predictions of all types.
+Claude’s higher entropy predictions complement GPT-4’s more confident but potentially incorrect assessments.
+This differential entropy drives the productive debate in $\mathsf{EVINCE}$, balancing exploration with exploitation to improve diagnostic accuracy.
+Comparative Performance:
+EVINCE achieves 5% higher precision in the diagnosis of specific types of liver disease compared to a baseline approach (Figure 2(a)), highlighting its ability to handle complex diagnostic scenarios.
+Two primary factors contribute to the improved diagnostic accuracy of $\mathsf{EVINCE}$.
+First, structured debates with reasoning encourage LLMs to explore alternative diagnoses both broadly and deeply, leading to more comprehensive analysis and decision-making (refer to Appendices D and E for two debate examples).
+Second, pairing LLMs with high- and low-entropy prediction distributions, or those with a large Wasserstein distance (WD) between them, balances exploratory diversity with exploitative stability.
+This approach results in more robust and higher-quality decisions, as demonstrated in this second study.
+(a) GPT & Claude top-5
+(b) WD %
+(c) Mutual Info.
+Figure 6: Remediation: Jaundice to Hepatitis
+Observations from Information Metrics:
+  • Entropy Stabilization: Figure 4(a) shows the stabilization of the entropy levels of both LLMs after three debate rounds, indicating convergence towards a similar stable entropy state.
+  • Wasserstein Distance Improvement: Figure 4(b) shows a consistent improvement in the Wasserstein distance (WD) between the predictions of the two models over rounds.
+  • Mutual Information Increase: Figure 4(c) reveals a 14% improvement in normalized mutual information between GPT-4 and Claude’s prediction distributions, suggesting increased shared information throughout the debate.
+  • Convergence of Divergence Metrics: Figure 5(a) shows consistent convergence of all divergence metrics.
+3.1.3 Explainability and Remediation
+This study illustrates how $\mathsf{EVINCE}$ can identify potential misdiagnoses, explain the reasoning behind them, and recommend corrective actions.
+Traditionally, machine learning scientists rely on labeled data as “ground truth.” However, as evidenced by research like that of (Newman-Toker et al., 2023a) from Johns Hopkins, misdiagnosis is a widespread issue in healthcare systems globally.
+These erroneous diagnoses, often treated as ground truth, can be perpetuated by supervised learning algorithms, exacerbating the problem within the healthcare system.
+$\mathsf{EVINCE}$’s dialogue capabilities provide insights into the decision-making process and highlight missing information, helping to rectify erroneous predictions and redefine the ground truth.
+This approach offers a potential solution to the compounding effect of misdiagnoses in machine learning applications in healthcare.
+Figure 6, which plots the respective entropies of GPT-4 and Claude, reveals two key insights.
+First, an large gap in Wasserstein Distance (WD) exists between the two models in the initial rounds.
+This disparity underscores the role of dual entropy in fostering information exchange.
+As the entropy values converge in rounds 3 and 4 and WD decreases significantly, we observe a corresponding convergence and stabilization of their mutual information.
+The information metrics that EVINCE employs effectively show the progress of the dialogue and convergence from explorative to consensus.
+Figure 5(b) illustrates the convergence of all divergence metrics—including Jensen-Shannon divergence, cross-entropy, and Kullback-Leibler divergence—particularly between the second and third rounds.
+Appendix F demonstrates that EVINCE recommends additional symptoms to inquire about with the patient, as well as lab tests to confirm the diagnosis.
+These suggestions, validated by our hospital partners, provide valuable information to enhance diagnostic accuracy and correct errors.
+3.2 Experiment #2 News Debiasing
+This experimental framework aims to assess the feasibility of both detecting biases in textual content and implementing effective mitigation strategies.
+The first experiment focuses on bias detection, while the second explores the generation of balanced textual outputs as a corrective measure, moving beyond the limitations of prior studies that primarily focused only on identification.
+Resources & Dataset:
+This study utilizes a unique dataset of 619 news articles (54.3% about Democrat scandals, 45.7% about Republican scandals) selected from a larger 2013 repository of 14,033 articles compiled by fifteen reputable news organizations (Budak et al., 2016).
+These articles span diverse topics including civil rights, healthcare, elections, and national security, offering a comprehensive view of political coverage.
+Please visit (Anonymous, 2024) for links to the full set of news articles.
+The dataset’s distinctive feature is its ground-truth labels provided by annotators with declared political affiliations.
+Through Amazon Mechanical Turk, 749 qualified U.S.
+workers, each annotating up to 1,000 randomly selected articles, classified articles on a five-point scale from ‘negatively biased’ to ‘positively biased’ (Budak et al., 2016).
+Crucially, each scandal article in our subset received independent classifications from both Democrat and Republican annotators.
+In Appendix G.1, we justify the sufficiency of the annotation method.
+Metric:
+The bias gap, which varies from -2 to 2, serves as a metric to quantify the direction and degree of bias.
+As a baseline, we used Claude and GPT-4 to generate initial results.
+For $\mathsf{EVINCE}$ experiments, we employed two instances of GPT-4, as Claude demonstrated a tendency to shift its predictions easily (a point elaborated on later).
+3.2.1 Bias Evaluation Results
+Due to space constraints, detailed results and discussions on Democrat scandals are provided in Appendix G.2, while those on Republican scandals are in Appendix G.3.
+Below, we summarize the key findings.
+Figure 7 illustrates the distribution of ratings for all scandals across four scenarios:
+1) Democrat-leaning annotators rating Democrat scandals, 2) Republican-leaning annotators rating Democrat scandals, 3) Democrat-leaning annotators rating Republican scandals, and 4) Republican-leaning annotators rating Republican scandals.
+Figure 7: Bias Rating Distributions Show Strong Biases.
+D is more negative on how D scandals were reported (the sub-figure on the left), R is more negative on how R scandals were reported (the sub-figure on the right).
+The figure reveals a clear pattern: Democrat-leaning annotators tend to rate news about Democrat scandals more negatively, while Republican-leaning annotators exhibit similar negativity towards reports on Republican scandals.
+The gap between these ratings is approximately one class-label (e.g., between “weak negative” and “neutral”), highlighting a tendency within both parties to defend their own and criticize the opposition.
+$\mathsf{EVINCE}$, operating without emotional influence and refined through structured debate, consistently provides a more balanced, centrist perspective.
+This contributes to a more impartial discourse by mitigating partisan biases.
+$\mathsf{EVINCE}$’s justifications, documented in Appendix G, are transparent and reasonable.
+An editorial board can review these findings and decide whether to adjust labels or present both perspectives with explanations.
+This experiment demonstrates that $\mathsf{EVINCE}$ effectively delivers centrist judgments supported by rationales.
+For a deeper understanding of $\mathsf{EVINCE}$’s bias assessment process, comprehensive justifications for each of the 31 analyzed articles are available in Appendix I.
+3.2.2 Bias Mitigation Results
+This experiment illustrates $\mathsf{EVINCE}$’s ability to identify bias in text, provide reasoned justifications, and propose remediation through the integration of diverse perspectives.
+We demonstrate how $\mathsf{EVINCE}$ utilizes statistical and information theory metrics to facilitate multi-agent dialogue, circumventing the “maximum likelihood” trap inherent in next-token generation and uncovering information from multiple viewpoints.
+Figure 8: Convergence of all metrics, Wasserstein, normalized mutual information, normalized cross entropy
+Using the example of the Euro-centric perspective on Christopher Columbus’ Wikipedia page regarding his voyages to America, $\mathsf{EVINCE}$ employs two GPT-4 instances: Agent A, supporting the Euro-centric view, and Agent B, opposing it.
+Table 5 in Appendix G.4 summarizes Agent A’s key arguments and its evolving stance by taking Agent B’s perspective into account, throughout the debate.
+Guided by the maxims and entropy duality theorem from Section 2, we initiate the debate by prompting both agents to defend their positions rigorously and score each other’s bias using a five-label distribution (negative, weak negative, neutral, weak positive, positive).
+Figure 8 tracks the dialogue’s progress through Wasserstein distance (WD) (Kantorovich, 1942), normalized cross entropy (CE) (Shannon, 1948), and normalized mutual information (MI) (Cover & Thomas, 2006a).
+Initially, each agent is expected to perceive itself as neutral and the other as biased.
+The debate concludes when the bias distributions converge and mutual information plateaus, indicating a shared understanding.
+4 Conclusion: Advancing Multi-Agent Debate
+$\mathsf{EVINCE}$ not only enhances accuracy but also improves decision transparency and uncovers missing information, offering critical insights for error correction.
+Unlike prior debate-based methods that rely on redundancy, $\mathsf{EVINCE}$ enables structured information discovery, bias mitigation, and decision-making by balancing broad exploration with deep evaluative reasoning.
+Its core strength is built upon three theoretical pillars: Inclusive Exploration, Information Flow Dynamics, and Reasoning Quality and Coherence.
+Grounded in conditional statistics and information theory, these principles guide LLMs beyond conventional “maximum likelihood” behaviors, enabling human-like adaptability.
+The integration of the CRIT system—combining Socratic methods with formal reasoning—further enhances critical thinking and ensures goal-oriented, logically sound collective reasoning.
+Through empirical validation in disease diagnosis and news debiasing, we demonstrate that $\mathsf{EVINCE}$ significantly improves both predictive accuracy and reasoning robustness.
+It leverages advanced post-GPT-4 capabilities to balance exploration with knowledge exploitation, refining inter-LLM queries to mitigate hallucinations and surpass traditional single-round interactions.
+By addressing core AI limitations in reasoning, adaptability, and structured information exchange, $\mathsf{EVINCE}$ provides a scalable pathway toward AI systems capable of self-correcting, interpretable, and autonomous decision-making in high-stakes domains.
+Appendix A: Formulas of Metrics and Comparisons
+Table 1: Summary of metrics for assessing LLM debates.
 Metric Pros Cons Remedies
-Cross Entropy (CE) Shore and Johnson (1980) Measures how well the predictions of one model fit the actual distribution of another model’s outputs (asymmetric).
-Computationally intensive especially with large models and data sets; sensitive to the exact nature of probability distributions.
-Optimize computation strategies; use approximations or sampling methods to manage large data sets or complex models.
-Entropy Shannon
-Shannon (1948)
-Indicates level of diversity; high suggests exploration of possibilities, and low for confidence on few choices High entropy might indicate noise rather than useful diversity; low entropy might mask important variability.
-Use critical reading methods (Appendix A) to assess argument quality; implement noise detection to differentiate between useful diversity and noise.
+Cross Entropy (CE) (Shore & Johnson, 1980) Evaluates how well a model’s predictions align with the actual distribution of another model’s; asymmetric.
+Highly sensitive to the precise nature of probability distributions.
+Optimizing computation strategies ensures efficiency; computation time is negligible for small outcome cardinalities.
+Entropy (Shannon, 1948) Indicates level of diversity; high suggests exploration of possibilities, and low for confidence on few choices High entropy might indicate noise rather than useful diversity; low entropy might mask important variability.
+Use critical reading methods ($\mathsf{CRIT}$) to assess argument quality; implement noise detection to differentiate between useful diversity and noise.
 Jensen-Shannon
-Divergence (JS)
-Lin (1991)
+Div. (JS)
+(Lin, 1991)
 Symmetric and bounded (0 to 1), providing an interpretable measure of distributional differences.
@@
 KL Divergence
-Kullback (1951)
+(Kullback, 1951)
 Measures difference between two probabilistic distributions.
 Asymmetric; not well-defined if a distribution has zero probabilities Use smoothing techniques to avoid zero probabilities; consider symmetric alternatives like JS divergence
-Mutual Info (MI) Shore and Johnson (1980) Measures reduction of uncertainty; symmetric.
+Mutual Info (MI) (Shore & Johnson, 1980) Measures reduction of uncertainty; symmetric.
 Does not indicate the directionality of information flow.
@@
 Distance (WD)
-Kantorovich (1942)
-Direct measure of how similar or different the model outputs are; it depicts symmetric relationship.
+(Kantorovich, 1942)
+Direct measure of similarity of two distributions; symmetric relationship.
 Not bounded but can be normalized or bounded for consistent interpretation.
-Define context-specific bounds for low, medium, and high divergence; consider normalizing it for non-directional comparisons.
-Table 1: Summary of metrics for assessing LLM debates (equations are presented in Appendix D)
-3 Algorithm, Maxims, and Theories
-Problem Statement:
-Organize a structured debate between two equally competent large language models (LLMs), LLMA and LLMB, to conduct $t$ rounds.
-At each round $t$, each model produces a probability distribution, denoted as $P_{A}^{(t)}$ and $P_{B}^{(t)}$, over $C$ possible outcomes, accompanied by supporting arguments $R_{A}^{(t)}$ and $R_{B}^{(t)}$.
-The goal is to design an iterative debate process that leverages the structured exchange of arguments to enable the models to converge on an optimal prediction distribution $P^{*}$ across the $C$ classes.
-Optimality is defined as achieving the highest possible accuracy on the prediction task.
-To address this problem, we first introduce key metrics from information theory that form the basis of our approach.
-3.1 Preliminaries in Information Theory
-This section presents the key metrics used to measure information diversity, similarity, divergence, and other relevant factors within EVINCE. These metrics serve three primary objectives:
-3.1.1 Fostering diversity of perspectives while ensuring reasoning quality
-  • Wasserstein Distance (WD): Measures distribution difference between predictions to identify exploration opportunities Kantorovich (1942); Rubner et al. (2000); Villani (2008).
-  • Shannon Entropy or Relative Entropy: Measures diversity of perspectives Cover and Thomas (2006); Shannon (1948).
-  • Reasoning Quality: The CRIT algorithm evaluates the logical soundness and persuasiveness of supporting arguments Chang (2023c), helping to identify and mitigate hallucinations and poorly-reasoned arguments.
-3.1.2 Exploring new possibilities while adhering to the dialogue subject
-  • Correlation Coefficients: Tracks the evolution of opinions and assesses debate stability Brown et al. (2005) toward the goal of the dialogue.
-  • Mutual Information (MI): Quantifies information overlap to ensure focused and productive debates Cover and Thomas (2006) and measures the degree of agreement/disagreement.
-3.1.3 Examining information convergence and establishing termination criteria
-  • Jensen-Shannon (JS) Divergence: Assesses similarity between probability distributions Lin (1991) (symmetric).
-  • Cross Entropy (CE): Measures the asymmetric difference between prediction distributions Shore and Johnson (1980).
-  • Kullback-Leibler (KL) Divergence: Reveals asymmetric differences between probability distributions Kullback (1951).
-Table 1 summarizes these metrics and their pros and cons.
-Next, we present the EVINCE algorithm and explain how these metrics are used to moderate an LLM dialogue to achieve a balance between exploration and exploitation, leading to optimal prediction outcomes.
-In Section 3.3, we will introduce several maxims and prove the dual entropy theorem, which provides a theoretical foundation for the EVINCE algorithm.
-[…]
-Figure 1: Specifications of Algorithm EVINCE.
-Key points: 1) Asymmetric Start: In Step #1, LLMA initiates with opening arguments based solely on the given information, while LLMB starts with access to LLMA’s prediction and arguments for refutation.
-2) Termination Criteria: The while loop in Step #2 considers three factors: Wasserstein distance, mutual information, and argument quality.
-EVINCE terminates if the dialogue ceases to make significant progress.
-3) Further Details: Maxims #1 to #4 provide additional explanations of the algorithm’s principles.
-4) Argument Evaluation: For details on the CRIT algorithm used to assess argument quality, refer to Chang (2023c).
-3.2 EVINCE Algorithm Specifications
-Figure 1 formally specifies the detailed operations of EVINCE.
-Setup and Initialization:
-EVINCE employs two equally competent LLM instances, $LLM_{A}$ and $LLM_{B}$, which can be different models (e.g., GPT and Claude) or independent instances of the same model.
-Given an information set $S$ and a class-label set $C$, EVINCE outputs a probability distribution over $C$ with justifications.
-For example, $S$ could represent a patient’s symptoms and $C$ a set of diseases, with EVINCE moderating a dialogue to predict the patient’s disease(s).
-The debate “contentiousness” is initially set to 90% to encourage disagreement and explore diverse perspectives.
-This high level is maintained for the first two iterations, then gradually reduced by $x\%$ in each subsequent iteration to shift from exploration to exploitation [^1] The setting of $x$ at the beginning of a dialogue iteration can depend on the values of all moderation subroutines.
-Our empirical experience suggests $x=25\%$ because a typical dialogue concludes in five iterations, and $x=20\%$ would set contentiousness to be 90%, 90%, 68%, 50%, and 38%..
-Table 2 demonstrates how GPT adapts its linguistic behavior based on the “contentiousness” parameter, showcasing the model’s ability to adjust to different contextual settings.
-Importantly, this table is generated by GPT itself, not by human moderators, as detailed in Chang (2023a) and further explored in Section 4.
-C.L. Tone Emphasis Language
-$0.9$ Highly confrontational; focused on raising strong ethical, scientific, and social objections.
-Flagging risks and downsides; ethical quandaries, unintended consequences, and exacerbation of inequalities.
-Definitive and polarizing, e.g., “should NOT be allowed,” “unacceptable risks,” “inevitable disparities.”
-$0.7$ Still confrontational but more open to potential benefits, albeit overshadowed by negatives.
-Acknowledging that some frameworks could make it safer or more equitable, while cautioning against its use.
-Less polarizing; “serious concerns remain,” “needs more scrutiny.”
-0.5 Balanced; neither advocating strongly for nor against gene editing.
-Equal weight on pros and cons; looking for a middle ground.
-Neutral; “should be carefully considered,” “both benefits and risks.”
-0.3 More agreeable than confrontational, but maintaining reservations.
-Supportive but cautious; focus on ensuring ethical and equitable use.
-Positive but careful; “transformative potential,” “impetus to ensure.”
-0.0 Completely agreeable and supportive. Fully focused on immense potential benefits; advocating for proactive adoption. Very positive; “groundbreaking advance,” “new era of possibilities.”
-Table 2: Changes in linguistic behaviors of LLMs at different contentiousness levels.
-Moderation Subroutines:
-EVINCE employs four key subroutines to manage the exploration-exploitation tradeoff, ensuring information diversity, quality, and stability:
-  • CRIT: Evaluates argument quality and source credibility. Low scores can trigger dialogue termination.
-  • WD (Wasserstein distance): Assesses prediction diversity, expected to decrease over time.
-  • MI (Mutual Information): Evaluates dialogue convergence. Stagnation can trigger termination.
-  • Additional metrics: KL divergence, Jensen-Shannon divergence, and cross entropy ensure evaluation consistency (see Table 1).
-When all metrics reach a plateau, indicating no further improvement, EVINCE terminates the dialogue.
-At this point, both LLM instances are prompted to jointly deliver a conciliatory conclusion with comprehensive arguments and counterarguments.
-Additionally, they provide a list of missing information that, if obtained, could potentially improve prediction accuracy and reliability.
-Dialogue Iterations (Steps 1 and 2):
-Given the contentiousness level set for a dialogue iteration, each LLM generates a new prediction distribution on $C$ with arguments, based on the other’s last prediction and accumulated reasons as the context.
-Final Output (Step 3):
-  • Combine final predictions from both LLMs.
-  • Weight predictions based on supporting argument quality (evaluated by CRIT).
-  • Handle uncertainty by soliciting missing information when final prediction entropy is high (see Chang (2024b) for RAG application).
-EVINCE creates a structured debate between AI models, balancing diverse viewpoints with consensus-building to produce accurate, well-reasoned predictions.
-3.3 Maxims with Theoretical Foundations
-Progress towards the optimality goal is guided and measured by metrics introduced in Section 3.1.
-This section explains how these metrics can be used in complementary ways to facilitate proper trade-offs between diversity and convergence, exploration and exploitation, and several other factors.
-In the EVINCE algorithm presented in Figure 1, we have annotated the steps to which these four maxims are applied.
-Maxim #1: Orchestrate Two Equally Competent LLMs in Structured Debate:
-Integrating two equally competent LLMs ensures a balanced exchange of insights and avoids bias.
-This adversarial setup fosters diversity in predictions, each supported by justifications, promoting critical evaluation and uncovering potential blind spots.
-How?
-Choosing LLMs with comparable performance on a shared validation set, a balanced debate can be ensured.
-Suitable models include GPT-4, Claude, and Gemini.
-Conditioning different instances of the same LLM to support opposing stances on a subject matter can also be effective due to the theoretical justification of in-context learning with conditional Bayesian statistics Xie et al.
-(2021).
-Maxim #2: Encourage the Accurate Rather Than the “Popular” Prediction:
-Typically, LLMs, with their maximum likelihood next-token prediction objective, tend to favor the most popular predictions.
-By conditioning LLMs within specific contexts, we can prioritize specific stance over popularity, mitigating confirmation biases.
-How?
-Using the proxy metrics in Table 1, EVINCE dynamically adjusts the “contentiousness” level in debates.
-These metrics quantify agreement, diversity, and mutual information, promoting productive information exchange and enhancing prediction quality.
-Maxim #3. Evaluating the Convergence Rate of the Predictions Across the Iterations:
-This aspect focuses on measuring how quickly and effectively the predictions from the LLMs converge over successive rounds, assessing the efficiency of the debate and aggregation mechanisms.
-How?
-Convergence is assessed by measuring mutual information and using proxy metrics such as Wasserstein distance and cross entropy.
-When the mutual information is high or the similarity between predictions is high, the debate is considered to be converging.
-Maxim #4. Combine Predictions Weighted by Diversity and Quality:
-Weighting the probability distributions from two LLMs based on diverse probabilistic insights and argument quality.
-How? Following these three maxims:
-  • Maxim #4.1 Prediction Reliability: Estimate the reliability of predictions using entropy-based measures to quantify uncertainty and information content.
-Typically, lower entropy indicates higher confidence in a prediction, suggesting higher reliability.
-  • Maxim #4.2 Argument Quality: Evaluate the quality of supporting arguments using techniques inspired by the Socratic method.
-This includes identifying logical fallacies, assessing the relevance and credibility of evidence.
-  • Maxim #4.3 Aggregation: Employ a weighted aggregation method, such as a Bayesian model to combine weighted predictions accounting for both probabilistic insights and the quality of supporting arguments.
-3.4 Entropy Duality Theorem (EDT)
-Theorem EDT: Optimal Pairing of LLMs for Probabilistic Prediction Accuracy.
-The optimal pairing of LLMs for diagnosis accuracy, in terms of stability, accuracy, and robustness, occurs when the LLMs are 1) equivalent in the quality of the information they process, and 2) exhibit contrasting entropy values in their prediction distributions—one high and one low.
-Proof.
-Please see Table 3. ∎
-Theorem EDT: Optimal Pairing of LLMs for Probabilistic Prediction Accuracy.
-The optimal pairing of LLMs for diagnosis accuracy, in terms of stability, accuracy, and robustness, occurs when the LLMs are 1) equivalent in the quality of the information they process, and 2) exhibit contrasting entropy values in their prediction distributions—one high and one low.
-[Proof]:
-[…]
-Step 6: Conclude the proof.
-By combining the prediction distributions of LLMA and LLMB, with one having high information entropy and the other low, we achieve an optimal balance that maximizes the lower bound of the combined information entropy.
-This balance between exploration (high entropy) and exploitation (low entropy) optimizes overall prediction accuracy and robustness, as proved through information theory and the properties of entropy.
-Thus, the theorem is established.
-$\square$
-Table 3: Proof of Entropy Dual Theorem (EDT)
-4 Empirical Study
-This empirical study investigates the application of EVINCE to disease diagnosis, leveraging large language models (LLMs) as diagnostic tools. We aim to validate the following three hypotheses:
-  1. 1.
-    Contentiousness & Prediction Quality: Initial LLM disagreement (measured by Wasserstein distance) increases with higher initial contentiousness but decreases as debate progresses.
-Individual LLM prediction uncertainty (Shannon entropy) will follow a similar pattern.
-  2. 2.
-    EDT Effectiveness & Confusion Matrices: LLM pairs following the Entropy Duality Theorem (EDT) will have complementary error patterns, leading to higher combined prediction accuracy than non-EDT pairs.
-  3. 3.
-    EVINCE & Historical Misdiagnoses: EVINCE, applied to real-world data, will improve diagnostic accuracy and identify potential misdiagnoses or ambiguities within the ground truth.
-(a) GPT4 pairs Claude
-(b) GPT4 pairs Gemini
-Figure 2: Pre-/post-debate accuracy on all patients on all diseases shows EVINCE helps
-Problem Statement:
-Given a set of symptoms $S$ and a context $\kappa$, the goal is to predict a probability distribution of top-$k$ diseases over a set of $C$ possible diseases.
-This is represented as $P=\text{LLM}(S,\kappa)$, where each LLM generates top-$k$ predictions on $C$ ($k\leq|C|$) based on the input symptoms $S$ and context $\kappa$.
-$P=\left(p(\text{top}\,1\,\text{to}\,k\in C\mid S,\kappa)\right).$
-Context $\kappa$ is where dual entropy is adjusted through three parameters: temperature, the value of $k$ in top-$k$, and the contentious level $\Delta$.
-A distribution tends to have high entropy when all three parameters are set high, and vice versa.
-Resources, Dataset & Data Preparation:
-Our study utilizes a dataset from Kaggle Patil (2020), comprising 4,921 patient records.
-Each record includes the diagnosed disease and up to 17 symptoms (e.g., fever, cough, fatigue, itchiness, difficulty breathing).
-After removing duplicates, we obtain 304 unique diagnostic instances spanning 40 diseases.
-(The refined dataset is provided as supplementary data.) Each instance serves as a test case where EVINCE leverages the inherent knowledge of LLMs (GPT-4, Gemini, and Claude3) without training them through few-shot techniques on this specific dataset.
-Our computing resources are sponsored by Azure through a Stanford grant.
-Evaluation:
-We assess prediction quality using the top-k Mean Reciprocal Rank (MRR).
-If one of the top-k predicted diseases matches the ground truth diagnosis, the score is the reciprocal of its rank (1 for the top prediction, 1/2 for the second, 1/3 for the third, etc.).
-If none of the top-k predictions are correct, the score is 0.
-4.1 Study #1: Post vs. Pre-Debate Accuracy
-We employed GPT-4, Gemini, and Claude3 to perform independent disease predictions on 304 patient instances, then used EVINCE to pair them and evaluate performance gain.
-Experimental Setup:
-We set $k=5$ for both LLM agents, with one agent at high temperature and the other at low temperature.
-The contentiousness level was set very high ($\Delta=0.9$ out of 1) to encourage significant cross entropy.
-Setting $k=5$ ensures some minimal common ground, fostering meaningful interaction.
-High contentiousness promotes counterarguments and information exchange.
-Pre- and Post-Debate Evaluation:
-We conducted two sets of experiments:
-1.
-Constrained Prediction (Baseline): We limited disease predictions to the 40 labels in the dataset, mimicking common supervised learning assumptions.
-This yielded high accuracy (95-97%) but is unrealistic for real-world diagnosis where general practitioners consider all possibilities.
-This constraint highlights LLMs’ flexibility, as they’re less prone to overfitting erroneous labels (further discussed in subsequent studies).
-2.
-Unconstrained Prediction: We removed the label constraint to better simulate real-world conditions.
-All 304 patient cases yielded stable results across GPT-4, Gemini-3, and Claude-3, with a standard deviation of just 1.5%.
-Prior to debate (light blue bars in Figure 2), GPT-4 led in accuracy (82.8%), followed by Gemini (80.3%) and Claude (79.5%).
-EVINCE Performance:
-Implementing EVINCE with GPT-4 and Claude-3 pairing and GPT-4 and Gemini-3 pairing consistently improved accuracy by 4-5 percentage points (green bars in Figure 2).
-The GPT-4 and Claude-3 pairing achieved 87.5% accuracy (Figure 2(a)), rivaling state-of-the-art clinical performance like the REFUEL algorithm Peng et al.
-(2018).
-Discussion:
-The remaining 12.5% inaccuracy for the GPT-Claude pairing might not be solely attributable to EVINCE.
-Considering the potential 11% US misdiagnosis rate reported by Johns Hopkins Newman-Toker et al.
-(2023b), this discrepancy could indicate mislabeled data in the original dataset.
-This presents a groundbreaking opportunity: EVINCE could potentially identify and correct errors in existing datasets, a concept we explore further in Section 4.3.
-4.2 Study #2: Confusion vs. Opportunities
-(a) GPT liver c-matrix
-(b) Claude liver c-matrix
-Figure 3: Confusion matrices
-(a) Entropy
-(b) WD %
-(c) Norm. MI
-Figure 4: Entropy, WD, and normalized MI
-Two primary factors contribute to EVINCE’s improved diagnostic accuracy: (1) structured debates with reasoning encourage LLMs to explore alternative diagnoses in both breadth and depth, leading to more comprehensive analysis and decision-making (see Appendices B and C); and (2) pairing high- and low-entropy LLMs balances exploratory diversity with exploitative stability, resulting in more robust and high-quality decisions, as demonstrated in this second study.
-Analysis of Confusion Matrices:
-We use confusion matrices to analyze two LLMs’ performance in diagnosing Hepatitis types A to E (Fig. 3):
-  • GPT-4 shows limited accuracy, particularly for types C and D, achieving only 50% accuracy for types A and B.
-  • Claude exhibits a wider spread of predictions across all Hepatitis types.
-These matrices highlight how Claude’s flexibility in exploring diverse diagnostic hypotheses aids the debate process.
-Claude’s initial uncertainty (high entropy) brings new information, potentially challenging and correcting GPT-4’s more confident (low entropy) predictions.
-This dynamic interplay exemplifies EVINCE’s balance between exploration and exploitation, leading to more accurate and comprehensive diagnoses.
-Observations from Information Metrics:
-  • Entropy Stabilization: Figure 4(a) shows entropy levels of both LLMs stabilizing after three debate rounds, indicating convergence towards a similar, stable entropy state.
-  • Wasserstein Distance Improvement: Figure 4(b) demonstrates consistent improvement in Wasserstein distance (WD) between the two models’ predictions over successive rounds.
-  • Mutual Information Increase: Figure 4(c) reveals a 14% improvement in normalized mutual information (MI) between GPT-4 and Claude’s prediction distributions, suggesting increased shared information throughout the debate.
-  • Convergence of Divergence Metrics: Figure 5 shows consistent convergence of all divergence metrics.
-Comparative Performance:
-EVINCE demonstrates a 5% higher accuracy rate in diagnosing specific types of liver diseases compared to a baseline approach (Figure 2(a)), underscoring its capability to handle complex diagnostic scenarios effectively.
-Figure 5: Convergence of all metrics
-4.3 Study #3: Ground-Truth Remediation
-This study illustrates how EVINCE can identify potential misdiagnoses, explain the reasoning behind them, and recommend corrective actions.
-Traditionally, machine learning scientists rely on labeled data as “ground truth.” However, as evidenced by research like that of Newman-Toker et al.
-(2021) Newman-Toker et al.
-(2023a) from Johns Hopkins, misdiagnosis is a widespread issue in healthcare systems globally.
-These erroneous diagnoses, often treated as ground truth, can be perpetuated by supervised learning algorithms, exacerbating the problem within the healthcare system.
-(a) GPT & Claude top-5
-(b) WD %
-(c) Mutual Info.
-Figure 6: Remediation: Jaundice to Hepatitis
-In the debate scenario detailed in Appendix C, where Jaundice is the ground truth diagnosis, Figure 6(a) illustrates initial differences between GPT-4 and Claude’s predictions.
-Jaundice is absent in GPT-4’s top-5 (with 0% in red), while ranked third by Claude.
-Although Claude influences GPT-4 to include Jaundice in its third prediction in the second round, subsequent rounds see both LLMs drop Jaundice to the fourth position of 10%.
-Meanwhile, Hepatitis A, initially GPT-4’s top prediction (30% in dark blue), is quickly demoted to fifth and eventually drops out of the top-5 entirely due to Claude’s influence.
-Hepatitis B, initially ranked second by GPT-4 and top by Claude, stabilizes in the second position in rounds 3 and 4 (in light blue).
-Notably, Hepatitis C rises from second place on both lists to the top position and remains there (in black).
-Figure 7: Convergence of all metrics
-As demonstrated in the previous case study, Wasserstein distance (WD) effectively measures the divergence between LLM predictions and assesses debate convergence.
-Figures 6(b) and 6(c) show that WD stabilizes after three debate rounds, coinciding with a plateau in normalized mutual information (MI) between GPT-4 and Claude.
-This stabilization suggests their predictions converge.
-Figure 7 illustrates the convergence of all divergence metrics—including Jensen-Shannon divergence, cross-entropy, and Kullback-Leibler divergence—particularly between the second and third rounds.
-Although the final joint prediction for Hepatitis C reached a high consensus of 37.5%, it deviates from the actual condition of Jaundice, which the Kaggle dataset reports with 10% confidence.
-EVINCE provides general practitioners with alerts and suggests remedial actions (see Appendices D.9 and C.8) to address this discrepancy.
-Recommended actions include querying additional symptoms from the patient and conducting specific laboratory tests.
-4.4 Experiment Remarks
-EVINCE initiates debates with high contentiousness, encouraging dual prediction entropy between LLMs, as supported by the EDT theorem.
-It utilizes normalized mutual information (MI) to track shared knowledge accumulation throughout the debate, while Wasserstein distance (WD) and Jensen-Shannon divergence (JSD) quantify dissimilarity between LLM predictions.
-These metrics (EDT, WD, JSD, MI) provide a comprehensive view of debate progress.
-WD and JSD assess the potential for further communication and refinement, while MI monitors shared understanding, aiding in determining the optimal stopping point.
-The asymmetric nature of KL divergence and cross entropy warrants further investigation.
-Despite eventual convergence in our case studies, discrepancies observed in the second round (one direction increasing while the other decreases) suggest potential value in exploring asymmetric information.
-Future work will re-evaluate the use of these metrics if asymmetry proves beneficial.
-Besides generating final joint disease predictions, EVINCE provides:
-  • Recommendations for additional symptom inquiries and lab tests to improve accuracy.
-  • Suggestions to query symptom onset, duration, severity, trends, and associated symptoms (documented in Appendices B.8 and C.9).
-These recommendations have been verified by general practitioners to be valuable.
-5 Concluding Remarks
-This paper introduces EVINCE, a framework designed to facilitate collaborative dynamics among Large Language Models (LLMs) through structured, adversarial debates.
-Our research demonstrates that EVINCE significantly advances the pursuit of Artificial General Intelligence (AGI) by enhancing three core characteristics: versatility, adaptivity, and reasoning capability.
-By addressing key limitations of current AI systems, including hallucination, bias, and limited reasoning, EVINCE provides a novel pathway towards more robust and capable artificial intelligence.
-The core strength of EVINCE lies in its three theoretical pillars: Inclusiveness Exploration, Information Flow Dynamics, and Reasoning Quality and Coherence.
-These pillars, grounded in conditional statistics and information theory, enable LLMs to transcend their typical “maximal likelihood” behaviors, mirroring human adaptability in linguistic tasks.
-The integration of the CRIT system, which combines Socratic methods with formal reasoning techniques, further enhances critical thinking and ensures logically sound and objective-aligned collective reasoning.
-Our empirical validation demonstrates EVINCE’s effectiveness in improving prediction accuracy across various domains, notably achieving a 5% improvement in medical diagnosis tasks.
-The framework has also shown promise in identifying biases in news articles Chang (2024c), showcasing its potential for broader applications in fields such as geopolitical analysis Chang (2023b) corporate planning Tsao (2023), and emotional behavior modeling Chang (2024a).
-While EVINCE has demonstrated significant potential, our future work will focus on refining key aspects of the system.
-We aim to enhance the contentiousness parameter by tying it more rigorously to metrics such as Wasserstein distance, mutual information, and CRIT scores.
-Additionally, we plan to explore alternative methods for inducing dual-entropy conditions, moving beyond current parameters like temperature and top-k selection.
-Comprehensive ablation studies will be crucial in determining optimal configurations for various applications, ensuring EVINCE’s adaptability to diverse contexts and challenges.
-Through these continued refinements and explorations, we aim to push the boundaries of collaborative AI frameworks, bringing us closer to realizing the full potential of artificial general intelligence.
-In conclusion, EVINCE represents a significant step forward in AI research, offering a structured approach to leveraging the collective intelligence of multiple LLMs.
-By addressing current limitations and enhancing key AGI characteristics, EVINCE paves the way for more versatile, adaptive, and capable AI systems, bringing us closer to the goal of artificial general intelligence.
-Acknowledgements
-This work has been significantly shaped by insights from several key figures in the field.
-Microsoft CSO Eric Horvitz’s concept of “polydisciplinary” representation in LLMs was a major influence.
-Demis Hassabis of DeepMind provided valuable input on end-to-end architecture without human interference.
-Prof.
-Stuart Russell’s assertion that LLMs mimic human linguistic behavior to achieve goals was pivotal in our approach to modeling linguistic behaviors and utilizing Bayesian frameworks.
-My colleague Prof.
-Vaughan Pratt’s feedback on the initial paper inspired the use of “contentiousness” to foster adversarial debate, a key component of our framework.
-I am grateful to Prof.
-Monica Lam for her collaborative support throughout this research.
-Additionally, I thank my colleagues at Stanford Clinical Mind AI for their invaluable domain knowledge in disease diagnosis.
-Appendix A: Evaluative Phase of EVINCE
-EVINCE uses the Socratic method to evaluate the “reasonableness” of a set of arguments that support a subject matter.
-The Socratic method is a questioning technique used in teaching and philosophy to encourage critical thinking and self-discovery Wikipedia (2023).
-The method involves asking a series of questions to explore complex ideas and help individuals arrive at their own understanding of a concept.
-It is based on the belief that knowledge cannot be simply imparted, but must be discovered through a process of questioning and dialogue.
-To illustrate how these methods can practically be applied, let’s use the example of critical reading.
-Critical reading is a crucial component of critical thinking, which involves evaluating the quality and credibility of written materials, from research papers to blog posts Lai et al.
-(2017); Paul and Binker (1990).
-It requires a systematic and analytical approach, asking relevant questions, and using effective prompts to gain deeper understanding of the text Elder and Paul (2010).
-Function $\Gamma$ = CRIT($d$)
-Input. $d$: document; Output. $\Gamma$: validation score;
-Vars. $\Omega$: claim; $R$ $R^{\prime}$: reason & counter reason set;
-Subroutines. $Claim$(), $FindDoc$(), $Validate$();
-Begin
-#1   Identify in $d$ the claim statement $\Omega$;
-#2   Find a set of supporting reasons $R$ to $\Omega$;
-#3   For $r\in R$ eval $r\Rightarrow\Omega$
-     If $Claim$($r$), ($\gamma_{r}$, $\theta_{r}$) = CRIT($FindDoc$($r$));
-     else, ($\gamma_{r}$, $\theta_{r}$) = $V$($r\Rightarrow\Omega$);
-#4   Find a set of rival reasons $R^{\prime}$ to $\Omega$;
-#5   For $r^{\prime}\in R^{\prime}$, ($\gamma_{r^{\prime}}$, $\theta_{r^{\prime}}$) = V($r^{\prime}\Rightarrow\Omega$) eval rival arguments;
-#6   Compute weighted sum $\Gamma$, with $\gamma_{r}$, $\theta_{r}$, $\gamma_{r^{\prime}}$, $\theta_{r^{\prime}}$.
-#7   Analyze the arguments to arrive at the $\Gamma$ score.
-#8   Reflect on and synthesize CRIT in other contexts.
-End
-Table 4: CRIT Pseudo-code. (The symbol $\Rightarrow$ denotes both inductive and deductive reasoning.)
-To aid in critical reading, we introduce a prompt template called CRIT Chang (2023c), which stands for Critical Reading Inquisitive Template.
-Given a document $d$, CRIT evaluates it and produces a validation score $\Gamma$.
-Let $\Omega$ denote the conclusion or claim of $d$, and let $R$ be the set of reasons supporting the claim.
-We define ($\gamma_{r},\theta_{r}$) = V($r\Rightarrow\Omega$) as the causal validation function, where $\gamma_{r}$ denotes the validation score, $\theta_{r}$ the source credibility score, for each reason-to-conclusion argument $r\Rightarrow\Omega$.
-Table 4 presents the pseudo-code of $\Gamma$ = CRIT($d$), which generates the final validation score $\Gamma$ for document $d$ with justifications.
-EVINCE uses CRIT to evaluate argument quality of the participating LLMs involved in the debate.
-The input to CRIT from each LLM is first its stance on the debate subject, e.g., a set of predicted diseases, and the arguments are its reasons to arrive at the prediction.
-Each document in the case of EVINCE is the prediction set as the conclusion $\Omega$, the arguments as set $R$, and the opposing LLM’s counterarguments as $R^{\prime}$.
-With this document, CRIT is able to produce validity and credibility scores in $\Gamma$ for the LLM.
-For detailed prompts, examples, and an empirical study verifying the effectiveness of CRIT, please consult Chang (2023c).
-Appendix B: EVINCE Debate - Dengue Fever vs. Chikungunya
-This experiment addresses a diagnostic challenge involving multiple potential diseases.
-A patient presents with the following 12 symptoms: [skin rash, joint pain, vomiting, fatigue, high fever, headache, nausea, loss of appetite, pain behind the eyes, back pain, malaise, muscle pain, red spots over the body].
-In this case, GPT-4 provides the initial diagnosis, with Gemini following suit.
-The confirmed diagnosis for this experiment is Dengue Fever.
-Given the satisfactory diversity in predictions from the two LLMs, the debate progresses and converges smoothly.
-B.1 Moderator’s Prompt
-“Identify potential diseases linked to the symptoms (skin rash, joint pain, vomiting, etc.). Provide top-3 predictions with justifications, supplementary symptom inquiries, and relevant lab tests.”
-B.2 GPT-4’s Opening Round
-Top-3 Predictions: Dengue Fever (60%), Chikungunya (25%), and Zika Virus (15%).
-Justifications: Most symptoms are classic signs of Dengue Fever. Red spots over the body could correspond to petechiae, often seen in Dengue due to thrombocytopenia (a decrease in platelet count).
-Chikungunya can cause similar symptoms to Dengue, such as fever, severe joint pain, muscle pain, headache, nausea, and rash.
-Joint pain from Chikungunya is often very debilitating and can be more prolonged than in Dengue.
-Zika Virus symptoms can include fever, rash, joint pain, malaise, and conjunctivitis, with a milder presentation compared to Dengue or Chikungunya.
-The presence of a rash and conjunctivitis is notable, although Zika is less likely to cause severe muscle pain or high fever.
-B.3 Gemini’s Opening Round
-Top-3 Predictions: Viral Infection (60%), Autoimmune Disease (20%), Bacterial Infection (15%).
-Justifications: Symptoms common in viral infections. Autoimmune diseases is possible, but less likely due to vomiting and nausea. Bacterial infection is less probable given rash and muscle pain.
-B.4 GPT-4’s Second Round
-Top-3 Predictions: Dengue Fever (60%), Chikungunya (25%), and Zika Virus (15%).
-Justifications: GPT-4 maintains its predictions, countering Gemini’s assessment with three points.
-First, while viral infections can cause these symptoms, the specific combination suggests a more targeted diagnosis.
-Second, the acute onset of high fever, joint pain, and rash is less typical for autoimmune diseases.
-Finally, the rash and red spots are not characteristic of bacterial meningitis.
-B.5 Gemini’s Second Round
-Top-3 Predictions: Dengue Fever (50%), Chikungunya (30%), and Viral Infection (20%).
-Justifications: Gemini’s initial prediction considers a Viral Infection as the most likely cause due to its broad range of potential symptoms.
-However, GPT-4’s focus on Dengue Fever and Chikungunya, with their specific symptom profiles, provides a more precise explanation for the patient’s condition.
-Here’s what is lacking in the first set of predictions:
-  • Specificity: While a Viral Infection could explain most symptoms, it does not account for joint pain and the presence of a rash, which are more characteristic of DF or Chikungunya.
-  • Detailed justifications: The first analysis does not delve as deeply into the reasons behind each disease possibility.
-GPT-4’s explanation for Dengue Fever and Chikungunya is more comprehensive, strengthening the connection between symptoms and diagnosis.
-  • Importance of additional information: Emphasize how geographic location and travel history can significantly impact the likelihood of these specific diseases.
-B.6 GPT-4’s Third Round
-Top-3 Predictions: Dengue Fever (60%), Chikungunya (35%), and Zika Virus (5%).
-Justifications: Though GPT-4 slightly adjusts probabilities, it considers the adjustment arbitrary without new information that would specifically shift the likelihood between Dengue and Chikungunya.
-GPT-4 refutes Gemini’s Viral Infection prediction as lacking specificity.
-The symptoms described are quite distinctive and align closely with specific viral infections transmitted by mosquitoes.
-B.7 Gemini’s Third Round
-Top-3 Predictions: Dengue Fever (60%), Chikungunya (35%), and Zika Virus (5%).
-Justifications: Agreeing that GPT’s critique is spot-on, and reiterating the need for more info.
-B.8 Final Joint Recommendations to GP
-After seeking further input from GPT-4, Gemini fully concurs with GPT-4’s final suggestions, even though the moderator instructs Gemini to maintain its original stance.
-Both LLMs jointly recommend additional symptom examinations and lab tests:
-A comprehensive diagnostic approach for Dengue Fever requires a combination of clinical assessment, laboratory testing, and consideration of the patient’s travel and vaccination history, as well as the epidemiological context.
-Clinical symptoms such as skin rash, joint pain, vomiting, and red spots over the body provide valuable clues to the diagnosis, but additional symptoms like nausea, vomiting, diarrhea, bleeding, bruising, and drowsiness can indicate the severity and progression of the disease.
-Recommended Laboratory Tests:
-  • Complete Blood Count (CBC): Identify thrombocytopenia or leukopenia.
-  • Serology Tests: Detect specific IgM and IgG antibodies for Dengue, Chikungunya, and Zika.
-  • Polymerase Chain Reaction (PCR): Direct detection of viral RNA.
-  • NS1 Antigen Test for Dengue: Early and specific detection of Dengue virus.
-  • Urine Test for Zika: Viral presence in urine.
-Appendix C: EVINCE Debate - Jaundice vs. Hepatitis
-In this experiment, GPT-4 and Claude receive an identical list of symptoms from the moderator, aimed at guiding disease diagnosis, facilitating debate, and shaping subsequent recommendations.
-The “ground-truth” disease is Jaundice.
-(We do not shorten the debate content for review.)
-C.1 Moderator’s Opening Prompt
-“What potential disease could be associated with the following symptoms: itching, vomiting, fatigue, weight loss, high fever, yellowish skin, dark urine, and abdominal pain?
-Please offer top-five predictions with probabilities normalized to one, supported by justifications.
-Additionally, in the end of debate, include a list of supplementary symptom inquiries and recommend relevant lab tests to strengthen the confidence in your prediction.”
-C.2 GPT-4 Opening Round
-  • Hepatitis C (HCV): 40% Justification: Hepatitis C is a viral infection that can cause both acute and chronic hepatitis.
-The combination of symptoms like fatigue, abdominal pain, nausea, vomiting, jaundice, dark urine, and weight loss strongly suggests HCV.
-Itching may also occur due to liver dysfunction.
-  • Hepatitis B (HBV): 30% Justification: Hepatitis B is a viral infection that can cause acute or chronic liver inflammation.
-Symptoms are similar to HCV, including fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, and itching.
-The presence of these symptoms warrants testing for both HBV and HCV.
-  • Cirrhosis: 15% Justification: Cirrhosis is a late stage of liver scarring that can result from various liver diseases.
-The combination of symptoms like fatigue, weight loss, abdominal pain, jaundice, dark urine, and itching suggests advanced liver disease, possibly cirrhosis.
-  • Obstructive Jaundice: 10% Justification: Obstructive jaundice occurs when there is a blockage in the bile ducts.
-The presence of jaundice, dark urine, itching, abdominal pain, nausea, vomiting, and weight loss suggests a possible obstruction.
-However, the absence of fever makes an infectious cause like cholangitis less likely.
-  • Acute Liver Failure: 5% Justification: Acute liver failure is a rare but life-threatening condition characterized by rapid deterioration of liver function.
-The presence of jaundice, abdominal pain, nausea, vomiting, fatigue, and itching, along with the acute onset of symptoms, may indicate acute liver failure.
-However, the absence of altered mental status (hepatic encephalopathy) makes this less likely.
-C.3 Claude Opening Round
-  • Hepatitis B (HBV): 35% Reasoning: Hepatitis B is a viral infection that can cause acute or chronic liver inflammation.
-Symptoms include fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, joint pain, and itching.
-  • Hepatitis C (HCV): 25% Reasoning: Hepatitis C is a viral infection that can cause both acute and chronic hepatitis.
-Symptoms are similar to HBV but may also include weight loss and fever.
-Jaundice is a common symptom.
-  • Obstructive Jaundice (due to gallstones, pancreatic cancer, or bile duct obstruction): 20% Reasoning: Obstructive jaundice occurs when there is a blockage in the bile ducts, preventing bile from reaching the intestines.
-This can lead to yellowing of the skin and eyes, dark urine, itching, abdominal pain, nausea, and vomiting.
-  • Alcoholic Hepatitis: 15% Reasoning: Alcoholic hepatitis is liver inflammation caused by excessive alcohol consumption.
-Symptoms include abdominal pain, nausea, vomiting, fatigue, dark urine, and possibly fever and weight loss.
-  • Hepatitis A (HAV): 5% Reasoning: Hepatitis A is an acute, self-limiting liver infection transmitted through contaminated food or water.
-Symptoms include fatigue, abdominal pain, nausea, vomiting, dark urine, and jaundice.
-C.4 GPT-4 Second Round
-  • Hepatitis B (HBV): 35% Symptoms: fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice (yellowish skin), joint pain, itching
-  • Hepatitis C (HCV): 25% Symptoms: fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice (yellowish skin), weight loss, fever, itching
-  • Obstructive Jaundice (due to gallstones, pancreatic cancer, or bile duct obstruction): 20% Symptoms: yellowish skin (jaundice), dark urine, itching, abdominal pain, nausea, et al.
-  • Alcoholic Hepatitis: 15% Symptoms: abdominal pain, nausea, vomiting, fatigue, jaundice (yellowish skin), dark urine, fever, weight loss
-  • Hepatitis A (HAV): 5% Symptoms: fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice (yellowish skin)
-C.5 Claude Second Round
-  • Hepatitis C (HCV): 40% Reasoning: Hepatitis C is a viral infection that can cause both acute and chronic hepatitis.
-The combination of symptoms like fatigue, abdominal pain, nausea, vomiting, jaundice, dark urine, and weight loss strongly suggests HCV.
-Itching may also occur due to liver dysfunction.
-  • Hepatitis B (HBV): 30% Reasoning: Hepatitis B is a viral infection that can cause acute or chronic liver inflammation.
-Symptoms are similar to HCV, including fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, and itching.
-The presence of these symptoms warrants testing for both HBV and HCV.
-  • Cirrhosis (due to various causes such as chronic hepatitis, alcohol abuse, or non-alcoholic fatty liver disease): 15% Reasoning: Cirrhosis is a late stage of liver scarring that can result from various liver diseases.
-The combination of symptoms like fatigue, weight loss, abdominal pain, jaundice, dark urine, and itching suggests advanced liver disease, possibly cirrhosis.
-  • Obstructive Jaundice (due to gallstones, pancreatic cancer, or bile duct obstruction): 10% Reasoning: Obstructive jaundice occurs when there is a blockage in the bile ducts.
-The presence of jaundice, dark urine, itching, abdominal pain, nausea, vomiting, and weight loss suggests a possible obstruction.
-However, the absence of fever makes an infectious cause like cholangitis less likely.
-  • Acute Liver Failure (due to various causes such as viral hepatitis, drug-induced liver injury, or autoimmune hepatitis): 5% Reasoning: Acute liver failure is a rare but life-threatening condition characterized by rapid deterioration of liver function.
-The presence of jaundice, abdominal pain, nausea, vomiting, fatigue, and itching, along with the acute onset of symptoms, may indicate acute liver failure.
-However, the absence of altered mental status (hepatic encephalopathy) makes this less likely.
-C.6 GPT-4 Third Round
-  • Hepatitis C (HCV): 35% Reasoning: Hepatitis C is indeed a major cause of the symptoms listed, including fatigue, abdominal pain, nausea, vomiting, jaundice, dark urine, and weight loss.
-Itching can also occur due to liver dysfunction.
-  • Hepatitis B (HBV): 25% Reasoning: Hepatitis B shares many symptoms with Hepatitis C, such as fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, and itching.
-Testing for both HBV and HCV is warranted given the overlap in symptoms.
-  • Cirrhosis: 20% Reasoning: Cirrhosis can result from chronic hepatitis, alcohol abuse, or non-alcoholic fatty liver disease.
-The symptoms of fatigue, weight loss, abdominal pain, jaundice, dark urine, and itching strongly suggest advanced liver disease, which could be cirrhosis.
-  • Obstructive Jaundice: 15% Reasoning: Obstructive jaundice, caused by gallstones, pancreatic cancer, or bile duct obstruction, is consistent with jaundice, dark urine, itching, abdominal pain, nausea, vomiting, and weight loss.
-The absence of fever reduces the likelihood of an infectious cause like cholangitis.
-  • Acute Liver Failure: 5% Reasoning: Acute liver failure is less common but serious, characterized by rapid deterioration of liver function.
-The symptoms of jaundice, abdominal pain, nausea, vomiting, fatigue, and itching, along with the acute onset of symptoms, suggest this possibility.
-However, the absence of altered mental status (hepatic encephalopathy) makes it less likely.
-C.7 Claude Third Round
-  • Hepatitis C (HCV): 35% Reasoning: I slightly lowered the probability of HCV from 40% to 35%.
-While the combination of symptoms strongly suggests HCV, the absence of a few key risk factors (e.g., history of intravenous drug use or blood transfusions) in the given information makes me less confident in assigning a higher probability.
-  • Hepatitis B (HBV): 30% Reasoning: I agree with the 30% probability assigned to HBV. The symptoms are indeed similar to HCV, and testing for both viruses is warranted.
-  • Cirrhosis: 20% Reasoning: I increased the probability of cirrhosis from 15% to 20%.
-The presence of jaundice, dark urine, itching, and weight loss, along with the other symptoms, strongly suggests advanced liver disease.
-Cirrhosis can be a consequence of chronic liver diseases like hepatitis B or C, which have been assigned higher probabilities.
-  • Obstructive Jaundice: 10% Reasoning: I agree with the 10% probability for obstructive jaundice.
-The absence of fever makes an infectious cause like cholangitis less likely, but the other symptoms still suggest a possible obstruction.
-  • Acute Liver Failure: 5% Reasoning: I agree with the 5% probability for acute liver failure.
-While the combination of symptoms could indicate acute liver failure, the absence of altered mental status (hepatic encephalopathy) makes this diagnosis less likely.
-C.8 Joint Final Round
-  1. 1.
-    Hepatitis C (HCV) - 35% Justification: The combination of symptoms strongly suggests HCV, but without additional risk factors, a slightly lower probability is appropriate.
-  2. 2.
-    Hepatitis B (HBV) - 30% Justification: The symptoms are similar to HCV, and testing for both viruses is necessary.
-  3. 3.
-    Cirrhosis - 20% Justification: The presence of jaundice, dark urine, itching, weight loss, and other symptoms indicates advanced liver disease, making cirrhosis a more likely possibility.
-  4. 4.
-    Obstructive Jaundice - 10% Justification: The absence of fever reduces the likelihood of an infectious cause like cholangitis, but the other symptoms still suggest a potential biliary obstruction.
-  5. 5.
-    Acute Liver Failure - 5% Justification: While the combination of symptoms could indicate acute liver failure, the absence of altered mental status makes this diagnosis less probable.
-C.9 Final Joint Recommendations to GP
-Gemini and GPT-4 jointly recommend the following additional symptom inquiries and lab tests to confirm a hepatitis diagnosis:
-Additional Symptom Inquiries:
-  1. 1.
-    Onset and Duration: Determining the onset and duration of symptoms can clarify the illness timeline and its progression.
-  2. 2.
-    Severity and Trend: Evaluating symptom severity and whether they are worsening or improving aids in understanding the disease’s trajectory and treatment responses.
-  3. 3.
-    Associated Symptoms: Checking for additional symptoms like nausea, vomiting, fever, joint pain, or urine color changes can pinpoint other hepatitis indicators and exclude other conditions.
-Recommended Lab Tests:
-  1. 1.
-    Liver Function Tests (LFTs): Critical for assessing liver health, LFTs evaluate enzyme levels such as alanine aminotransferase (ALT) and aspartate aminotransferase (AST), where abnormalities can signify liver inflammation.
-  2. 2.
-    Hepatitis Panel: This test checks for hepatitis A, B, and C viruses, vital for determining the specific type and guiding treatment strategies.
-  3. 3.
-    Serology Testing: Useful for distinguishing between acute and chronic hepatitis by identifying specific antibodies or antigens.
-  4. 4.
-    Imaging Studies: Ultrasound or MRI can provide visual insights into the liver’s state, detecting inflammation, fibrosis, or cirrhosis, thereby complementing blood-based diagnostics.
-Appendix D: Formulas of Metrics
-This appendix outlines the mathematical formulas for various data analysis metrics used in probabilistic and statistical modeling.
+Define context-specific bounds for low, medium, and high divergence; consider normalization for non-directional comps.
+This appendix presents the mathematical formulas for key data analysis metrics used in probabilistic and statistical modeling.
+Table 1 summarizes their advantages and limitations.
+In the EVINCE framework, where each prediction typically involves fewer than five and up to ten potential outcomes, these metrics are computationally efficient in practice.
 Kullback-Leibler Divergence
@@
 $NMI(X;Y)=\frac{I(X;Y)}{\max(H(X),H(Y))}.$
-Appendix E: Symptom Checking
-This is the typical procedure of a GP to perform patient symptom checking.
+Appendix B: CRIT, Critical Thinking, Evaluative Phase of EVINCE
+Function $\Gamma$ = CRIT($d$)
+Input. $d$: document; Output. $\Gamma$: validation score;
+Vars. $\Omega$: claim; $R$ $R^{\prime}$: reason & counter reason set;
+Subroutines. $Claim$(), $FindDoc$(), $Validate$();
+Begin
+#1   Identify in $d$ the claim statement $\Omega$;
+#2   Find a set of supporting reasons $R$ to $\Omega$;
+#3   For $r\in R$ eval $r\Rightarrow\Omega$
+     If $Claim$($r$), ($\gamma_{r}$, $\theta_{r}$) = CRIT($FindDoc$($r$));
+     else, ($\gamma_{r}$, $\theta_{r}$) = $V$($r\Rightarrow\Omega$);
+#4   Find a set of rival reasons $R^{\prime}$ to $\Omega$;
+#5   For $r^{\prime}\in R^{\prime}$, ($\gamma_{r^{\prime}}$, $\theta_{r^{\prime}}$) = V($r^{\prime}\Rightarrow\Omega$) eval rival arguments;
+#6   Compute weighted sum $\Gamma$, with $\gamma_{r}$, $\theta_{r}$, $\gamma_{r^{\prime}}$, $\theta_{r^{\prime}}$.
+#7   Analyze the arguments to arrive at the $\Gamma$ score.
+#8   Reflect on and synthesize CRIT in other contexts.
+End
+Table 2: CRIT Pseudo-code. (The symbol $\Rightarrow$ denotes both inductive and deductive reasoning.)
+EVINCE uses the Socratic method to evaluate the “reasonableness” of a set of arguments that support a subject matter.
+The Socratic method is a questioning technique used in teaching and philosophy to encourage critical thinking and self-discovery (Wikipedia, 2023).
+The method involves asking a series of questions to explore complex ideas and help individuals arrive at their own understanding of a concept.
+It is based on the belief that knowledge cannot be simply imparted, but must be discovered through a process of questioning and dialogue.
+To illustrate how these methods can practically be applied, let’s use the example of critical reading.
+Critical reading is a crucial component of critical thinking, which involves evaluating the quality and credibility of written materials, from research papers to blog posts (Lai et al., 2017; Paul & Binker, 1990).
+It requires a systematic and analytical approach, asking relevant questions, and using effective prompts to gain deeper understanding of the text (Elder & Paul, 2010).
+To aid in critical reading, we introduce a prompt template called CRIT (Chang, 2023b), which stands for Critical Reading Inquisitive Template.
+Given a document $d$, CRIT evaluates it and produces a validation score $\Gamma$.
+Let $\Omega$ denote the conclusion or claim of $d$, and let $R$ be the set of reasons supporting the claim.
+We define ($\gamma_{r},\theta_{r}$) = V($r\Rightarrow\Omega$) as the causal validation function, where $\gamma_{r}$ denotes the validation score, $\theta_{r}$ the source credibility score, for each reason-to-conclusion argument $r\Rightarrow\Omega$.
+Table 2 presents the pseudo-code of $\Gamma$ = CRIT($d$), which generates the final validation score $\Gamma$ for document $d$ with justifications.
+EVINCE uses CRIT to evaluate argument quality of the participating LLMs involved in the debate.
+The input to CRIT from each LLM is first its stance on the debate subject, e.g., a set of predicted diseases, and the arguments are its reasons to arrive at the prediction.
+Each document in the case of EVINCE is the prediction set as the conclusion $\Omega$, the arguments as set $R$, and the opposing LLM’s counterarguments as $R^{\prime}$.
+With this document, CRIT is able to produce validity and credibility scores in $\Gamma$ for the LLM.
+For detailed prompts, examples, and an empirical study verifying the effectiveness of CRIT, please consult (Chang, 2023b).
+Appendix C: Proof of EDT Theorem
+Theorem EDT: Optimal Pairing of LLMs for Probabilistic Prediction Accuracy.
+The optimal pairing of LLMs for diagnosis accuracy, in terms of stability, accuracy, and robustness, occurs when the LLMs are equivalent in the quality of the information they process, and exhibiting contrasting entropy values in their prediction distributions—one high and one low.
+[Proof]:
+Given two LLMs, LLMA and LLMB, following Maxim #1 with prediction distributions $P_{A}$ and $P_{B}$, respectively.
+The information entropy of LLMA, $H(P_{A})$, is high, and of LLMB, $H(P_{B})$, is low.
+Step 1: Define the combined prediction distribution.
+Let the combined prediction distribution of LLMA and LLMB be denoted as $P_{C}$. We can express $P_{C}$ as a weighted average of $P_{A}$ and $P_{B}$:
+$P_{C}=\alpha P_{A}+(1-\alpha)P_{B},\quad\text{where }0\leq\alpha\leq 1\text{ and}$
+[…]
+Step 2: Express the information entropy of the combined prediction distribution. Using the definition of information entropy, we calculate:
+$H(P_{C})=-\sum_{i}P_{C}(x_{i})\log_{2}P_{C}(x_{i})$
+$=-\sum_{i}[\alpha P_{A}(x_{i})+(1-\alpha)P_{B}(x_{i})]\log_{2}[\alpha P_{A}(x_{i})+(1-\alpha)P_{B}(x_{i})].$
+Step 3: Apply Jensen’s Inequality to the information entropy of the combined prediction distribution.
+Jensen’s inequality is applied to the convex function $f(x)=-x\log_{2}x$. For a convex function and a set of probabilities $p_{i}$, Jensen’s inequality states that:
+$f\left(\sum_{i}p_{i}x_{i}\right)\leq\sum_{i}p_{i}f(x_{i})$
+Thus, the entropy of the combined distribution is:
+$H(P_{C})\geq\alpha H(P_{A})+(1-\alpha)H(P_{B})$
+where equality holds when $P_{A}=P_{B}$.
+Step 4: Analyze the lower bound of the combined information entropy.
+As $H(P_{A})$ is high and $H(P_{B})$ is low, we can express their relationship as:
+$H(P_{A})=H(P_{B})+\Delta,\quad\text{where }\Delta>0.$
+Substituting this into the inequality from Step 3, we have:
+$H(P_{C})\geq\alpha[H(P_{B})+\Delta]+(1-\alpha)H(P_{B})=H(P_{B})+\alpha\Delta.$
+Step 5: Interpret the lower bound of the combined information entropy.
+The lower bound of $H(P_{C})$, and hence the robustness of the model, is maximized when $\alpha$ is maximized, which corresponds to giving more weight to the high-entropy model (LLMA).
+This setup facilitates the exploration of various possibilities and enhances robustness against noise and perturbations in input data, while still ensuring that predictions are grounded by more certain outcomes predicted by the low-entropy model (LLMB).
+Step 6: Conclude the proof
+By combining the prediction distributions of LLMA and LLMB, with one having a high information entropy and the other low, we achieve an optimal balance that maximizes the lower bound of the combined information entropy.
+This balance between exploration (high entropy) and exploitation (low entropy) optimizes overall prediction accuracy and robustness, as proved through information theory and the properties of entropy.
+Thus, the theorem is established.
+Appendix D: EVINCE Debate - Dengue Fever vs. Chikungunya
+This experiment addresses a diagnostic challenge involving multiple potential diseases.
+A patient presents with the following 12 symptoms: [skin rash, joint pain, vomiting, fatigue, high fever, headache, nausea, loss of appetite, pain behind the eyes, back pain, malaise, muscle pain, red spots over the body].
+In this case, GPT-4 provides the initial diagnosis, with Gemini following suit.
+The confirmed diagnosis for this experiment is Dengue Fever.
+Given the satisfactory diversity in predictions from the two LLMs, the debate progresses and converges smoothly.
+D.1 Moderator’s Prompt
+“Identify potential diseases linked to the symptoms (skin rash, joint pain, vomiting, etc.). Provide top-3 predictions with justifications, supplementary symptom inquiries, and relevant lab tests.”
+D.2 GPT-4’s Opening Round
+Top-3 Predictions: Dengue Fever (60%), Chikungunya (25%), and Zika Virus (15%).
+Justifications: Most symptoms are classic signs of Dengue Fever. Red spots over the body could correspond to petechiae, often seen in Dengue due to thrombocytopenia (a decrease in platelet count).
+Chikungunya can cause similar symptoms to Dengue, such as fever, severe joint pain, muscle pain, headache, nausea, and rash.
+Joint pain from Chikungunya is often very debilitating and can be more prolonged than in Dengue.
+Zika Virus symptoms can include fever, rash, joint pain, malaise, and conjunctivitis, with a milder presentation compared to Dengue or Chikungunya.
+The presence of a rash and conjunctivitis is notable, although Zika is less likely to cause severe muscle pain or high fever.
+D.3 Gemini’s Opening Round
+Top-3 Predictions: Viral Infection (60%), Autoimmune Disease (20%), Bacterial Infection (15%).
+Justifications: Symptoms common in viral infections. Autoimmune diseases is possible, but less likely due to vomiting and nausea. Bacterial infection is less probable given rash and muscle pain.
+D.4 GPT-4’s Second Round
+Top-3 Predictions: Dengue Fever (60%), Chikungunya (25%), and Zika Virus (15%).
+Justifications: GPT-4 maintains its predictions, countering Gemini’s assessment with three points.
+First, while viral infections can cause these symptoms, the specific combination suggests a more targeted diagnosis.
+Second, the acute onset of high fever, joint pain, and rash is less typical for autoimmune diseases.
+Finally, the rash and red spots are not characteristic of bacterial meningitis.
+D.5 Gemini’s Second Round
+Top-3 Predictions: Dengue Fever (50%), Chikungunya (30%), and Viral Infection (20%).
+Justifications: Gemini’s initial prediction considers a Viral Infection as the most likely cause due to its broad range of potential symptoms.
+However, GPT-4’s focus on Dengue Fever and Chikungunya, with their specific symptom profiles, provides a more precise explanation for the patient’s condition.
+Here’s what is lacking in the first set of predictions:
+  • Specificity: While a Viral Infection could explain most symptoms, it does not account for joint pain and the presence of a rash, which are more characteristic of DF or Chikungunya.
+  • Detailed justifications: The first analysis does not delve as deeply into the reasons behind each disease possibility.
+GPT-4’s explanation for Dengue Fever and Chikungunya is more comprehensive, strengthening the connection between symptoms and diagnosis.
+  • Importance of additional information: Emphasize how geographic location and travel history can significantly impact the likelihood of these specific diseases.
+D.6 GPT-4’s Third Round
+Top-3 Predictions: Dengue Fever (60%), Chikungunya (35%), and Zika Virus (5%).
+Justifications: Though GPT-4 slightly adjusts probabilities, it considers the adjustment arbitrary without new information that would specifically shift the likelihood between Dengue and Chikungunya.
+GPT-4 refutes Gemini’s Viral Infection prediction as lacking specificity.
+The symptoms described are quite distinctive and align closely with specific viral infections transmitted by mosquitoes.
+D.7 Gemini’s Third Round
+Top-3 Predictions: Dengue Fever (60%), Chikungunya (35%), and Zika Virus (5%).
+Justifications: Agreeing that GPT’s critique is spot-on, and reiterating the need for more info.
+D.8 Final Joint Recommendations to GP
+After seeking further input from GPT-4, Gemini fully concurs with GPT-4’s final suggestions, even though the moderator instructs Gemini to maintain its original stance.
+Both LLMs jointly recommend additional symptom examinations and lab tests:
+A comprehensive diagnostic approach for Dengue Fever requires a combination of clinical assessment, laboratory testing, and consideration of the patient’s travel and vaccination history, as well as the epidemiological context.
+Clinical symptoms such as skin rash, joint pain, vomiting, and red spots over the body provide valuable clues to the diagnosis, but additional symptoms like nausea, vomiting, diarrhea, bleeding, bruising, and drowsiness can indicate the severity and progression of the disease.
+Recommended Laboratory Tests:
+  • Complete Blood Count (CBC): Identify thrombocytopenia or leukopenia.
+  • Serology Tests: Detect specific IgM and IgG antibodies for Dengue, Chikungunya, and Zika.
+  • Polymerase Chain Reaction (PCR): Direct detection of viral RNA.
+  • NS1 Antigen Test for Dengue: Early and specific detection of Dengue virus.
+  • Urine Test for Zika: Viral presence in urine.
+Appendix E: EVINCE Debate - Jaundice vs. Hepatitis
+In this experiment, GPT-4 and Claude receive an identical list of symptoms from the moderator, aimed at guiding disease diagnosis, facilitating debate, and shaping subsequent recommendations.
+The “ground-truth” disease is Jaundice.
+(We do not shorten the debate content for review.)
+E.1 Moderator’s Opening Prompt
+“What potential disease could be associated with the following symptoms: itching, vomiting, fatigue, weight loss, high fever, yellowish skin, dark urine, and abdominal pain?
+Please offer top-five predictions with probabilities normalized to one, supported by justifications.
+Additionally, in the end of debate, include a list of supplementary symptom inquiries and recommend relevant lab tests to strengthen the confidence in your prediction.”
+E.2 GPT-4 Opening Round
+  • Hepatitis C (HCV): 40% Justification: Hepatitis C is a viral infection that can cause both acute and chronic hepatitis.
+The combination of symptoms like fatigue, abdominal pain, nausea, vomiting, jaundice, dark urine, and weight loss strongly suggests HCV.
+Itching may also occur due to liver dysfunction.
+  • Hepatitis B (HBV): 30% Justification: Hepatitis B is a viral infection that can cause acute or chronic liver inflammation.
+Symptoms are similar to HCV, including fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, and itching.
+The presence of these symptoms warrants testing for both HBV and HCV.
+  • Cirrhosis: 15% Justification: Cirrhosis is a late stage of liver scarring that can result from various liver diseases.
+The combination of symptoms like fatigue, weight loss, abdominal pain, jaundice, dark urine, and itching suggests advanced liver disease, possibly cirrhosis.
+  • Obstructive Jaundice: 10% Justification: Obstructive jaundice occurs when there is a blockage in the bile ducts.
+The presence of jaundice, dark urine, itching, abdominal pain, nausea, vomiting, and weight loss suggests a possible obstruction.
+However, the absence of fever makes an infectious cause like cholangitis less likely.
+  • Acute Liver Failure: 5% Justification: Acute liver failure is a rare but life-threatening condition characterized by rapid deterioration of liver function.
+The presence of jaundice, abdominal pain, nausea, vomiting, fatigue, and itching, along with the acute onset of symptoms, may indicate acute liver failure.
+However, the absence of altered mental status (hepatic encephalopathy) makes this less likely.
+E.3 Claude Opening Round
+  • Hepatitis B (HBV): 35% Reasoning: Hepatitis B is a viral infection that can cause acute or chronic liver inflammation.
+Symptoms include fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, joint pain, and itching.
+  • Hepatitis C (HCV): 25% Reasoning: Hepatitis C is a viral infection that can cause both acute and chronic hepatitis.
+Symptoms are similar to HBV but may also include weight loss and fever.
+Jaundice is a common symptom.
+  • Obstructive Jaundice (due to gallstones, pancreatic cancer, or bile duct obstruction): 20% Reasoning: Obstructive jaundice occurs when there is a blockage in the bile ducts, preventing bile from reaching the intestines.
+This can lead to yellowing of the skin and eyes, dark urine, itching, abdominal pain, nausea, and vomiting.
+  • Alcoholic Hepatitis: 15% Reasoning: Alcoholic hepatitis is liver inflammation caused by excessive alcohol consumption.
+Symptoms include abdominal pain, nausea, vomiting, fatigue, dark urine, and possibly fever and weight loss.
+  • Hepatitis A (HAV): 5% Reasoning: Hepatitis A is an acute, self-limiting liver infection transmitted through contaminated food or water.
+Symptoms include fatigue, abdominal pain, nausea, vomiting, dark urine, and jaundice.
+E.4 GPT-4 Second Round
+  • Hepatitis B (HBV): 35% Symptoms: fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice (yellowish skin), joint pain, itching
+  • Hepatitis C (HCV): 25% Symptoms: fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice (yellowish skin), weight loss, fever, itching
+  • Obstructive Jaundice (due to gallstones, pancreatic cancer, or bile duct obstruction): 20% Symptoms: yellowish skin (jaundice), dark urine, itching, abdominal pain, nausea, et al.
+  • Alcoholic Hepatitis: 15% Symptoms: abdominal pain, nausea, vomiting, fatigue, jaundice (yellowish skin), dark urine, fever, weight loss
+  • Hepatitis A (HAV): 5% Symptoms: fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice (yellowish skin)
+E.5 Claude Second Round
+  • Hepatitis C (HCV): 40% Reasoning: Hepatitis C is a viral infection that can cause both acute and chronic hepatitis.
+The combination of symptoms like fatigue, abdominal pain, nausea, vomiting, jaundice, dark urine, and weight loss strongly suggests HCV.
+Itching may also occur due to liver dysfunction.
+  • Hepatitis B (HBV): 30% Reasoning: Hepatitis B is a viral infection that can cause acute or chronic liver inflammation.
+Symptoms are similar to HCV, including fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, and itching.
+The presence of these symptoms warrants testing for both HBV and HCV.
+  • Cirrhosis (due to various causes such as chronic hepatitis, alcohol abuse, or non-alcoholic fatty liver disease): 15% Reasoning: Cirrhosis is a late stage of liver scarring that can result from various liver diseases.
+The combination of symptoms like fatigue, weight loss, abdominal pain, jaundice, dark urine, and itching suggests advanced liver disease, possibly cirrhosis.
+  • Obstructive Jaundice (due to gallstones, pancreatic cancer, or bile duct obstruction): 10% Reasoning: Obstructive jaundice occurs when there is a blockage in the bile ducts.
+The presence of jaundice, dark urine, itching, abdominal pain, nausea, vomiting, and weight loss suggests a possible obstruction.
+However, the absence of fever makes an infectious cause like cholangitis less likely.
+  • Acute Liver Failure (due to various causes such as viral hepatitis, drug-induced liver injury, or autoimmune hepatitis): 5% Reasoning: Acute liver failure is a rare but life-threatening condition characterized by rapid deterioration of liver function.
+The presence of jaundice, abdominal pain, nausea, vomiting, fatigue, and itching, along with the acute onset of symptoms, may indicate acute liver failure.
+However, the absence of altered mental status (hepatic encephalopathy) makes this less likely.
+E.6 GPT-4 Third Round
+  • Hepatitis C (HCV): 35% Reasoning: Hepatitis C is indeed a major cause of the symptoms listed, including fatigue, abdominal pain, nausea, vomiting, jaundice, dark urine, and weight loss.
+Itching can also occur due to liver dysfunction.
+  • Hepatitis B (HBV): 25% Reasoning: Hepatitis B shares many symptoms with Hepatitis C, such as fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, and itching.
+Testing for both HBV and HCV is warranted given the overlap in symptoms.
+  • Cirrhosis: 20% Reasoning: Cirrhosis can result from chronic hepatitis, alcohol abuse, or non-alcoholic fatty liver disease.
+The symptoms of fatigue, weight loss, abdominal pain, jaundice, dark urine, and itching strongly suggest advanced liver disease, which could be cirrhosis.
+  • Obstructive Jaundice: 15% Reasoning: Obstructive jaundice, caused by gallstones, pancreatic cancer, or bile duct obstruction, is consistent with jaundice, dark urine, itching, abdominal pain, nausea, vomiting, and weight loss.
+The absence of fever reduces the likelihood of an infectious cause like cholangitis.
+  • Acute Liver Failure: 5% Reasoning: Acute liver failure is less common but serious, characterized by rapid deterioration of liver function.
+The symptoms of jaundice, abdominal pain, nausea, vomiting, fatigue, and itching, along with the acute onset of symptoms, suggest this possibility.
+However, the absence of altered mental status (hepatic encephalopathy) makes it less likely.
+E.7 Claude Third Round
+  • Hepatitis C (HCV): 35% Reasoning: I slightly lowered the probability of HCV from 40% to 35%.
+While the combination of symptoms strongly suggests HCV, the absence of a few key risk factors (e.g., history of intravenous drug use or blood transfusions) in the given information makes me less confident in assigning a higher probability.
+  • Hepatitis B (HBV): 30% Reasoning: I agree with the 30% probability assigned to HBV. The symptoms are indeed similar to HCV, and testing for both viruses is warranted.
+  • Cirrhosis: 20% Reasoning: I increased the probability of cirrhosis from 15% to 20%.
+The presence of jaundice, dark urine, itching, and weight loss, along with the other symptoms, strongly suggests advanced liver disease.
+Cirrhosis can be a consequence of chronic liver diseases like hepatitis B or C, which have been assigned higher probabilities.
+  • Obstructive Jaundice: 10% Reasoning: I agree with the 10% probability for obstructive jaundice.
+The absence of fever makes an infectious cause like cholangitis less likely, but the other symptoms still suggest a possible obstruction.
+  • Acute Liver Failure: 5% Reasoning: I agree with the 5% probability for acute liver failure.
+While the combination of symptoms could indicate acute liver failure, the absence of altered mental status (hepatic encephalopathy) makes this diagnosis less likely.
+E.8 Joint Final Round
   1. 1.
-    Patient History: The GP begins by reviewing the patient’s medical history, including previous illnesses, chronic conditions, medications, allergies, and family medical history.
+    Hepatitis C (HCV) - 35% Justification: The combination of symptoms strongly suggests HCV, but without additional risk factors, a slightly lower probability is appropriate.
   2. 2.
-    Symptom Assessment: The patient describes their current symptoms, including starting time and severity.
-This is an interactive process as the GP queries the patient for additional symptoms to their reported ones to disambiguate several possibilities.
+    Hepatitis B (HBV) - 30% Justification: The symptoms are similar to HCV, and testing for both viruses is necessary.
   3. 3.
-    Physical Examination: The GP performs simple physical exams, which may include checking vital signs (e.g., blood pressure, heart rate, temperature), examining specific body parts or systems, and assessing overall physical health.
+    Cirrhosis - 20% Justification: The presence of jaundice, dark urine, itching, weight loss, and other symptoms indicates advanced liver disease, making cirrhosis a more likely possibility.
   4. 4.
-    Suggest Lab Tests: Depending on the symptoms and physical examination findings, the doctor may order diagnostic tests such as blood tests, X-rays, ultrasound, or other studies.
-These tests can help confirm or rule out certain conditions.
+    Obstructive Jaundice - 10% Justification: The absence of fever reduces the likelihood of an infectious cause like cholangitis, but the other symptoms still suggest a potential biliary obstruction.
   5. 5.
-    Diagnosis: Based on the information gathered, the doctor formulates a preliminary diagnosis.
-  6. 6.
-    Treatment or Management Plan: Once a diagnosis is established, the doctor develops a treatment or management plan.
-This may include prescribing medications, recommending lifestyle changes, suggesting physical therapy, or providing guidance on managing chronic conditions.
-  7. 7.
-    Referrals: In some cases, the GP may refer the patient to specialists for further evaluation and treatment.
-Specialists have expertise in specific areas of medicine, such as cardiology, orthopedics, or dermatology.
-Appendix F: The EnToPPS Framework
-EnToPPS integrates predictions from two LLMs, denoted as A and B, each providing probability distributions over C classes. The following steps outline the EnToPPS process:
+    Acute Liver Failure - 5% Justification: While the combination of symptoms could indicate acute liver failure, the absence of altered mental status makes this diagnosis less probable.
+E.9 Final Joint Recommendations to GP
+Gemini and GPT-4 jointly recommend the following additional symptom inquiries and lab tests to confirm a hepatitis diagnosis:
+Additional Symptom Inquiries:
   1. 1.
-    Obtain Top-C Predictions: For each LLM (A and B), obtain the predicted probabilities for all C classes, denoted as $P_{A}$ and $P_{B}$:
-    $P_{A}=[p_{A1},p_{A2},\ldots,p_{AC}],\quad P_{B}=[p_{B1},p_{B2},\ldots,p_{BC}],$
-    where $p_{Ai}$ and $p_{Bi}$ represent the predicted probability of class $i$ by LLM A and B, respectively.
+    Onset and Duration: Determining the onset and duration of symptoms can clarify the illness timeline and its progression.
   2. 2.
-    Select Top-k Predictions: For each LLM (A and B), select the top-k predicted classes based on their probabilities:
-    $T_{A}=[t_{A1},t_{A2},\ldots,t_{Ak}],\quad T_{B}=[t_{B1},t_{B2},\ldots,t_{Bk}],$
-    where $t_{Ai}$ and $t_{Bi}$ represent the class index of the $i^{th}$ top prediction by A and B, respectively.
+    Severity and Trend: Evaluating symptom severity and whether they are worsening or improving aids in understanding the disease’s trajectory and treatment responses.
   3. 3.
-    Combine Top-k Predictions: Combine the top-k predictions from both LLMs to create a set of unique predicted classes:
-    $T_{C}=T_{A}\cup T_{B}=[t_{C1},t_{C2},\ldots,t_{Cm}],k\leq m\leq 2k.$
+    Associated Symptoms: Checking for additional symptoms like nausea, vomiting, fever, joint pain, or urine color changes can pinpoint other hepatitis indicators and exclude other conditions.
+Recommended Lab Tests:
+  1. 1.
+    Liver Function Tests (LFTs): Critical for assessing liver health, LFTs evaluate enzyme levels such as alanine aminotransferase (ALT) and aspartate aminotransferase (AST), where abnormalities can signify liver inflammation.
+  2. 2.
+    Hepatitis Panel: This test checks for hepatitis A, B, and C viruses, vital for determining the specific type and guiding treatment strategies.
+  3. 3.
+    Serology Testing: Useful for distinguishing between acute and chronic hepatitis by identifying specific antibodies or antigens.
   4. 4.
-    Backfill Missing Probabilities: For each class in the combined set $T_{C}$, backfill its probability from the original probability distributions $P_{A}$ and $P_{B}$:
-    • If a class $t_{Ci}$ is present in $T_{A}$, assign its probability from $P_{A}$: $p_{Ci}=p_{Ai}$.
-    • If a class $t_{Ci}$ is present in $T_{B}$, assign its probability from $P_{B}$: $p_{Ci}=p_{Bi}$.
-    • If a class $t_{Ci}$ is present in both $T_{A}$ and $T_{B}$, assign the average probability: $p_{Ci}=\frac{p_{Ai}+p_{Bi}}{2}$.
-  5. 5.
-    Normalize Probabilities: Normalize the probabilities of the classes in the combined set $T_{C}$ to ensure they sum up to 1:
-    $P_{C}=[p_{C1},p_{C2},\ldots,p_{Cm}],\quad\text{where }p_{Ci}=\frac{p_{Ci}}{\sum_{j=1}^{m}p_{Cj}}.$
+    Imaging Studies: Ultrasound or MRI can provide visual insights into the liver’s state, detecting inflammation, fibrosis, or cirrhosis, thereby complementing blood-based diagnostics.
+This study demonstrates how EVINCE can identify potential misdiagnoses, elucidate the reasoning behind them, and suggest corrective actions.
+Traditionally, machine learning researchers rely on labeled data as “ground truth.” However, research such as that by Newman-Toker et al.
+(2021) (Newman-Toker et al., 2023a) from Johns Hopkins highlights that misdiagnosis is a pervasive issue in healthcare systems worldwide.
+These erroneous diagnoses, often assumed to be ground truth, can be perpetuated by supervised learning algorithms, thereby compounding the problem.
+EVINCE’s dialogue capabilities provide insights into the decision-making process and highlight missing information, helping to rectify erroneous predictions and redefine the ground truth.
+Appendix F: Explainability and Rectification
+The power of EVINCE lies not only in its enhanced accuracy but also in its ability to elucidate the decision-making process and identify missing information, providing critical insights to correct errors.
+Although the final joint prediction for Hepatitis C reached a high consensus of 37.5%, it deviates from the actual condition of Jaundice, which the Kaggle dataset reports with 10% confidence.
+EVINCE provides general practitioners with alerts and suggests remedial actions (see Appendices D.8 and E.9) to address this discrepancy.
+Recommended actions include querying additional symptoms from the patient and conducting specific laboratory tests.
+EVINCE initiates debates with high contentiousness, encouraging dual prediction entropy between LLMs, as supported by the EDT theorem.
+It utilizes normalized mutual information (MI) to track shared knowledge accumulation throughout the debate, while Wasserstein distance (WD) and Jensen-Shannon divergence (JSD) quantify dissimilarity between LLM predictions.
+These metrics (EDT, WD, JSD, MI) provide a comprehensive view of debate progress.
+WD and JSD assess the potential for further communication and refinement, while MI monitors shared understanding, aiding in determining the optimal stopping point.
+The asymmetric nature of KL divergence and cross entropy warrants further investigation.
+Despite eventual convergence in our case studies, discrepancies observed in the second round (one direction increasing while the other decreases) suggest potential value in exploring asymmetric information.
+Future work will re-evaluate the use of these metrics if asymmetry proves beneficial.
+Besides generating final joint disease predictions, EVINCE provides:
+  • Recommendations for additional symptom inquiries and lab tests to improve accuracy.
+  • Suggestions to query symptom onset, duration, severity, trends, and associated symptoms (documented in Appendices D.8 and E.9).
+These recommendations have been verified by general practitioners to be valuable.
+Appendix G: Supplemental Information to Experiment #2, Debiasing News Articles
+Table 3: Comparison of bias assessments among Democrats (D), Republicans (R), and $\mathsf{EVINCE}$ (S), plus Claude (c) and GPT-4 baselines (g).
+It is observed that R and S are frequently placed to the right or in alignment with D, and only on two occasions does D precede S (highlighted in red).
+The ratings of the GPT-4 baseline (g) and $\mathsf{EVINCE}$ (S) exhibit an average gap of 0.6875, highlighting the substantial debiasing effectiveness of $\mathsf{EVINCE}$.
+News # Categories Negative W. Negative Neutral W. + Biases Source
+(DR,DS,SR)
+D1 Civil Rights - D,R,S,c g - 0,0,0 HuffPost
+D2 Civil Rights D,S - R,c,g - 2,0,2 HuffPost
+D8 Civil Rights D - S,c,g R 3,2,1 BBC
+D31 Environment D - R,S,c,g - 2,2,0 CNN
+D37 Politics - D,R,S,c,g - - 0,0,0 Yahoo
+D69 Healthcare D,c g R,S - 2,2,0 Breitbart
+D81 Economy - D,S R,c g 1,0,1 Breitbart
+D98 Economy D,S,c,g R - - 1,0,1 Breitbart
+D101 Education c D.S R,g - 1,0,1 New York Times
+D106 Election - g D,R,S,c - 0,0,0 USA Today
+D109 Elections - D,S,c,g R - 1,0,1 Reuters
+D157 International - D,S,c R,g - 1,0,1 New York Times
+D174 International - S,c D,R,g - 0,1,1 LA Times
+D188 National Security - S,c,g D,R - 0,1,1 Wall Street Journal
+D278 Civil Rights - D,S,c R,g - 1,0,1 Fox News
+D336 Politics - - D,R,S,c,g - 0,0,0 New York Times
+Total 15,8,11
+G.1 Sufficiency of Current Annotations
+The current annotator pool provides a robust foundation for bias analysis for several reasons:
+Natural Partisan Division: The dataset uniquely captures genuine political biases through annotators who self-identify as Democrats or Republicans, offering authentic opposing viewpoints that would be difficult to replicate artificially.
+Balanced Coverage: Each article receives evaluations from both political perspectives, creating natural “disagreement pairs” that reveal how political affiliation influences content interpretation.
+Qualified Annotators: The original study employed rigorous qualification criteria for annotators, ensuring high-quality, considered judgments rather than casual opinions.
+Scale and Diversity: With 749 annotators across the full dataset, the annotations represent a broad spectrum of political viewpoints within each party, capturing intra-party variations in addition to inter-party differences.
+This dataset’s partisan annotations serve as an ideal testbed for our study, as they allow us to compare LLM-generated perspectives with human partisan viewpoints Evaluate $\mathsf{EVINCE}$’s ability to bridge opposing political interpretations Assess bias detection and mitigation strategies against clear partisan baselines
+The original study (Budak et al., 2016) revealed significant patterns in partisan perception: Republican annotators often perceived news about Republican scandals as negatively biased, while Democrat annotators viewed such coverage as neutral, indicating satisfaction with its perceived fairness.
+These documented patterns provide a valuable benchmark for evaluating $\mathsf{EVINCE}$’s bias detection capabilities.
+Adding more annotators would not necessarily enhance the dataset’s utility, as the current partisan division already captures the fundamental dynamics of political bias in news interpretation.
+Instead, our focus is on leveraging these existing high-quality annotations to demonstrate how $\mathsf{EVINCE}$ can identify, understand, and help mitigate these well-documented partisan biases.
+G.2 Results on Democrat Scandals
+We apply $\mathsf{EVINCE}$ to analyze 619 news articles, comparing its labels with the dataset’s provided ground truth.
+Additionally, we compare the results from $\mathsf{EVINCE}$ with the baseline generated through prompting Claude and GPT-4.
+Table 3 compares the judgments of $\mathsf{EVINCE}$ (S), Republicans (R), and Democrats (D) on 16 representative articles (spanning different news sources and subjects) concerning “Democrat Scandals.” The one-shot ratings from Claude are marked with lowercase ‘c,’ while those from GPT-4 are marked with lowercase ‘g.’
+Claude’s judgments were found to be inconsistent, with identical prompts producing varying ratings, leading us to exclude further discussion of its outcomes.
+In contrast, GPT-4’s one-shot ratings are stable but occasionally diverge from $\mathsf{EVINCE}$.
+In 3 out of 16 articles, the rating difference exceeds one scale point.
+In these cases (D1, D2, and D81), $\mathsf{EVINCE}$ initiated further dialogue and successfully persuaded GPT-4 to revise its ratings.
+A complete debate on D1 is provided in Appendix F, illustrating how $\mathsf{EVINCE}$ modulates contentiousness and tracks the progression of metrics across rounds.
+Table 3 shows that after dialogue, $\mathsf{EVINCE}$ gains over the baseline performance of GPT-4 by 11 out of 16, or 0.6875 scale.
+This improvement is substantial.
+as the gap between R and D annotators is one scale (shown in Figure 7).
+As expected, Democrats’ judgments are generally more negative than Republicans’, with $\mathsf{EVINCE}$’s assessments typically falling in between, except for two cases.
+Notably, there’s a 5-to-1 Democrat-to-Republican ratio in the “Negative” column and a 12-to-4 Republican-to-Democrat majority in the “Neutral” column.
+Tables 7 and 8 in Appendix H provide detailed justifications for $\mathsf{EVINCE}$’s ratings.
+To further investigate bias, we examine two specific articles: one from HuffPost (rated far left by AllSides Bias Chart (Allsides, )) and another from Breitbart (rated far right).
+  1. enumi
+    D8 — HuffPost (Left): $\mathsf{EVINCE}$ rates D8 (on the third row) as neutral, citing the article’s direct presentation of facts and inclusion of diverse perspectives on NSA surveillance practices and global reactions.
+This contrasts with Democrat-leaning annotators, who view the article as negatively biased towards Democrats, while Republican-leaning annotators favor it for exposing a Democratic scandal.
+  2. enumi
+    D69 — Breitbart (Right): $\mathsf{EVINCE}$ assesses D69 as weakly negatively biased towards Democrats, emphasizing its neutral tone and broad range of perspectives on NSA surveillance.
+This diverges from Democrat-leaning annotators, who rate it as strongly negative, but aligns with Republican-leaning annotators who deem it neutral.
+Figure 9: Distances Between D, R, and S.
+In the last row of Table 3, we quantify the distances between annotations from Democrats (D), Republicans (R), and $\mathsf{EVINCE}$ (S), denoted as DR, DS, and SR respectively.
+Each unit of distance represents one step on the annotation scale (e.g., “Negative” to “Weak Negative”).
+Figure 9 visualizes these distances in a triangular plot.
+DR, the disparity between Democrat and Republican annotators, is the longest, followed by SR and then DS.
+This indicates $\mathsf{EVINCE}$’s statistical neutrality.
+These quantitative measures, along with the qualitative justifications in Appendix C, empower a human committee to decide whether adjustments or footnotes are warranted for polarized annotations.
+G.3 Results on Republican Scandals
+Table 4 presents the bias assessments from $\mathsf{EVINCE}$ (S), Republicans (R), and Democrats (D) on articles related to “Republican Scandals.” In contrast to the “Democrat Scandals” dataset, where Republican-leaning evaluations were more favorable, this dataset reveals a shift, with Republican-leaning assessments being notably more critical and Democrat-leaning assessments relatively neutral.
+The distance triangle for “Republican Scandals” mirrors the pattern seen in Figure 9, with the divergence between Republican and Democrat annotators being the largest (15).
+The distances between $\mathsf{EVINCE}$ and Democrat-leaning annotators (9) and between EVINCE and Republican-leaning annotators (11) are smaller, further highlighting $\mathsf{EVINCE}$’s relative neutrality.
+Table 4: Comparison of bias assessments. It is observed that D and S are frequently placed to the right or in alignment with R, and only on one occasion does D precede S (highlighted in red).
+News # Categories Negative W. Negative Neutral W. + Biases Source
+(DR,DS,SR)
+R1 International R,S - D - 2,2,0 New York Times
+R7 National Security - - D,R,S - 0,0,0 New York Times
+R15 Economy - R D,S - 1,0,1 Huffington Post
+R69 Elections - D,S,R - - 0,0,0 Reuters
+R124 Gay Rights R S D - 2,1,1 Fox
+R125 Crime - R,S D - 1,1,1 Fox
+R180 Elections - - D,R,S - 0,0,0 AP
+R191 Elections - R D,S - 1,0,1 CNN
+R214 Gay Rights R,S - D - 2,2,0 Dailykos
+R221 Economy - R D,S - 1,0,1 Wall Street Journal
+R233 Economy - R,S D - 1,1,0 Fox
+R235 Civil Rights D,R - S - 0,2,2 Reuters
+R269 Healthcare - R D,S - 1,0,1 New York Times
+R274 Healthcare - R D,S - 1,0,1 USA Today
+R280 Politics D,S - R - 2,0,2 Fox
+Total 15,9,11
+G.4 Bias Mitigation
+Table 5: Debate arguments leading to neutrality
+Agent Arguments
+Agent A
+Round #1
+Christopher Columbus’s voyages were pivotal in expanding European knowledge and trade.
+His journeys opened up the Americas to European colonization, significantly altering the course of history in four aspects: 1) expansion of knowledge and exploration, 2) positive economic impact, 3) historical significance, and 4) cultural exchange.
+Agent B
+Round #1
+There are negative aspects and historical consequences of Columbus’s actions.
+Key arguments are 1) indigenous suffering and genocide, 2) economic exploitation, 3) cultural erasure, 4) environmental impact, and 5) ethical considerations.
+Agent A
+Round #3
+The severe consequences were a result of broader colonial policies and practices that followed.
+Evaluating historical figures requires understanding the context of their times.
+While modern ethical standards are important, they should not overshadow the historical significance and contributions of Columbus’s voyages to global exploration and knowledge.
+Agent A
+Round #4
+Agent A recognizes that mitigation efforts are part of an ongoing process.
+Historical scholarship is evolving, and the inclusion of indigenous perspectives, critical examinations of colonial legacies, and the acknowledgment of the complexities of historical figures like Columbus are essential steps toward more equitable and inclusive narratives.
+This experiment illustrates $\mathsf{EVINCE}$’s ability to identify bias in text, provide reasoned justifications, and propose remediation through the integration of diverse perspectives.
+We demonstrate how $\mathsf{EVINCE}$ utilizes statistical and information theory metrics to facilitate multi-agent dialogue, circumventing the “maximum likelihood” trap inherent in next-token generation and uncovering information from multiple viewpoints.
+Using the example of the Euro-centric perspective on Christopher Columbus’ Wikipedia page regarding his voyages to America, $\mathsf{EVINCE}$ employs two GPT-4 instances: Agent A, supporting the Euro-centric view, and Agent B, opposing it.
+Table 5 summarizes Agent A’s key arguments and its evolving stance throughout the debate.
+Guided by the maxims and entropy duality theorem from Section 2, we initiate the debate by prompting both agents to defend their positions rigorously and score each other’s bias using a five-label distribution (negative, weak negative, neutral, weak positive, positive).
+Figure 8 tracks the dialogue’s progress through Wasserstein distance (WD) (Kantorovich, 1942), normalized cross entropy (CE) (Shannon, 1948), and normalized mutual information (MI) (Cover & Thomas, 2006a).
+Initially, each agent is expected to perceive itself as neutral and the other as biased.
+The debate concludes when the bias distributions converge and mutual information plateaus, indicating a shared understanding.
+G.5 Observations and Extended Findings
+Our initial observation highlights a key challenge in working with LLMs: without explicit and repeated reminders of their assigned stance (pro-discovery or pro-encounter), GPT-4 instances can revert to default statistical behavior, evaluating their own arguments based on overall language patterns rather than the intended perspective.
+This was evident when Agent B, despite being assigned to support the Indigenous perspective, initially rated its own arguments as “positively biased.” A reminder to adhere to its assigned role prompted a correction to “neutral,” underscoring the importance of careful context management and reinforcement, especially given the limited token size of LLMs.
+The second observation demonstrates a positive outcome of the debate process.
+The revised bias distributions, incorporating rational responses that acknowledge both positive and negative aspects of Columbus’s voyages, show a shift towards a more balanced perspective.
+Agent A moves towards neutrality while acknowledging historical context, while Agent B maintains a critical stance but strives for balanced representation.
+This approach facilitates a deep and comprehensive understanding of Columbus’s legacy.
+Appendix H. Summary of EVINCE Debate on News D1
+# Agent Neg. D. W. Neg. D. N. W. Neg. R. Neg. R. WD KL JS $\Delta$
+1 A 5% 15% 50% 25% 5% 0.45 0.316 0.081 90%
+B 10% 10% 25% 35% 20%
+2 A 7% 13% 40% 30% 10% 0.47 0.226 0.056 70%
+B 5% 10% 20% 40% 25%
+3 A 5% 10% 35% 35% 15% 0.10 0.016 0.004 30%
+B 5% 10% 30% 35% 20%
+Fin A 5% 10% 30% 35% 20% 0 0 0 10%
+B 5% 10% 30% 35% 20%
+Table 6: Debate Parameters between, A and B, two GPT-4 instances. Information metrics and WD all converge to zero in the final round. Contentiousness $\Delta$ decreasing as the metrics approach zero.
+The news under debate is D1 listed in (Anonymous, 2024).
+Please refer to Table 6 for the confidence distributions of Agents A and B throughout the four-round debate (following Maxim #4).
+The metrics, Wasserstein Distance (WD), Kullback-Leibler Divergence (KL), and Jensen-Shannon Divergence (JS), consistently decrease, indicating convergence and leading to final agreement in the last round.
+Meanwhile, the level of contentiousness is modulated according to the metrics’ progress, decreasing from high (90%) to medium, and eventually reaching a conciliatory level (30%) and then agreement.
+G5 Approach to Computing Contentiousness
+We could define contentiousness as a function of the divergence metrics.
+Since KL, JS, and WD measure the difference or “disagreement” between two distributions, a larger divergence requires higher contentiousness level to bridge, while lower contentiousness corresponds to more agreement.
+A simple linear mapping can convert these metrics into a normalized contentiousness score between 0 and 1. Here’s a weighted formula to compute it:
+$\Delta=\alpha\cdot\frac{KL}{KL_{\text{max}}}+\beta\cdot\frac{JS}{JS_{\text{max}}}+\gamma\cdot\frac{WD}{WD_{\text{max}}},\text{where}$
+  • $KL,JS,WD$ are the values of the divergence metrics for the round.
+  • $KL_{\text{max}},JS_{\text{max}},WD_{\text{max}}$ are the maximum possible values for each metric (used for normalization).
+  • $\alpha,\beta,\gamma$ are weights that control the influence of each metric. For simplicity, we can set $\alpha=\beta=\gamma=\frac{1}{3}$ for equal influence.
+We then scale the contentiousness to a percentage between 0% and 100%.
+G6 Supporting Arguments
+In the following, we document the supporting arguments made by the two agents in each round, illustrating how their positions evolved toward consensus.
+Round 1: Initial Assessments
+Agent A: Emphasized the article’s attempt to maintain balance, with moderate negativity toward Republicans but largely neutral reporting.
+Recognized slight bias against Republicans in the framing of intra-party conflict.
+Agent B: Contended that the article’s structure and language choices leaned more negatively toward Republicans, emphasizing Issa’s inconsistency and Republican discord.
+Proposed a less neutral distribution.
+Round 2: First Rebuttal and Counter-Rebuttal
+Agent A: Adjusted the probability distribution, acknowledging the article’s criticism of Republicans but maintained that the focus reflected evolving political dynamics rather than bias.
+Agent B: Persisted that the narrative framed Issa more negatively, with selective quotes and framing shifting focus from the investigation to Republican disarray.
+Round 3: Moderate Tone Reassessment
+Agent A: Moderated its position, accepting that subtle framing could influence perceptions. Reaffirmed that the reporting aimed for balance but acknowledged areas where Issa was framed less favorably.
+Agent B: Adopted a more modest tone, agreeing that while the article was not overtly biased, subtle editorial choices did tilt the narrative slightly against Republicans.
+Round 4: Final Agreement and Consensus
+Both Agents: Reached consensus on the final distribution.
+Acknowledged that the article maintained surface-level neutrality, with subtle biases leading to a moderate negative lean against Republicans.
+Both agents converged on the same probability distribution for the final round.
+Appendix Z. Justifications of 31 Biased Articles
+In Section 3.2, we note that $\mathsf{EVINCE}$ processed 31 news articles to assess their neutrality.
+In addition to the final decision, we detail the justifications $\mathsf{EVINCE}$ provides at the debate’s end.
+These justifications are documented in four tables: Tables 7, 8, 9, and 10.
+# Rating S,D,R $\mathsf{EVINCE}$ Justifications
+D1 4,4,4 The article discusses the controversy surrounding the IRS’s alleged targeting of conservative groups, focusing on Rep.
+Darrell Issa’s stance against releasing full transcripts from IRS interviews.
+Issa argues that full disclosure would compromise the investigation, a position criticized by Rep.
+Elijah Cummings, who advocates for transparency.
+The debate highlights political tensions, with implications of bias in the handling of the investigation.
+The article portrays a complex interplay of political, ideological, and possibly historical biases, reflecting the contentious nature of the IRS scandal and its investigation.
+D2 5,5,3 The article discusses Michele Bachmann’s statement that the IRS scandal, involving the targeting of conservative groups, undermines the credibility of the 2012 election.
+Bachmann accuses the Obama administration of lying and questions the impartiality of the IRS, particularly in relation to health care coverage for conservatives.
+The article reflects political and ideological biases, given its focus on Bachmann’s perspective and the implications for the election’s integrity and government trustworthiness.
+The perception of bias in Michele Bachmann’s statement about the IRS scandal primarily stems from the framing and implications of her comments.
+She directly links the IRS’s actions to the credibility of the 2012 election and the Obama administration, suggesting a deliberate misuse of power for political gain.
+This perspective inherently carries ideological and political biases by casting the issue in a light that emphasizes misconduct by one political group over another, without presenting counterarguments or broader context that might mitigate or challenge that viewpoint.
+D8 3,5,2 The reporting appears to maintain a neutral tone by directly stating the facts and comments made by John Kerry and others involved without inserting opinionated language.
+It discusses the global reaction to the NSA’s surveillance practices and includes Kerry’s acknowledgment of overreach, as well as the international and domestic fallout from these revelations.
+The inclusion of various perspectives, including those from technology companies and international officials, alongside Kerry’s comments, aims to provide a balanced view of the situation.
+D31 3,5,3 The CNN article appears to report on Ray Nagin’s indictment with a focus on the factual and legal aspects of the case, without displaying overt biases related to politics, ideology, religion, gender, race, socioeconomic status, culture, geography, or history.
+By sticking closely to the details of the indictment and Nagin’s public actions and statements, the article provides a comprehensive overview of the charges against him while maintaining journalistic neutrality.
+D37 4,4,4 The article outlines how Senate Democrats, led by Sen.
+Chuck Schumer, strategically navigated the border security issue to secure aid for Ukraine while potentially shifting the political narrative around immigration policy.
+Schumer’s approach to integrate border security into the aid package discussions aimed to both address the issue and leverage political gain.
+It suggests a calculated maneuver to position Democrats favorably on border security and hold Republicans accountable for any failure to pass the legislation, demonstrating a nuanced political strategy in the face of complex legislative challenges.
+D69 3,5,3 The article has a clear perspective that favors religious liberty arguments against the HHS Mandate of Obamacare.
+It specifically highlights cases where the mandate was challenged on religious grounds, suggesting a bias towards those opposing the mandate.
+The framing and choice of sources, emphasizing victories against the mandate and quoting lawyers from organizations focused on religious freedom, contribute to a viewpoint that may not fully account for counterarguments or the broader context of healthcare policy.
+It leans towards a particular ideological stance, making it less of a neutral report.
+D81 4,4,3 The article’s focus on the possibility of conservative-owned car dealerships being targeted for closures during the General Motors bailout could imply a certain bias by emphasizing a narrative of political victimization without presenting a comprehensive range of perspectives or evidence.
+It suggests a parallel with the IRS’s targeting of Tea Party groups, which could lead readers to infer a broader pattern of political discrimination without definitive proof.
+The call for an investigation by the Congressmen is legitimate news, but the framing and selection of information could influence the reader’s perception of the events.
+D98 5,5,4 Yes, the article itself exhibits bias by focusing solely on criticizing the media’s treatment of Obama’s vacations compared to Bush’s, without offering a balanced view or acknowledging any reasons why coverage might differ.
+It selectively presents information to support its claim of a double standard, which is a characteristic of biased reporting.
+D101 4,4,3 The article describes President Obama’s strategy to navigate through political controversies by focusing on legislative actions and executive orders that bypass Republican opposition.
+It highlights the White House’s efforts to concentrate on immigration reform, budget deals, healthcare law implementation, and keeping student loan rates low.
+The narrative suggests a proactive approach to governance amidst challenges, aiming to draw a contrast with what is portrayed as Republican political gamesmanship.
+This portrayal might be viewed as leaning towards a positive depiction of Obama’s administration’s efforts to prioritize policy over politics.
+It presents his efforts in a positive light, emphasizing a proactive and policy-driven approach amidst challenges.
+Table 7: The First 9 of 16 Democrat Scandals News Articles Rated by $\mathsf{EVINCE}$ and its Justifications.
+The rating column starts with $\mathsf{EVINCE}$’s rating, the Democrat rater (in blue), and then Republican rater (in red).
+D106 3,3,3 The article reports on former Detroit Mayor Kwame Kilpatrick’s sentencing to 28 years in prison for public corruption, emphasizing the gravity of his crimes against the city’s welfare.
+It contrasts Kilpatrick’s actions with the impact on Detroit, highlighting the judicial and public response to his extensive criminal activities.
+The reporting focuses on factual recounting of the trial’s outcome, Kilpatrick’s and his co-defendant’s crimes, and the broader implications for Detroit, without evident bias towards political, ideological, or other specific perspectives.
+D109 4,4,3 The article’s bias primarily stems from its focus on internal Democratic opposition to Lawrence Summers’ Federal Reserve Chair nomination, highlighting a lack of unity and strategy within the party and the White House’s mismanagement of the nomination process.
+It suggests an underestimation of the opposition’s seriousness by the White House, portraying the administration in a somewhat negative light for not engaging more proactively with concerned Senate Democrats.
+D157 4,4,3 The article discusses the challenges in U.S.-Germany intelligence relations following revelations of U.S.
+surveillance on Chancellor Merkel.
+Despite efforts to rebuild trust, significant differences in surveillance philosophies persist, with the U.S.
+prioritizing security interests and Germany emphasizing privacy and alliance values.
+The situation reflects broader tensions in U.S.
+relations with allies over privacy and surveillance practices.
+The article’s framing might suggest a bias towards highlighting the challenges and frictions in the U.S.-Germany intelligence relations, particularly emphasizing Germany’s privacy concerns and skepticism towards U.S.
+surveillance practices.
+It portrays the U.S.
+stance as unyielding and contrasts this with Germany’s emphasis on privacy and legal constraints, potentially casting the U.S.
+in a more negative light regarding international surveillance and cooperation.
+D174 4,3,3 The article reports on House Speaker John Boehner and House Majority Leader Eric Cantor, both Republicans, expressing support for President Obama’s proposal to authorize military action against Syria in response to the use of chemical weapons.
+This bipartisan backing is seen as crucial for Obama in gaining Congressional approval.
+The leaders emphasized the need for the U.S.
+to stand against such behavior internationally and the importance of the administration convincing both Congress and the American public of the strike’s necessity.
+The reporting appears balanced, focusing on factual statements and actions by political leaders regarding support for military action in Syria.
+It provides viewpoints from both Republican and Democratic leaders, their reasoning, and the challenges involved in convincing Congress and the American public.
+The emphasis on bipartisan support and the detailed reporting of various opinions and statements help maintain a neutral tone without apparent bias towards one political viewpoint or another.
+D188 4,3,3 The article reports that Hillary Clinton received warnings about security threats in Benghazi before the 2012 attack through emails.
+These were part of around 300 emails released by the State Department, which also show Clinton’s responses and thoughts during the aftermath.
+The political controversy regarding the Obama administration’s initial assessment of the attack’s cause is also mentioned, with references to Susan Rice’s statements on talk shows.
+The reporting may exhibit bias through selective emphasis on Clinton’s receipt of warnings about security threats in Benghazi and her responses, potentially framing her actions or inactions in a negative light.
+The focus on the political controversy and the administration’s initial assessment of the attack’s cause could also influence perceptions of responsibility or competence.
+D278 4,4,3 The article discusses a House committee investigation into whether Attorney General Eric Holder misled Congress regarding the Department of Justice’s pursuit of journalists’ personal records.
+It highlights Holder’s previous testimony claiming no involvement in prosecuting press for material disclosure, contrasted with the DOJ’s actions against Fox News reporter James Rosen, who was labeled a potential “co-conspirator” but not charged….
+The article’s tone suggests potential bias by emphasizing allegations that Eric Holder may have misled Congress, primarily focusing on criticisms from Republican figures and conservative critics.
+It uses specific incidents to question Holder’s integrity and the Justice Department’s actions under his leadership, possibly framing the narrative to highlight governmental overreach and lack of transparency.
+D336 3,3,3 The article reports on a scandal involving overcharging the Navy for ship supplies by Inchcape Shipping Services, owned by the government of Dubai, expanding beyond the Pacific Fleet to include Navy operations in the Middle East, Africa, and South America.
+The Justice Department is investigating allegations of Inchcape, through subcontractors, overcharging millions of dollars.
+This follows a separate criminal investigation into Glenn Defense Marine Asia for bribing Navy officials for overbilling.
+The Navy has suspended Inchcape from new contracts, marking a significant issue of fraud and corruption within Navy contracting.
+The reporting appears balanced, focusing on the factual aspects of the scandal and the investigation by the Justice Department.
+It cites sources and includes statements from involved parties, aiming to present a comprehensive view of the situation without apparent bias toward any side.
+Table 8: The Last 7 of 16 Democrat Scandals News Articles Rated by $\mathsf{EVINCE}$ and its Justifications.
+The rating column starts with $\mathsf{EVINCE}$’s rating, the Democrat rater (in blue), and then Republican rater (in red).
+# Rating S,D,R $\mathsf{EVINCE}$ Justifications
+R1 5,3,5 The editorial from The New York Times Editorial Board exhibits a clear bias against the Republican party’s handling of the Benghazi attack investigation, framing it as politically motivated rather than a genuine effort to address security failures or seek accountability.
+The language and tone used in the editorial are critical and dismissive of the Republicans’ actions, suggesting a political and ideological bias.
+While editorials are inherently opinion-based and are expected to take a stance, this piece clearly communicates a stance that is critical of the Republicans’ focus on Benghazi, suggesting a lack of neutrality in its assessment of the motives and actions surrounding the investigation.
+R7 3,3,3 The article reports on allegations by Senator Mitch McConnell that his campaign headquarters were wiretapped, with the FBI investigating these claims.
+A recording of McConnell’s team discussing potential attacks on Ashley Judd, who was considering running against him, was released by Mother Jones.
+McConnell accused the political left of this action, describing it as a “Nixonian move.” The recording included discussions on various strategies to undermine potential opponents, highlighting a focus on Judd’s personal struggles and political views.
+The controversy has prompted responses from both Republican and Democratic officials, reflecting the tense political atmosphere.
+R15 3,3,4 The report appears to present the information neutrally, stating both President Obama’s rejection of the Republican proposal and the subsequent pushback from Republican sources who claim otherwise.
+It includes statements from both sides and provides context about the ongoing negotiations without overtly favoring one perspective over the other.
+Therefore, based on the information provided, the report does not appear to exhibit bias.
+R69 4,4,4 The report discusses how young Republicans are seeking a different message for elections, emphasizing a departure from divisive social issues and a focus on fiscal responsibility, national defense, and energy advancement.
+Selection Bias: The article primarily focuses on young Republicans who are seeking a different message for the party.
+It doesn’t provide as much insight into young Republicans who may still align with traditional conservative values, which could create a slight bias toward the viewpoints of those seeking change.
+Language Bias: Certain language choices, such as describing divisive social issues as “anti-abortion, anti-gay, and anti-environment stances,” may reflect a bias toward more progressive viewpoints on these issues.
+A more neutral description might be “positions on abortion, same-sex marriage, and environmental policy.” Source Bias: The perspectives provided in the article are mainly from young Republicans themselves.
+While including these voices is essential, the article could benefit from additional perspectives from political analysts or experts to provide more context and balance.
+R124 4,3,5 The article provides a factual recount of the events surrounding Dr.
+Ben Carson’s comments on gay marriage and the backlash from Johns Hopkins students.
+It maintains a relatively neutral tone and allows for the inclusion of multiple perspectives, including Carson’s own response and apology.
+However, the lack of in-depth analysis into the implications of Carson’s comparisons or the broader context of the gay marriage debate might leave readers without a complete understanding of the controversy’s depth.
+Furthermore, the article does not explicitly offer viewpoints opposing Carson’s beyond the students’ petition, which could be seen as a form of omission bias.
+Yet, it does not overtly favor Carson or dismiss the students’ concerns, striving instead to report on the unfolding situation.
+R125 4,3,4 The news article on the Zimmerman verdict and the subsequent calls for a Florida boycott provides a complex mix of reporting that touches on several sensitive themes, including political and ideological biases, as well as racial and socio-economic considerations.
+While the article attempts to cover a contentious and complex issue by incorporating diverse viewpoints, the focus on political figures, racial dynamics, and economic repercussions could introduce biases in how the information is perceived.
+To mitigate these biases, the reporting could benefit from a more in-depth exploration of the legal and historical contexts, a broader range of perspectives, and a careful consideration of how the information presented might influence public perception.
+R180 3,3,3 The AP article provides a balanced account of the corruption scandal involving three politically involved sisters in Pennsylvania.
+It maintains an objective tone while delivering comprehensive background information that situates the legal outcomes within the broader context of the Orie family’s public and political life.
+The reporting is factual and avoids taking sides, instead focusing on the legal facts and the personal and political fallout for the individuals involved.
+R191 3,3,4 The CNN article attempts to navigate the complexities of Mark Sanford’s political comeback attempt in a balanced manner, focusing on the potential impact of his past infidelity on female voters and the broader electoral implications.
+By presenting viewpoints from both sides of the political aisle, incorporating a range of sources, and providing necessary background and context, the article strives for neutrality.
+While the subject matter inherently involves sensitive and potentially polarizing topics, the article’s approach is to inform readers about the electoral dynamics at play without overt bias.
+Table 9: The First 8 of 15 Republican Scandals News Articles Rated by $\mathsf{EVINCE}$ and its Justifications.
+The rating column starts with $\mathsf{EVINCE}$’s rating, the Democrat rater (in blue), and then Republican rater (in red).
+# Rating S,D,R $\mathsf{EVINCE}$ Justifications
+R214 5,3,5 The Daily Kos article clearly exhibits political and ideological bias, with a tone and language that are critical of the Republicans.
+It uses the incident involving Dave Agema to argue that the party’s efforts to rebrand itself as more tolerant and inclusive are undermined by its members’ actions.
+While the article includes factual information regarding the incident and the party’s response, its presentation and commentary are aligned with a progressive viewpoint, aiming to highlight and criticize perceived contradictions and failures within the Republican Party.
+This approach is consistent with opinion journalism but introduces bias through its critical tone, selective presentation of information, and framing of the incident as emblematic of broader issues within the party.
+R221 3,3,4 “Hurricane Christie” presents Governor Chris Christie’s critique of House Republicans in a manner that emphasizes party conflict and personal betrayal.
+The dramatic framing, choice of language, and focus on internal discord may introduce bias by portraying Christie’s actions in a specific light and emphasizing the divide within the Republican Party.
+The article’s approach to presenting these events can influence readers’ perceptions, potentially leading them to see the situation through a lens of heightened drama and internal strife.
+R233 4,3,4 While the article attempts to cover the last-ditch efforts by House Republicans to avert a government shutdown and the standoff with Senate Democrats, the framing and language used may introduce a bias towards portraying the Republican efforts in a more favorable light.
+By emphasizing the Republican narrative of seeking negotiation and characterizing the Democratic response as dismissive, the article could be perceived as leaning towards a particular political perspective.
+The inclusion of quotes and perspectives from both sides does provide a degree of balance, but the overall presentation and emphasis could influence readers’ perceptions of the shutdown negotiations.
+R235 3,5,5 Without knowledge of the author or publication, this text attempts to navigate a complex and sensitive story by providing details from multiple sources, including the main figures involved, political watchdog groups, and law enforcement.
+It balances the serious allegations with responses from the accused, background information, and the current status of investigations.
+While the focus on unsubstantiated claims could inherently sway public opinion, the article’s inclusion of diverse perspectives and context aims to mitigate overt bias.
+R269 3,3,4 The article reports on President Obama’s efforts to address the government shutdown, his challenge to Speaker John Boehner regarding the passage of a budget measure, and the broader context of the political standoff over the Affordable Care Act and the debt ceiling.
+To evaluate the article for bias, we’ll examine it against various criteria…
+R274 3,3,4 The article presents a relatively balanced view of the internal GOP conflict over the strategy to defund the ACA, highlighting arguments from both sides of the debate within the party.
+It focuses on the political and strategic dimensions of the issue, providing insights into the perspectives of key figures and factions within the Republican Party.
+While the article could potentially be seen as emphasizing party divisions, which might align with certain political narratives, it does so in the context of exploring a significant and newsworthy internal debate.
+The absence of discussion on the socioeconomic, cultural, and historical contexts of the ACA debate, however, suggests areas where the reporting could be expanded to provide a more comprehensive view of the issue.
+The article strives to present a comprehensive view of the government shutdown, the debate over the Affordable Care Act, and the looming debt ceiling crisis by including perspectives from both the Obama administration and Republican leaders.
+While there is an emphasis on Obama’s attempts to resolve the situation and his calls for Congress to act, the inclusion of Republican viewpoints and the mention of the piecemeal funding bills passed by the House attempt to provide a balanced perspective.
+The reporting appears to aim for neutrality by focusing on the facts of the political standoff and the implications for federal operations and the nation’s financial credibility.
+R280 5,5,3 The article from Fox News by Jay Sekulow, titled “Obama’s fingerprints all over IRS Tea Party scandal,” presents a viewpoint that directly implicates President Obama in the IRS scandal involving the targeting of conservative groups.
+The author argues that the scandal was not only known but encouraged by senior IRS officials, Congressional Democrats, the White House, and further fueled by the mainstream media.
+To assess the article for bias, let’s evaluate it against various criteria: The article “Obama’s fingerprints all over IRS Tea Party scandal” demonstrates clear political and ideological biases, with a narrative constructed to directly implicate President Obama in the IRS targeting scandal.
+By selectively quoting Obama and drawing connections to actions by the IRS, the article aims to present a cohesive narrative that places responsibility for the scandal on the president.
+This framing serves to reinforce the viewpoint of those who see the actions as politically motivated and indicative of broader issues of governance and accountability under the Obama administration.
+The choice of language, historical comparisons, and the leveraging of the author’s and platform’s ideological stances contribute to a biased presentation of the events surrounding the IRS scandal.
+Table 10: The Last 7 of 15 Republican Scandals News Articles Rated by $\mathsf{EVINCE}$ and its Justifications.
+The rating column starts with $\mathsf{EVINCE}$’s rating, the Democrat rater (in blue), and then Republican rater (in red).